### v4

# web_appendix

Abstract
We propose difference-in-differences estimators in designs where the treatment is continuously distributed at every period, as is often the case when one studies the effects of taxes, tariffs, or prices. We assume that between consecutive periods, the treatment of some units, the switchers, changes, while the treatment of other units remains constant. We show that under a placebo-testable parallel-trends assumption, averages of the slopes of switchers’ potential outcomes can be nonparametrically estimated. We generalize our estimators to the instrumental-variable case. We use our estimators to estimate the price-elasticity of gasoline consumption.
Keywords: differences-in-differences, continuous treatment, two-way fixed effects regressions, heterogeneous treatment effects, panel data, policy evaluation, instrumental variable.
JEL Codes: C21, C23
1 Introduction
A popular method to estimate the effect of a treatment on an outcome is to estimate a two-way fixed effects (TWFE) regression that controls for unit and time fixed effects:
$Y_{i,t}=\alpha_{i}+\gamma_{t}+\beta_{TWFE}D_{i,t}+u_{i,t},$
where $D_{i,t}$ is the treatment of unit $i$ at time $t$. de Chaisemartin and D’Haultfœuille (2023a) find that 26 of the 100 most cited papers published by the American Economic Review from 2015 to 2019 have estimated at least one TWFE regression. de Chaisemartin and D’Haultfœuille (2020), Goodman-Bacon (2021), and Borusyak et al. (2021) have shown that under a parallel trends assumption, TWFE regressions are not robust to heterogeneous effects: they may estimate a weighted sum of treatment effects across periods and units, with some negative weights. Owing to the negative weights, $\beta_{TWFE}$ could be, say, negative, even if the treatment effect is positive for every unit $\times$ period. Importantly, the result in de Chaisemartin and D’Haultfœuille (2020) applies to binary, discrete, and continuous treatments.
Several alternative heterogeneity-robust difference-in-difference (DID) estimators have been proposed (see Table 2 of de Chaisemartin and D’Haultfœuille, 2023b). Some apply to binary and staggered treatments (see Sun and Abraham, 2021; Callaway and Sant’Anna, 2021; Borusyak et al., 2021). Some apply to designs where all units start with a treatment equal to 00, and then get treated with heterogeneous, potentially continuously distributed treatment intensities (see de Chaisemartin and D’Haultfœuille, 2023a; Callaway et al., 2021; de Chaisemartin and D’Haultfœuille, 2024). However, treatments continuously distributed at every period, including the first one in the data, are ubiquitous in applied work. For instance, taxes (see Li et al., 2014) or tariffs (see Fajgelbaum et al., 2020) are often continuously distributed at all periods. No difference-in-difference estimator robust to heterogeneous effects is available for such designs. Proposing such estimators is the purpose of this paper.
We assume that we have a panel data set, whose units could be geographical locations such as counties. We start by considering the case where the panel has two time periods. From period one to two, the treatment of some units, hereafter referred to as the switchers, changes. On the other hand, the treatment of other units, hereafter referred to as the stayers, does not change. We consider two target parameters. The first one is the average slope of switchers’ period-two potential outcome function, from their period-one to their period-two treatment, hereafter referred to as the Average of Slopes (AS). Our second target is a weighted average of switchers’ slopes, where switchers receive a weight proportional to the absolute value of their treatment change, hereafter referred to as the Weighted Average of Slopes (WAS). We propose a novel parallel trends assumption on the outcome evolution of switchers and stayers with the same period-one treatment, in the counterfactual where switchers’ treatment would not have changed. Because it conditions on units’ period-one treatment, this parallel-trends assumption does not impose any restriction on effect heterogeneity. This parallel trends assumption is also placebo testable, by comparing switchers’ and stayers’ outcome evolutions before switchers’ treatment changes. We view the possibility of placebo-testing it as an important advantage of our assumption, as placebo tests are an essential step in establishing the credibility of an identifying assumption in observational studies (Imbens et al., 2001; Imbens and Xu, 2024). We show that under our placebo-testable parallel-trends assumption, the AS and the WAS are identified. This contrasts with other target parameters, like the dose-response function or the average marginal effect, which can only be identified under non-placebo-testable assumptions.
Economically, the AS and the WAS can serve different purposes, so neither parameter dominates the other. Under shape restrictions on the potential outcome function, the AS can be used to infer the effect of other treatment changes than those that took place from period one to two. Instead, the WAS can be used to conduct a cost-benefit analysis of the treatment changes that effectively took place. On the other hand, when it comes to estimation, the WAS unambiguously dominates the AS. First, we show that it can be estimated at the standard parametric rate even if switchers can experience an arbitrarily small change of their treatment between consecutive periods. Second, we show that under some conditions, the asymptotic variance of the WAS estimator is strictly lower than that of the AS estimator. Third, unlike the AS, the WAS is amenable to doubly-robust estimation, which comes with a number of advantages.
Then, we consider the instrumental-variable (IV) case. For instance, one may be interested in estimating the price-elasticity of a good’s consumption. If prices respond to demand shocks, the counterfactual consumption trends of units experiencing and not experiencing a price change may not be parallel. On the other hand, the counterfactual consumption trends of units experiencing and not experiencing a tax change may be parallel. Then, taxes may be used as an instrument for prices. In such cases, we show that the reduced-form WAS effect of the instrument on the outcome divided by the first-stage WAS effect is equal to a weighted average of switchers’ outcome-slope with respect to the treatment, where switchers with a larger first-stage effect receive more weight. Hereafter, we refer to this effect as the IV-WAS effect. The ratio of the reduced-form and first-stage AS effects is also equal to a weighted average of slopes, with arguably less natural weights, so in the IV case the WAS seems both economically and statistically preferable to the AS. Importantly, we show that the reduced-form parallel-trends assumption implicitly restricts treatment-effect heterogeneity. Such restrictions can be alleviated by controlling for groups baseline treatment in the IV specification, but the resulting estimator still restricts effects’ heterogeneity across units.
We consider other extensions. First, we extend our results to applications with more than two time periods. Importantly, with several time periods our estimators rely on a parallel-trends assumption over consecutive periods, rather than over the entire duration of the panel. Second, we propose a placebo estimator of the parallel-trends assumption underlying our estimators.
Finally, we use the yearly, 1966 to 2008 US state-level panel dataset of Li et al. (2014) to estimate the effect of gasoline taxes on gasoline consumption and prices. Using the WAS estimators, we find a significantly negative effect of taxes on gasoline consumption, and a significantly positive effect on prices. The AS estimators are close to, and not significantly different from, the WAS estimators, but they are also markedly less precise: their standard errors are almost three times larger than that of the WAS estimators. This shows that our theoretical result on the precision ranking of the WAS and AS estimators, which we derive under strong assumptions, still holds in a real-life application where those assumptions probably do not hold. The precision losses attached to using the AS have consequences. The AS estimator of the effect of taxes on prices is not statistically significant, so with that estimator one cannot use taxes as an instrument to estimate the price-elasticity of consumption, because the instrument does not have a first-stage. This contrasts with the WAS, whose first-stage t-stat is around 7. We compute an IV-WAS estimator of the price elasticity of gasoline consumption, and find a fairly small elasticity of -0.67, in line with previous literature (for instance, Hausman and Newey, 1995, find a long-run elasticity of -0.81). Our placebo estimators are small, insignificant, and fairly precisely estimated, thus suggesting that our parallel trends assumption is plausible.
Our estimators are computed by the did_multiplegt_stat Stata package, available from the SSC repository. Our package allows estimators with control variables and weights, see the help file and the package’s companion paper for further details.
Related Literature.
Our paper builds upon several previous papers in the panel data literature. Chamberlain (1982) seems to be the first paper to have proposed an estimator of the AS parameter. Under the assumption of no counterfactual time trend, the estimator therein is a before-after estimator. Then, our paper is closely related to the work of Graham and Powell (2012), who also propose DID estimators of the AS (see their Equation (21)) when the treatment is continuously distributed at every time period. Their estimators rely on a linear effect assumption and assume that units experience the same evolution of their treatment effect over time, a parallel-trends-on-treatment-effects assumption. By contrast, our estimator of the AS does not place any restriction on treatment effects. But our main contribution to this literature is to introduce the WAS, and to contrast the pros and cons of the AS and WAS estimators. Our results are also related to Hoderlein and White (2012), who consider the average marginal effect of a continuous treatment with panel data. However, their target parameters and identifying assumptions are different. For instance, they rule out systematic changes of the outcome over time.
With respect to the aforementioned heterogeneity-robust DID literature, we make two contributions. First, in the non-IV case we propose estimators that can be used even if units’ treatment varies at baseline. Thus we usefully complement previous literature, that has mostly focused on the case where all units have a baseline treatment equal to zero (see Sun and Abraham, 2021; Callaway and Sant’Anna, 2021; Borusyak et al., 2021; Callaway et al., 2021; de Chaisemartin and D’Haultfœuille, 2024). One exception predating this paper is de Chaisemartin and D’Haultfœuille (2020), who allow for a non-binary discrete treatment at baseline in their Web Appendix, and propose estimators comparing switchers and stayers with the same baseline treatment. However, that paper does not allow for continuously distributed treatments, and comparing switchers and stayers with the same baseline treatment is no longer feasible with a continuously distributed treatment. [^1] Another exception, posterior to this paper, is de Chaisemartin and D’Haultfœuille (2023a), who extend the estimators in this paper to models with dynamic effects in Section 1.10 of their Web Appendix. Second, in the IV case, previous IV-DID literature has only considered classical designs with two periods and binary instrument and treatment (de Chaisemartin, 2010; Hudson et al., 2017), as well as fuzzy DID designs, a special case of IV-DIDs (de Chaisemartin and D’Haultfœuille, 2018). Instead, this paper proposes broadly applicable IV-DID estimators, and also highlights that IV-DID estimators impose restrictions on treatment effect-heterogeneity, which can be mitigated by controlling for the baseline treatment.
Importantly, our estimators require that there be some stayers, whose treatment does not change between consecutive periods. This assumption is unlikely to be met when the treatment is say, precipitations: for instance, US counties never experience the exact same precipitations over two consecutive years. In de Chaisemartin et al. (2023), we discuss the (non-trivial) extension of the results in this paper to applications without stayers.
Organization of the paper.
In Section 2, we present the set-up, introduce notation and discuss our main assumptions. In Section 3, we introduce the AS and discuss its identification and estimation. Section 4 then turns to the WAS. Section 5 extends our previous results to an instrumental variable set-up. We consider other extensions in Section 6. Finally, our application is developed in Section 7. The proofs are collected in the appendix.
2 Set-up, assumptions, and building-block identification result
2.1 Set-up
A representative unit is drawn from an infinite super population, and observed at two time periods. This unit could be an individual or a firm, but it could also be a geographical unit, like a county or a region. [^2] In that case, one may want to weight the estimation by counties’ or regions’ populations. Extending the estimators we propose to allow for such weighting is a mechanical extension. All expectations below are taken with respect to the distribution of variables in the super population. We are interested in the effect of a continuous and scalar treatment variable on that unit’s outcome. Let $D_{1}$ (resp. $D_{2}$) denote the unit’s treatment at period 1 (resp. 2), and let $\mathcal{D}_{1}$ (resp. $\mathcal{D}_{2}$) be the set of values $D_{1}$ (resp. $D_{2}$) can take, i.e. its support. Let $S=1\{D_{2}\neq D_{1}\}$ be an indicator equal to 1 if the unit’s treatment changes from period one to two, i.e. if they are a switcher.
For any $d\in\mathcal{D}_{1}\cup\mathcal{D}_{2}$, let $Y_{1}(d)$ and $Y_{2}(d)$ respectively denote the unit’s potential outcomes at periods 1 and 2 with treatment $d$, and let $Y_{1}$ and $Y_{2}$ denote their observed outcomes at periods 1 and 2. Our potential outcome notation assumes that $Y_{1}$ does not depend on units’ period-two treatment, thus ruling out anticipation effects, a commonly-made assumption in the DID literature. Our notation also assumes that $Y_{2}$ does not depend on units’ period-one treatment, thus ruling out dynamic effects. When the treatment is continuously distributed at period one, allowing for dynamic effects opens up the so-called initial-conditions problem. As units receive heterogeneous doses at period one, they may have experienced treatment changes before period one. With dynamic effects such changes may still affect their outcome over the study period, but they cannot be accounted for because they are not observed. Ruling out dynamic effects allows us to abstract from this thorny issue, but could yield misleading results if dynamic effects are present. To alleviate this concern, in Section 6.3 we propose a modified version of our estimators, robust to dynamic effects up to a pre-specified treatment lag.
In what follows, all equalities and inequalities involving random variables are required to hold almost surely. Finally, for any random variable observed at the two time periods $(X_{1},X_{2})$, let $\Delta X=X_{2}-X_{1}$ denote the change of $X$ from period $1$ to $2$.
2.2 Assumptions
We make the following assumptions. [^3] Throughout the paper, we implicitly assume that all potential outcomes have an expectation.
Assumption 1
(Parallel trends) For all $d_{1}\in\mathcal{D}_{1}$, $E(\Delta Y(d_{1})|D_{1}=d_{1},D_{2})=E(\Delta Y(d_{1})|D_{1}=d_{1})$.
Assumption 1 is a parallel trends assumption, requiring that $\Delta Y(d_{1})$ be mean independent of $D_{2}$, conditional on $D_{1}=d_{1}$.
Assumption 2
(Bounded treatment, Lipschitz and bounded potential outcomes)
  1. 1.
    $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ are bounded subsets of $\mathbb{R}$.
  2. 2.
    For all $t\in\{1,2\}$ and for all $(d,d^{\prime})\in\mathcal{D}_{t}^{2}$, there is a random variable $\overline{Y}\geq 0$ such that $|Y_{t}(d)-Y_{t}(d^{\prime})|\leq\overline{Y}|d-d^{\prime}|$, with $\sup_{(d_{1},d_{2})\in\text{Supp}(D_{1},D_{2})}E[\overline{Y}|D_{1}=d_{1},D_{2}=d_{2}]<\infty$.
Assumption 2 ensures that all the expectations below are well defined. It requires that the set of values that the period-one and period-two treatments can take be bounded. It also requires that the potential outcome functions be Lipschitz (with a unit-specific Lipschitz constant). This will automatically hold if $d\mapsto Y_{2}(d)$ is differentiable with respect to $d$ and has a bounded derivative.
Finally, for estimation and inference we assume we observe an iid sample with the same distribution as $(Y_{1},Y_{2},D_{1},D_{2})$:
Assumption 3
(iid sample) We observe $(Y_{i,1},Y_{i,2},D_{i,1},D_{i,2})_{1\leq i\leq n}$, that are independent and identically distributed vectors with the same probability distribution as $(Y_{1},Y_{2},D_{1},D_{2})$.
Importantly, Assumption 3 allows for the possibility that $Y_{1}$ and $Y_{2}$ (resp. $D_{1}$ and $D_{2}$) are serially correlated, as is commonly assumed in DID studies (see Bertrand et al., 2004).
2.3 Building-block identification result
Assumption 1 implies the following lemma, our building-block identification result.
Lemma 1
If Assumption 1 holds, then for all $(d_{1},d_{2})\in\mathcal{D}_{1}\times\mathcal{D}_{2}$ such that $d_{1}\neq d_{2}$ and $P(S|D_{1}=d_{1})<1$,
$\text{TE}(d_{1},d_{2}|d_{1},d_{2}):=$ $E\left(\frac{Y_{2}(d_{2})-Y_{2}(d_{1})}{d_{2}-d_{1}}\middle|D_{1}=d_{1},D_{2}=d_{2}\right)$
$=$ $E\left(\frac{\Delta Y-E(\Delta Y|D_{1}=d_{1},S=0)}{d_{2}-d_{1}}\middle|D_{1}=d_{1},D_{2}=d_{2}\right).$
Proof:
$E\left(Y_{2}(d_{2})-Y_{2}(d_{1})\middle|D_{1}=d_{1},D_{2}=d_{2}\right)$
$=$ $E\left(\Delta Y\middle|D_{1}=d_{1},D_{2}=d_{2}\right)-E\left(\Delta Y(d_{1})\middle|D_{1}=d_{1},D_{2}=d_{2}\right)$
$=$ $E\left(\Delta Y\middle|D_{1}=d_{1},D_{2}=d_{2}\right)-E\left(\Delta Y(d_{1})\middle|D_{1}=d_{1},D_{2}=d_{1}\right)$
$=$ $E\left(\Delta Y\middle|D_{1}=d_{1},D_{2}=d_{2}\right)-E\left(\Delta Y\middle|D_{1}=d_{1},S=0\right)$
$=$ $E\left(\Delta Y-E\left(\Delta Y\middle|D_{1}=d_{1},S=0\right)\middle|D_{1}=d_{1},D_{2}=d_{2}\right),$
where the second equality follows from Assumption 1. This proves the result
Intuitively, under Assumption 1 the counterfactual outcome evolution switchers would have experienced if their treatment had not changed is identified by the outcome evolution of stayers with the same period-one treatment. If a unit’s treatment changes from two to five, we can recover its counterfactual outcome evolution if its treatment had not changed, by using the average outcome evolution of all stayers with a baseline treatment of two. Then, a DID estimand comparing switchers’ and stayers’ outcome evolutions identifies $E\left(Y_{2}(d_{2})-Y_{2}(d_{1})\middle|D_{1}=d_{1},D_{2}=d_{2}\right)$, and we can scale that effect by $d_{2}-d_{1}$ to identify a slope rather than an unnormalized effect.
Note that in a canonical DID design where $\mathcal{D}_{1}=0$ and $\mathcal{D}_{2}\in\{0,1\}$, Lemma 1 only applies to $(d_{1},d_{2})=(0,1)$, $\text{TE}(0,1|0,1)$ reduces to the ATT, and the estimand reduces to the canonical DID estimand comparing the outcome evolutions of treated and untreated units. Note also that in a design where all units are untreated at period one, $d_{1}=0$, and
$\text{TE}(0,d_{2}|0,d_{2})=E\left(\frac{Y_{2}(d_{2})-Y_{2}(0)}{d_{2}}\middle|D_{2}=d_{2}\right),$
an effect closely related to the $ATT(d|d)$ effect in Callaway et al. (2021). Thus, the effects we consider are extensions of those effects to applications with a treatment continuous at all periods.
Lemma 1 shows that under Assumption 1,
$(d_{1},d_{2})\mapsto\text{TE}(d_{1},d_{2}|d_{1},d_{2})$
is identified. Of course, one may be interested in other parameters, like
$(d,d^{\prime})\mapsto\text{TE}(d,d^{\prime}):=E\left(\frac{Y_{2}(d)-Y_{2}(d^{\prime})}{d-d^{\prime}}\right),$
a function which, unlike $(d_{1},d_{2})\mapsto\text{TE}(d_{1},d_{2}|d_{1},d_{2}),$ applies to the entire population rather than to specific subpopulation that depend on $(d_{1},d_{2})$. Alternatively, one could also be interested in the average marginal effect
$E\left(Y^{\prime}_{2}(D_{2})\right).$
What is the appeal of $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ with respect to those other parameters? Conditional on $D_{1}=d_{1},D_{2}=d_{2}$, $Y_{2}(d_{2})$ is observed, so estimating $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ only requires estimating $Y_{2}(d_{1})$, switchers’ counterfactual outcomes if their treatment had not changed. By definition, $Y_{1}(d_{1})$ is observed. If the data contains a third period 00 and the treatment of some units does not change from period 00 to $1$, then $Y_{0}(d_{1})$ is also observed for some switchers and stayers. Then, as explained in further details in Section 6.2, one can placebo-test Assumption 1, by comparing the outcome evolutions of switchers and stayers from period 00 to $1$. This shows that $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ is identified under a placebo-testable parallel-trends assumption. On the other hand, estimating $\text{TE}(d,d^{\prime})$ requires estimating, for most units, two unobserved counterfactual outcomes. This cannot be achieved under a placebo-testable assumption as we only observe one potential outcome at each date. When $D_{1}=0$, Callaway et al. (2021) propose a “strong parallel-trends” assumption under which the dose-response function, a parameter closely related to $\text{TE}(0,d^{\prime})$ are identified, but their “strong parallel-trends” assumption is not placebo testable. Similarly, estimating $E\left(\frac{Y_{2}(D_{2})-Y_{2}(d^{\prime})}{D_{2}-d^{\prime}}\right)$ requires estimating $Y_{2}(d^{\prime})$, which cannot be achieved under a placebo-testable assumption because $Y_{1}(d^{\prime})$ is not observed for all units. As $Y^{\prime}_{2}(D_{2})=\lim_{d^{\prime}\rightarrow D_{2}}\frac{Y_{2}(D_{2})-Y_{2}(d^{\prime})}{D_{2}-d^{\prime}}$, the same issue applies to $E\left(Y^{\prime}_{2}(D_{2})\right)$.
Variability in $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ across values of $(d_{1},d_{2})$ conflates a dose-response relationship that may be of economic interest, and a selection bias due to the fact that units with different period one and two treatments may have heterogeneous treatment effects (Callaway et al., 2021). Moreover, Lemma 1 shows that estimating $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ requires estimating the values of two conditional expectations with respect to continuous variables, at points $D_{1}=d_{1},D_{2}=d_{2}$ and $D_{1}=d_{1}$. Unless one is willing to make parametric functional-form assumptions, the resulting estimator will converge at a slower rate than the standard $\sqrt{n}-$ parametric rate. For these two reasons, in this paper we focus on averages of the slopes $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$, that can be estimated non-parametrically at the standard $\sqrt{n}-$ parametric rate, and we do not focus on the function $(d_{1},d_{2})\mapsto\text{TE}(d_{1},d_{2}|d_{1},d_{2}).$
Finally, our DID estimands compare switchers and stayers with the same period-one treatment. Instead, one could propose estimands comparing switchers and stayers, without conditioning on their period-one treatment. To recover the counterfactual outcome trend of a switcher going from two to five units of treatment, one could use a stayer with treatment equal to three at both dates. On top of Assumption 1, such estimands rest on two supplementary conditions:
  1. (i)
    $E(\Delta Y(d)|D_{1}=d)=E(\Delta Y(d))$.
  2. (ii)
    For all $(d,d^{\prime})\in\mathcal{D}_{1}^{2}$, $E(\Delta Y(d))=E(\Delta Y(d^{\prime}))$.
(i) requires that all units experience the same evolution of their potential outcome with treatment $d$, while Assumption 1 only imposes that requirement for units with the same baseline treatment. Assumption 1 may be more plausible: units with the same period-one treatment may be more similar and more likely to be on parallel trends than units with different period-one treatments. (ii) requires that the trend affecting all potential outcomes be the same: to rationalize a DID estimand comparing a switcher going from two to five units of treatment to a stayer with treatment equal to three, $E(\Delta Y(2))$ and $E(\Delta Y(3))$ should be equal. Rearranging, (ii) is equivalent to
$E(Y_{2}(d)-Y_{2}(d^{\prime}))=E(Y_{1}(d)-Y_{1}(d^{\prime})):$ (1)
the treatment effect should be constant over time, a strong restriction on treatment effect heterogeneity. Assumption 1, on the other hand, does not impose any restriction on treatment effect heterogeneity, as it only restricts one potential outcome per unit.
3 Estimating the average of switchers’ slopes
3.1 Target parameter
In this section, our target parameter is
$\delta_{1}$ $:=E\left(\frac{Y_{2}(D_{2})-Y_{2}(D_{1})}{D_{2}-D_{1}}\middle|S=1\right),$ (2)
the average of the slopes of switchers’ potential outcome functions, between their period-one and their period-two treatments. Hereafter, $\delta_{1}$ is referred to as the Average of Slopes (AS).
The AS is a local effect: it only applies to switchers, and it measures the effect of changing their treatment from its period-one to its period-two value, not of other changes of their treatment. Still, the AS can be used to point or partially identify the effect of other treatment changes under shape restrictions. First, assume that the potential outcomes are linear: for $t\in\{1,2\}$,
$Y_{t}(d)=Y_{t}(0)+B_{t}d,$
where $B_{t}$ is a slope that may vary across units and may change over time. Then, $\delta_{1}=E\left(B_{2}\middle|S=1\right)$: the AS is equal to the average, across switchers, of the slopes of their potential outcome functions at period 2. Therefore, for all $d\neq d^{\prime}$,
$E(Y_{2}(d)-Y_{2}(d^{\prime})|S=1)=(d-d^{\prime})\delta_{1}:$
under linearity, knowing the AS is sufficient to recover the ATE of any uniform treatment change among switchers. Of course, this only holds under linearity, which may not be a plausible assumption. Assume instead that $d\mapsto Y_{2}(d)$ is convex. Then, for any $\epsilon>0$,
$E\left(Y_{2}(D_{2}+\epsilon)-Y_{2}(D_{2})\middle|S=1\right)\geq\epsilon\delta_{1}.$
Accordingly, under convexity one can use the AS to obtain lower bounds of the effect of changing the treatment from $D_{2}$ to larger values than $D_{2}$. For instance, in Fajgelbaum et al. (2020), one can use this strategy to derive a lower bound of the effect of increasing tariffs’ to even higher levels than those decided by the Trump administration. Under convexity, one can also use the AS to derive an upper bound of the effect of changing the treatment from $D_{1}$ to a lower value than $D_{1}$. And under concavity, one can derive an upper (resp. lower) bound of the effect of changing the treatment from $D_{2}$ (resp. $D_{1}$) to a larger (resp. lower) value. [^4] See D’Haultfœuille et al. (2023) for bounds of the same kind obtained under concavity or convexity. Importantly, the AS is identified even if those linearity or convexity/concavity conditions fail. But those non-placebo testable conditions are necessary to use the AS to identify or bound the effects of alternative policies.
3.2 Identification
To identify the AS, we use a DID estimand comparing switchers and stayers with the same period-one treatment. This requires that there be no value of the period-one treatment $D_{1}$ such that only switchers have that value, as stated formally below.
Assumption 4
(Support condition for AS identification) $P(S=1)>0$, $P(S=1|D_{1})<1$.
Assumption 4 implies that $P(S=0)>0$, meaning that there are stayers whose treatment does not change. While we assume that $D_{1}$ and $D_{2}$ are continuous, we also assume that the treatment is persistent, and thus $\Delta D$ has a mixed distribution with a mass point at zero.
To identify the AS, we also start by assuming that there are no quasi-stayers: the treatment of all switchers changes by at last $c$ from period one to two, for some strictly positive $c$.
Assumption 5
(No quasi-stayers) $\exists c>0$: $P(|\Delta D|>c|S=1)=1$.
We relax Assumption 5 just below.
Theorem 1
If Assumptions 1-5 hold,
$\delta_{1}$ $=E\left(\frac{\Delta Y-E(\Delta Y|D_{1},S=0)}{\Delta D}\middle|S=1\right).$
If there are quasi-stayers, the AS is still identified. For any $\eta>0$, let $S_{\eta}=1\{|\Delta D|>\eta\}$ be an indicator for switchers whose treatment changes by at least $\eta$ from period one to two.
Theorem 2
If Assumptions 1-4 hold,
$\delta_{1}$ $=\lim_{\eta\downarrow 0}E\left(\frac{\Delta Y-E(\Delta Y|D_{1},S=0)}{\Delta D}\middle|S_{\eta}=1\right).$
If there are quasi-stayers whose treatment change is arbitrarily close to 00 (i.e. $f_{|\Delta D||S=1}(0)>0$), the denominator of $(\Delta Y-E(\Delta Y|D_{1},S=0))/\Delta D$ is close to 00 for them. On the other hand,
$\Delta Y-E(\Delta Y|D_{1},S=0)$
$=$ $Y_{2}(D_{2})-Y_{2}(D_{1})+\Delta Y(D_{1})-E(\Delta Y(D_{1})|D_{1},S=0)$
$\approx$ $\Delta Y(D_{1})-E(\Delta Y(D_{1})|D_{1},S=0),$
so the ratio’s numerator may not be close to 00. Then, under weak conditions,
$E\left(\left|\frac{\Delta Y-E(\Delta Y|D_{1},S=0)}{\Delta D}\right|\;\middle|S=1\right)=+\infty.$
Therefore, we need to trim quasi-stayers from the estimand in Theorem 1, and let the trimming go to 0, as in Graham and Powell (2012) who consider a related estimand with some quasi-stayers. Accordingly, with quasi-stayers the AS is irregularly identified by a limiting estimand.
3.3 Estimation and inference
With no quasi-stayers, $E\left((\Delta Y-E(\Delta Y|D_{1},S=0))/\Delta D\middle|S=1\right)$ can be estimated in three steps. First, one estimates $E(\Delta Y|D_{1},S=0)$ using a non-parametric regression of $\Delta Y_{i}$ on $D_{i,1}$ among stayers. Second, for each switcher, one computes $\widehat{E}(\Delta Y|D_{1}=D_{i,1},S=0)$, its predicted outcome evolution given its baseline treatment, according to the non-parametric regression estimated among stayers. Third, one lets
$\widehat{\delta}_{1}:=$ $\frac{1}{n_{s}}\sum_{i:S_{i}=1}\frac{\Delta Y_{i}-\widehat{E}(\Delta Y|D_{1}=D_{i,1},S=0)}{\Delta D_{i}},$
where $n_{s}=\#\{i:S_{i}=1\}$.
To estimate $E(\Delta Y|D_{1},S=0)$, we consider a series estimator based on polynomials in $D_{1}$, $(p_{k,K_{n}}(D_{1}))_{1\leq k\leq K_{n}}$. We make the following technical assumption.
Assumption 6
(Conditions for asymptotic normality of AS estimator)
  1. 1.
    $D_{1}$ is continuously distributed on a compact interval $I$, with $\inf_{d\in I}f_{D_{1}}(d)>0$.
  2. 2.
    $E[\Delta Y^{2}]<\infty$ and $d\mapsto E[\Delta Y^{2}|D_{1}=d]$ is bounded on $I$.
  3. 3.
    $P(S=1)>0$ and $\sup_{d\in I}P(S=1|D_{1}=d)<1$.
  4. 4.
    The functions $d\mapsto E[(1-S)\Delta Y|D_{1}=d]$, $d\mapsto E[S|D_{1}=d]$ and $d\mapsto E[S/\Delta D|D_{1}=d]$ are four times continuously differentiable.
  5. 5.
    The polynomials $d\mapsto p_{k,K_{n}}(d)$, $1\leq k\leq K_{n}$, are orthonormal on $I$ and $K_{n}^{12}/n\to+\infty$, $K_{n}^{7}/n\to 0$.
Point 3 is a slight reinforcement of Assumption 4. In Point 5, $K_{n}^{12}/n\to\infty$ requires that $K_{n}$, the order of the polynomial in $D_{1}$ we use to approximate $E(\Delta Y|D_{1},S=0)$, goes to +$\infty$+ ∞ when the sample size grows, thus ensuring that the bias of our series estimator of $E(\Delta Y|D_{1},S=0)$ tends to zero. $K_{n}^{7}/n\to 0$ ensures that $K_{n}$ does not go to infinity too fast, thus preventing overfitting.
Theorem 3
If Assumptions 1-3 and 5-6 hold,
$\sqrt{n}\left(\widehat{\delta}_{1}-\delta_{1}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{1})),$
where
$\psi_{1}$ $:=\frac{1}{E(S)}\left\{\left(\frac{S}{\Delta D}-E\left(\left.\frac{S}{\Delta D}\right|D_{1}\right)\frac{(1-S)}{E[1-S|D_{1}]}\right)[\Delta Y-E(\Delta Y|D_{1},S=0)]-\delta_{1}S\right\}.$
Theorem 3 shows that without quasi-stayers, the AS can be estimated at the $\sqrt{n}-$rate, and gives an expression of its estimator’s asymptotic variance. With quasi-stayers, we conjecture that the AS cannot be estimated at the $\sqrt{n}-$rate. This conjecture is based on a result from Graham and Powell (2012). Though their result applies to a broader class of estimands, it implies in particular that with quasi-stayers,
$\lim_{\eta\downarrow 0}E\left(\frac{\Delta Y-E(\Delta Y|S=0)}{\Delta D}\middle|S_{\eta}=1\right)$
cannot be estimated at a faster rate than $n^{1/3}$. The estimand in the previous display is closely related to our estimand
$\lim_{\eta\downarrow 0}E\left(\frac{\Delta Y-E(\Delta Y|D_{1},S=0)}{\Delta D}\middle|S_{\eta}=1\right)$
in Theorem 2, and is equal to it if $E(\Delta Y|D_{1},S=0)=E(\Delta Y|S=0)$. Then, even though the assumptions in Graham and Powell (2012) differ from ours, it seems reasonable to assume that their general conclusion still applies to our set-up: here as well, owing to $\delta_{1}$’s irregular identification, this parameter can probably not be estimated at the parametric $\sqrt{n}-$rate with quasi-stayers. This is one of the reasons that lead us to consider, in the next section, another target parameter that can be estimated at the parametric $\sqrt{n}-$rate with quasi-stayers.
4 Estimating a weighted average of switchers’ slopes
4.1 Target parameter
In this section, our target parameter is
$\delta_{2}:=$ $E\left(\frac{|D_{2}-D_{1}|}{E(|D_{2}-D_{1}||S=1)}\times\frac{Y_{2}(D_{2})-Y_{2}(D_{1})}{D_{2}-D_{1}}\middle|S=1\right)$
$=$ $\frac{E\left(\text{sgn}(D_{2}-D_{1})(Y_{2}(D_{2})-Y_{2}(D_{1}))\middle|S=1\right)}{E(|D_{2}-D_{1}||S=1)}$
$=$ $\frac{E\left(\text{sgn}(D_{2}-D_{1})(Y_{2}(D_{2})-Y_{2}(D_{1}))\right)}{E(|D_{2}-D_{1}|)}.$
$\delta_{2}$ is a weighted average of the slopes of switchers’ potential outcome functions from their period-one to their period-two treatments, where slopes receive a weight proportional to switchers’ absolute treatment change from period one to two. Accordingly, we refer to $\delta_{2}$ as the Weighted Average of Slopes (WAS). $\delta_{2}=\delta_{1}$ if and only if
$\text{cov}\left(\frac{Y_{2}(D_{2})-Y_{2}(D_{1})}{D_{2}-D_{1}},|D_{2}-D_{1}|\middle|S=1\right)=0:$ (3)
the WAS and AS are equal if and only if switchers’ slopes are uncorrelated with $|D_{2}-D_{1}|$.
Economically, the AS and WAS serve different purposes. As discussed above, under shape restrictions on the potential outcome function, the AS can be used to identify or bound the effect of other treatment changes than the actual change switchers experienced from period one to two. The WAS cannot serve that purpose, but under some assumptions, it may be used to conduct a cost-benefit analysis of the treatment changes that took place from period one to two. To simplify the discussion, let us assume in the remainder of this paragraph that $D_{2}\geq D_{1}$. Assume also that the outcome is a measure of output, such as agricultural yields or wages, expressed in monetary units. Finally, assume that the treatment is costly, with a cost linear in dose, uniform across units, and known to the analyst: the cost of giving $d$ units of treatment to a unit at period $t$ is $c_{t}\times d$ for some known $(c_{t})_{t\in\{1,2\}}$. Then, $D_{2}$ is beneficial relative to $D_{1}$ if and only if $E(Y_{2}(D_{2})-c_{2}D_{2})>E(Y_{2}(D_{1})-c_{2}D_{1})$ or, equivalently,
$\delta_{2}>c_{2}:$
comparing $\delta_{2}$ to $c_{2}$ is sufficient to evaluate if changing the treatment from $D_{1}$ to $D_{2}$ was beneficial.
4.2 Identification
Let $S_{+}=1\{D_{2}-D_{1}>0\}$, $S_{-}=1\{D_{2}-D_{1}<0\}$ and
$\delta_{2+}$ $:=\frac{E\left(Y_{2}(D_{2})-Y_{2}(D_{1})|S_{+}=1\right)}{E(D_{2}-D_{1}|S_{+}=1)},$
$\delta_{2-}$ $:=\frac{E\left(Y_{2}(D_{1})-Y_{2}(D_{2})|S_{-}=1\right)}{E(D_{1}-D_{2}|S_{-}=1)}.$
Hereafter, units with $S_{+}=1$ are referred to as “switchers up”, while units with $S_{-}=1$ are referred to as “switchers down”. Thus, $\delta_{2+}$ is the WAS of switchers up, and $\delta_{2-}$ is the WAS of switchers down. One has
$\delta_{2}=$ $\frac{P(S_{+}=1|S=1)E(D_{2}-D_{1}|S_{+}=1)}{E(|D_{2}-D_{1}||S=1)}\delta_{2+}$
$+$ $\frac{P(S_{-}=1|S=1)E(D_{1}-D_{2}|S_{-}=1)}{E(|D_{2}-D_{1}||S=1)}\delta_{2-}.$ (4)
To identify $\delta_{2+}$ (resp. $\delta_{2-}$) we use DID estimands comparing switchers up (resp. switchers down) to stayers with the same period-one treatment. This requires that there be no value of $D_{1}$ such that some switchers up (resp. switchers down) have that baseline treatment while there is no stayer with the same baseline treatment, as stated in Point 1 (resp. 2) of Assumption 7 below.
Assumption 7
(Support conditions for WAS identification)
  1. 1.
    $0<P(S_{+}=1)$, and $0<P(S_{+}=1|D_{1})$ implies that $0<P(S=0|D_{1})$.
  2. 2.
    $0<P(S_{-}=1)$, and $0<P(S_{-}=1|D_{1})$ implies that $0<P(S=0|D_{1})$.
Theorem 4
  1. 1.
    If Assumptions 1-2 and Point 1 of Assumption 7 hold,
    $\delta_{2+}$ $=\frac{E\left(\Delta Y-E(\Delta Y|D_{1},S=0)\middle|S_{+}=1\right)}{E(\Delta D|S_{+}=1)}$ (5)
    $=\frac{E\left(\Delta Y|S_{+}=1\right)-E\left(\Delta Y\frac{P(S_{+}=1|D_{1})}{P(S=0|D_{1})}\frac{P(S=0)}{P(S_{+}=1)}\middle|S=0\right)}{E(\Delta D|S_{+}=1)}.$ (6)
  2. 2.
    If Assumptions 1-2 and Point 2 of Assumption 7 hold,
    $\delta_{2-}$ $=\frac{E\left(\Delta Y-E(\Delta Y|D_{1},S=0)\middle|S_{-}=1\right)}{E(\Delta D|S_{-}=1)}$ (7)
    $=\frac{E\left(\Delta Y|S_{-}=1\right)-E\left(\Delta Y\frac{P(S_{-}=1|D_{1})}{P(S=0|D_{1})}\frac{P(S=0)}{P(S_{-}=1)}\middle|S=0\right)}{E(\Delta D|S_{-}=1)}.$ (8)
  3. 3.
    If Assumptions 1-2 and Assumption 7 hold,
    $\delta_{2}=$ $\frac{E\left[\text{sgn}(\Delta D)\left(\Delta Y-E(\Delta Y|D_{1},S=0)\right)\right]}{E[|\Delta D|]}$ (9)
    $=$ $\frac{E\left[\text{sgn}(\Delta D)\Delta Y\right]-E\left[\Delta Y\frac{P(S_{+}=1|D_{1})-P(S_{-}=1|D_{1})}{P(S=0|D_{1})}P(S=0)\middle|S=0\right]}{E[|\Delta D|]}.$ (10)
Point 1 of Theorem 4 shows that $\delta_{2+}$, the WAS of switchers-up, is identified by two estimands, a regression-based and a propensity-score-based estimand. Point 2 of Theorem 4 shows that $\delta_{2-}$, the WAS of switchers down, is identified by two estimands similar to those identifying $\delta_{2+}$, replacing $S_{+}$ by $S_{-}$. Finally, if the conditions in Point 1 and 2 of Theorem 4 jointly hold, it directly follows from (4.2) that $\delta_{2}$, the WAS of all switchers, is identified by a weighted average of the estimands in Equations (5) and (7), and by a weighted average of the estimands in Equations (6) and (8). Those weighted averages simplify into the expressions given in Point 3 of Theorem 4. Point 3 of Theorem 4 also implies that $\delta_{2}$ is identified by the following doubly-robust estimand:
$\frac{E\left[\left(S_{+}-S_{-}-\frac{P(S_{+}=1|D_{1})-P(S_{-}=1|D_{1})}{P(S=0|D_{1})}(1-S)\right)(\Delta Y-E(\Delta Y|D_{1},S=0))\right]}{E[\left|\Delta D\right|]}.$ (11)
4.3 Estimation and inference
The regression-based estimands identifying $\delta_{2+}$ and $\delta_{2-}$ can be estimated following almost the same steps as in Section 3.3. Specifically, let
$\widehat{\delta}^{r}_{2+}:=$ $\frac{\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\left(\Delta Y_{i}-\widehat{E}(\Delta Y|D_{1}=D_{i,1},S=0)\right)}{\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta D_{i}}$
$\widehat{\delta}^{r}_{2-}:=$ $\frac{\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\left(\Delta Y_{i}-\widehat{E}(\Delta Y|D_{1}=D_{i,1},S=0)\right)}{\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\Delta D_{i}},$
where $n_{+}=\#\{i:S_{i+}=1\}$ and $n_{-}=\#\{i:S_{i-}=1\}$, and where $\widehat{E}(\Delta Y|D_{1},S=0)$ is the series estimator of $E(\Delta Y|D_{1},S=0)$ defined in Section 3.3 of the paper. Then, let
$\widehat{w}_{+}=\frac{\frac{n_{+}}{n}\times\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta D_{i}}{\frac{n_{+}}{n}\times\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta D_{i}-\frac{n_{-}}{n}\times\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\Delta D_{i}},$
and let
$\widehat{\delta}^{r}_{2}=$ $\widehat{w}_{+}\widehat{\delta}^{r}_{2+}+(1-\widehat{w}_{+})\widehat{\delta}^{r}_{2-}$
be the corresponding estimator of $\delta_{2}$.
We now propose estimators of the propensity-score-based estimands identifying $\delta_{2+}$ and $\delta_{2-}$ in Equations (6) and (8). Let $\widehat{P}(S_{+}=1)=n_{+}/n$ (resp. $\widehat{P}(S_{-}=1)=n_{-}/n$, $\widehat{P}(S=0)=(n-n_{s})/n$) be an estimator of $P(S_{+}=1)$ (resp. $P(S_{-}=1)$, $P(S=0)$). Let $\widehat{P}(S_{+}=1|D_{1})$ (resp. $\widehat{P}(S_{-}=1|D_{1})$, $\widehat{P}(S=0|D_{1})$) be a non-parametric estimator of $P(S_{+}=1|D_{1})$ (resp. $P(S_{-}=1|D_{1})$, $P(S=0|D_{1})$) using a series logistic regression of $S_{i+}$ (resp. $S_{i-}$, $1-S_{i}$) on polynomials in $D_{1}$ $(p_{k,K_{n}}(D_{1}))_{1\leq k\leq K_{n}}$. We make the following technical assumption.
Assumption 8
(Conditions for asymptotic normality of propensity-score WAS estimator)
  1. 1.
    $D_{1}$ is continuously distributed on a compact interval $I$, with $\inf_{d\in I}f_{D_{1}}(d)>0$.
  2. 2.
    $E[\Delta Y^{2}]<\infty$ and $d\mapsto E[\Delta Y^{2}|D_{1}=d]$ is bounded on $I$
  3. 3.
    $10<E[S_{+}]<$ 1, $10<E[S_{-}]<$ 1, $E[S]>0$ and $\sup_{d\in I}E[S|D_{1}=d]<1$.
  4. 4.
    The functions $d\mapsto E[\Delta Y(1-S)|D_{1}=d]$, $d\mapsto E[S|D_{1}=d]$, $d\mapsto E[S_{+}|D_{1}=d]$ and $d\mapsto E[S_{-}|D_{1}=d]$ are four times continuously differentiable.
  5. 5.
    The polynomials $d\mapsto p_{k,K_{n}}(d)$, $k\leq 1\leq K_{n}$ are orthonormal on $I$ and $K_{n}=Cn^{\nu}$ where $61/10<\nu<1/$ 6.
Let
$\widehat{\delta}^{ps}_{2+}:=$ $\frac{\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta Y_{i}-\frac{1}{n-n_{s}}\sum_{i:S_{i}=0}\Delta Y_{i}\frac{\widehat{P}(S_{+}=1|D_{1}=D_{i1})}{\widehat{P}(S=0|D_{1}=D_{i1})}\frac{\widehat{P}(S=0)}{\widehat{P}(S_{+}=1)}}{\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta D_{i}}$
$\widehat{\delta}^{ps}_{2-}:=$ $\frac{\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\Delta Y_{i}-\frac{1}{n-n_{s}}\sum_{i:S_{i}=0}\Delta Y_{i}\frac{\widehat{P}(S_{-}=1|D_{1}=D_{i1})}{\widehat{P}(S=0|D_{1}=D_{i1})}\frac{\widehat{P}(S=0)}{\widehat{P}(S_{-}=1)}}{\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\Delta D_{i}},$
and let
$\widehat{\delta}^{ps}_{2}=$ $\widehat{w}_{+}\widehat{\delta}^{ps}_{2+}+(1-\widehat{w}_{+})\widehat{\delta}^{ps}_{2-}$
be the corresponding estimator of $\delta_{2}$. Let
$\psi_{2+}$ $:=\frac{1}{E(\Delta DS_{+})}\left\{\left(S_{+}-E(S_{+}|D_{1})\frac{(1-S)}{E(1-S|D_{1})}\right)(\Delta Y-E(\Delta Y|D_{1},S=0))-\delta_{2+}\Delta DS_{+}\right\}$
$\psi_{2-}$ $:=\frac{1}{E(\Delta DS_{-})}\left\{\left(S_{-}-E(S_{-}|D_{1})\frac{(1-S)}{E(1-S|D_{1})}\right)(\Delta Y-E(\Delta Y|D_{1},S=0))-\delta_{2-}\Delta DS_{-}\right\}$
$\psi_{2}$ $:=\frac{1}{E(\left|\Delta D\right|)}\left\{\left(S_{+}-S_{-}-E(S_{+}-S_{-}|D_{1})\frac{(1-S)}{E(1-S|D_{1})}\right)\right.$
×$\qquad\times(\Delta Y-E(\Delta Y|D_{1},S=0))-\delta_{2}\left|\Delta D\right|\bigg{\}}.$
Theorem 5
  1. 1.
    If Assumptions 1-3 and 6 hold,
    $\sqrt{n}\left((\widehat{\delta}^{r}_{2+},\widehat{\delta}^{r}_{2-})^{\prime}-(\delta_{2+},\delta_{2-})^{\prime}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V((\psi_{2+},\psi_{2-})^{\prime})).$
    and
    $\sqrt{n}\left(\widehat{\delta}^{r}_{2}-\delta_{2}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{2})).$
  2. 2.
    If Assumptions 1-3 and 8 hold,
    $\sqrt{n}\left((\widehat{\delta}^{ps}_{2+},\widehat{\delta}^{ps}_{2-})^{\prime}-(\delta_{2+},\delta_{2-})^{\prime}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V((\psi_{2+},\psi_{2-})^{\prime})).$
    and
    $\sqrt{n}\left(\widehat{\delta}^{ps}_{2}-\delta_{2}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{2})).$
Based on (11), we can also estimate $\delta_{2}$ using the following doubly-robust estimator:
$\hat{\delta}^{dr}_{2}=\frac{\sum_{i}\left(S_{i+}-S_{i-}-\frac{\hat{P}(S_{+}=1|D_{1}=D_{1i})-P(S_{i-}=1|D_{1}=D_{1i})}{P(S_{i}=0|D_{1}=D_{1i})}(1-S_{i})\right)(\Delta Y_{i}-\hat{E}(\Delta Y_{i}|D_{1}=D_{1i},S_{i}=0))}{\sum_{i}\left|\Delta D_{i}\right|}.$
This estimator has an important advantage. While our regression-based (resp. propensity-score-based) estimator is nominally non-parametric, in practice it requires choosing a polynomial order to estimate $E(\Delta Y|D_{1},S=0)$ (resp. $P(S_{+}=1|D_{1})$ and $P(S_{-}=1|D_{1})$), and the rate conditions on $K_{n}$ in Assumptions 6 (resp. 8) do not give specific guidance on the choice of this tuning parameter. With the doubly-robust estimator above, one can choose this tuning parameter in a data-driven manner, using cross-validation (CV). Results in Section 4 of Andrews (1991) imply that a series estimator of a nonparametric regression model with a polynomial order chosen by CV is optimal: the ratio of its mean-squared error and that of an oracle estimator using the best polynomial order given the sample size converges to one. Then, as $D_{1}$ is a scalar variable, series estimators of $E(\Delta Y|D_{1},S=0)$, $P(S_{+}=1|D_{1})$, and $P(S_{-}=1|D_{1})$ with CV-chosen polynomial orders converge at a rate strictly faster than $n^{1/4}$, as long as one assumes that those nuisance functions are twice differentiable. Then, we conjecture that one can show, following arguments similar to those in Farrell (2015), that $\hat{\delta}^{dr}_{2}$ with CV-chosen polynomial orders in the estimation of the nuisance functions is $\sqrt{n}-$consistent, with asymptotic variance $V(\psi_{2})$.
Finally, we now show that under some assumptions, the asymptotic variance of the WAS estimator is lower than that of the AS estimator.
Proposition 1
If Assumption 1 holds, $(Y_{2}(D_{2})-Y_{2}(D_{1}))/(D_{2}-D_{1})=\delta$ for some real number $\delta$, $V(\Delta Y(D_{1})|D_{1},D_{2})=\sigma^{2}$ for some real number $\sigma^{2}>0$, $D_{2}\geq D_{1}$, and $\Delta D\perp\!\!\!\perp D_{1}$,
$V(\psi_{1})=$ $\sigma^{2}\left[\frac{E(1/(\Delta D)^{2}|S=1)}{P(S=1)}+\frac{\left(E(1/\Delta D|S=1)\right)^{2}}{P(S=0)}\right]$
$\geq$ $\sigma^{2}\frac{1}{\left(E(\Delta D|S=1)\right)^{2}}\left[\frac{1}{P(S=1)}+\frac{1}{P(S=0)}\right]=V(\psi_{2}),$
with equality if and only if $V(\Delta D|S=1)=0$.
Of course, the constant treatment effect and the homoscedasticity assumptions underlying Proposition 1 are strong, but one often has to make strong assumptions to be able to rank estimators’ variances. The question then is whether this ranking still holds in real-life applications, where those assumptions are unlikely to hold. Put differently, all models are wrong but some are useful, and the question is whether Proposition 1 is useful. In our empirical application, we find that the variance of $\hat{\delta}_{1}$ is indeed much larger than that of $\hat{\delta}^{dr}_{2}$, as predicted by Proposition 1.
5 Instrumental-variable estimation
There are instances where the parallel-trends condition in Assumption 1 is implausible, but one has at hand an instrument satisfying a similar parallel-trends condition. For instance, one may be interested in estimating the price-elasticity of a good’s consumption, but prices respond to supply and demand shocks, and therefore do not satisfy Assumption 1. On the other hand, taxes may not respond to supply and demand shocks and may satisfy a parallel-trends assumption.
5.1 Notation and assumptions
Let $(Z_{1},Z_{2})$ denote the instrument’s values at period one and two and $\mathcal{Z}_{t}$ be the support of $Z_{t}$. For any $z\in\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$, let $D_{1}(z)$ and $D_{2}(z)$ respectively denote the unit’s potential treatments at periods 1 and 2 with instrument $z$. Let $SC=1\{D_{2}(Z_{2})\neq D_{2}(Z_{1}),Z_{2}\neq Z_{1}\}$ be an indicator equal to 1 for switchers-compliers, namely units whose instrument changes from period one to two and whose treatment is affected by that change in the instrument.
We replace Assumption 1 by the following assumption. [^5] Note that with our notation where potential outcomes do not depend on $z$, we also implicitly impose the usual exclusion restriction.
Assumption 9
(Reduced-form and first-stage parallel trends) For all $z\in\mathcal{Z}_{1}$,
  1. 1.
    $E(Y_{2}(D_{2}(z))-Y_{1}(D_{1}(z))|Z_{1}=z,Z_{2},D_{1})=E(Y_{2}(D_{2}(z))-Y_{1}(D_{1}(z))|Z_{1}=z,D_{1})$.
  2. 2.
    $E(D_{2}(z)-D_{1}(z)|Z_{1}=z,Z_{2},D_{1})=E(D_{2}(z)-D_{1}(z)|Z_{1}=z,D_{1})$.
Point 1 of Assumption 9 requires that $Y_{2}(D_{2}(z))-Y_{1}(D_{1}(z))$, units’ outcome evolutions in the counterfactual where their instrument does not change from period one to two, be mean independent of $Z_{2}$, conditional on $Z_{1}$ and $D_{1}$. Unlike Assumption 1, this condition imposes some restrictions on treatment effect heterogeneity, and the goal of conditioning on $D_{1}$ is to minimize the stringency of those restrictions. To see this, note that the two following conditions are sufficient for Point 1 of Assumption 9 to hold:
$E(Y_{2}(D_{1}(z))-Y_{1}(D_{1}(z))|Z_{1}=z,Z_{2},D_{1})=E(Y_{2}(D_{1}(z))-Y_{1}(D_{1}(z))|Z_{1}=z,D_{1})$ (12)
$E(Y_{2}(D_{2}(z))-Y_{2}(D_{1}(z))|Z_{1}=z,Z_{2},D_{1})=E(Y_{2}(D_{2}(z))-Y_{2}(D_{1}(z))|Z_{1}=z,D_{1}).$ (13)
(12) requires that $Y_{2}(D_{1}(z))-Y_{1}(D_{1}(z))$, units’ outcome evolutions in the counterfactual where their instrument and their treatment does not change from period one to two, be mean independent of $Z_{2}$, conditional on $Z_{1}$ and $D_{1}$. Thanks to the conditioning on $D_{1}$, (12) is a standard parallel trends assumption that does not impose any restriction on treatment effect heterogeneity, like Assumption 1. If $D_{1}$ was not conditioned upon, (12) would require parallel trends among units with different baseline treatments, which implicitly assumes homogeneous treatment effects over time, as discussed in Section 2. (13), on the other hand, is a restriction on treatment effect heterogeneity across units. Essentially, it requires that switching the treatment from $D_{1}(Z_{1})$ to $D_{2}(Z_{1})$, the natural treatment change happening over time even without any change in the instrument, has an effect on the outcome that is mean independent of $Z_{2}$ conditional on $Z_{1}$ and $D_{1}$. Importantly, note that Point 1 of Assumption 9 is placebo testable, by comparing the outcome evolutions of instrument-switchers and instrument-stayers before instrument-switchers experience a change of their instrument. Finally, Point 2 of Assumption 9 requires that units’ treatment evolutions under $Z_{1}$ be mean independent of $Z_{2}$, conditional on $Z_{1}$ and $D_{1}$. Because $D_{1}$ is conditioned upon, this parallel trends condition is equivalent to a sequential exogeneity assumption (see Robins, 1986; Bojinov et al., 2021).
Point 1 of Assumption 9 is related to identifying assumptions previously proposed in the literature. de Chaisemartin (2010) and Hudson et al. (2017) also consider IV-DID estimands, in classical designs with two periods and a binary instrument that turns on for some units at period two. Both papers introduce a “reduced-form” parallel trends assumption similar to Point 1 of Assumption 9, but without noting that it imposes restrictions on effects’ heterogeneity, even in the simple designs considered by those papers.
We also make the following assumptions.
Assumption 10
(Monotonicity and strictly positive first-stage) i) For all $(z,z^{\prime})\in\mathcal{Z}_{2}^{2}$, $z\geq z^{\prime}\Rightarrow D_{2}(z)\geq D_{2}(z^{\prime})$, and ii) $E(|D_{2}(Z_{2})-D_{2}(Z_{1})|)>0$.
i) is a monotonicity assumption similar to that in Imbens and Angrist (1994). It requires that increasing the period-two instrument weakly increases the period-two treatment. This condition is plausible when the instrument is taxes and the treatment is prices, as is the case in our application. ii) requires that the instrument has a strictly positive first stage.
Assumption 11
(Bounded instrument, Lipschitz and bounded reduced-form potential outcomes and potential treatments)
  1. 1.
    $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ are bounded subsets of $\mathbb{R}$.
  2. 2.
    For all $t\in\{1,2\}$ and for all $(z,z^{\prime})\in\mathcal{Z}_{t}^{2}$, there is a random variable $\overline{Y}\geq 0$ such that $|Y_{t}(D_{t}(z))-Y_{t}(D_{t}(z^{\prime}))|\leq\overline{Y}|z-z^{\prime}|$, with $\sup_{(z_{1},z_{2})\in\text{Supp}(Z_{1},Z_{2})}E[\overline{Y}|Z_{1}=z_{1},Z_{2}=z_{2}]<\infty$.
  3. 3.
    For all $t\in\{1,2\}$ and for all $(z,z^{\prime})\in\mathcal{Z}_{t}^{2}$, there is a random variable $\overline{D}\geq 0$ such that $|D_{t}(z)-D_{t}(z^{\prime})|\leq\overline{D}|z-z^{\prime}|$, with $\sup_{(z_{1},z_{2})\in\text{Supp}(Z_{1},Z_{2})}E[\overline{D}|Z_{1}=z_{1},Z_{2}=z_{2}]<\infty$.
Assumption 11 is an adaptation of Assumption 2 to the IV setting we consider in this section.
Assumption 12
(iid sample) We observe $(Y_{i,1},Y_{i,2},D_{i,1},D_{i,2},Z_{i,1},Z_{i,2})_{1\leq i\leq n}$, that are independent and identically distributed with the same probability distribution as $(Y_{1},Y_{2},D_{1},D_{2},Z_{1},Z_{2})$.
5.2 Target parameter
In this section, our target parameter is
$\delta_{IV}:=$ $E\left(\frac{|D_{2}(Z_{2})-D_{2}(Z_{1})|}{E(|D_{2}(Z_{2})-D_{2}(Z_{1})||SC=1)}\times\frac{Y_{2}(D_{2}(Z_{2}))-Y_{2}(D_{2}(Z_{1}))}{D_{2}(Z_{2})-D_{2}(Z_{1})}\middle|SC=1\right).$
$\delta_{IV}$ is a weighted average of the slopes of compliers-switchers’ period-two potential outcome functions, from their period-two treatment under their period-one instrument, to their period-two treatment under their period-two instrument. Slopes receive a weight proportional to the absolute value of compliers-switchers’ treatment response to the instrument change. $\delta_{IV}$ is just equal to the reduced-form WAS effect of the instrument on the outcome, divided by the first-stage WAS effect of the instrument on the treatment. With a binary instrument, such that $Z_{1}=0$ and $Z_{2}\in\{0,1\}$, our IV-WAS effect coincides with that identified in Corollary 2 of Angrist et al. (2000), in a cross-sectional IV model.
We could also consider a reduced-form AS divided by a first-stage AS. The resulting target is a weighted average of the slopes $\frac{Y_{2}(D_{2}(Z_{2}))-Y_{2}(D_{2}(Z_{1}))}{D_{2}(Z_{2})-D_{2}(Z_{1})}$, with weights proportional to $\frac{D_{2}(Z_{2})-D_{2}(Z_{1})}{Z_{2}-Z_{1}}$. It seems more natural to us to weight compliers-switchers’ slopes by the absolute value of their first-stage than by the slope of their first-stage. [^6] If the first-stage effect is homogenous and linear, the weights in the IV-AS effect reduce to one, and one recovers a standard AS effect. However, linearity and homogeneity of the first-stage effect are strong assumptions.
5.3 Identification
Let $S^{I}=1\{Z_{2}-Z_{1}\neq 0\}$, $S^{I}_{+}=1\{Z_{2}-Z_{1}>0\}$, and $S^{I}_{-}=1\{Z_{2}-Z_{1}<0\}$.
Assumption 13
(Support conditions for IV-WAS identification)
  1. 1.
    $0<P(S^{I}_{+}=1)$, and $0<P(S^{I}_{+}=1|Z_{1},D_{1})$ implies that $0<P(S^{I}=0|Z_{1},D_{1})$.
  2. 2.
    $0<P(S^{I}_{-}=1)$, and $0<P(S^{I}_{-}=1|Z_{1},D_{1})$ implies that $0<P(S^{I}=0|Z_{1},D_{1})$.
Theorem 6
If Assumptions 9-11 and 13 hold,
$\delta_{IV}=$ $\frac{E\left[\text{sgn}(\Delta Z)\left(\Delta Y-E(\Delta Y|Z_{1},D_{1},S^{I}=0)\right)\right]}{E\left[\text{sgn}(\Delta Z)\left(\Delta D-E(\Delta D|Z_{1},D_{1},S^{I}=0)\right)\right]}$ (14)
$=$ $\frac{E\left[\text{sgn}(\Delta Z)\Delta Y\right]-E\left[\Delta Y\frac{P(S^{I}_{+}=1|Z_{1},D_{1})-P(S^{I}_{-}=1|Z_{1},D_{1})}{P(S^{I}=0|Z_{1},D_{1})}P(S^{I}=0)\middle|S^{I}=0\right]}{E\left[\text{sgn}(\Delta Z)\Delta D\right]-E\left[\Delta D\frac{P(S^{I}_{+}=1|Z_{1},D_{1})-P(S^{I}_{-}=1|Z_{1},D_{1})}{P(S^{I}=0|Z_{1},D_{1})}P(S^{I}=0)\middle|S^{I}=0\right]}.$ (15)
The regression-based (resp. propensity-score-based) estimand identifying $\delta_{IV}$ is just equal to the regression-based (resp. propensity-score-based) estimand identifying the reduced-form WAS effect of the instrument on the outcome controlling for $D_{1}$, divided by the regression-based (resp. propensity-score-based) estimand identifying the first-stage WAS effect controlling for $D_{1}$.
5.4 Estimation and inference
Let
$\widehat{\delta}^{r}_{IV}=$ $\frac{\frac{1}{n}\sum_{i=1}^{n}\text{sgn}(\Delta Z_{i})\left(\Delta Y_{i}-\widehat{E}(\Delta Y|Z_{1}=Z_{i,1},D_{1}=D_{i,1},S^{I}=0)\right)}{\frac{1}{n}\sum_{i=1}^{n}\text{sgn}(\Delta Z_{i})\left(\Delta D_{i}-\widehat{E}(\Delta D|Z_{1}=Z_{i,1},D_{1}=D_{i,1},S^{I}=0)\right)},$ (16)
where $\widehat{E}(\Delta Y|Z_{1},D_{1},S^{I}=0)$ and $\widehat{E}(\Delta D|Z_{1},D_{1},S^{I}=0)$ are series estimators of $E(\Delta Y|Z_{1},D_{1},S^{I}=0)$ and $E(\Delta D|Z_{1},D_{1},S^{I}=0)$ defined analogously to the series estimator in Section 3.3.
Let us define
$\widehat{\delta}^{ps}_{IV}=$ $\frac{\frac{1}{n}\sum_{i=1}^{n}\text{sgn}(\Delta Z_{i})\Delta Y_{i}-\frac{1}{n}\sum_{i:S^{I}_{i}=0}\Delta Y_{i}\frac{\widehat{P}(S^{I}_{+}=1|Z_{1}=Z_{i1},D_{1}=D_{i1})-\widehat{P}(S^{I}_{-}=1|Z_{1}=Z_{i1},D_{1}=D_{i1})}{\widehat{P}(S^{I}=0|Z_{1}=Z_{i1},D_{1}=D_{i1})}}{\frac{1}{n}\sum_{i=1}^{n}\text{sgn}(\Delta Z_{i})\Delta D_{i}-\frac{1}{n}\sum_{i:S^{I}_{i}=0}\Delta D_{i}\frac{\widehat{P}(S^{I}_{+}=1|Z_{1}=Z_{i1},D_{1}=D_{i1})-\widehat{P}(S^{I}_{-}=1|Z_{1}=Z_{i1},D_{1}=D_{i1})}{\widehat{P}(S^{I}=0|Z_{1}=Z_{i1},D_{1}=D_{i1})}},$ (17)
where $\widehat{P}(S^{I}_{+}=1|Z_{1},D_{1})$ (resp. $\widehat{P}(S^{I}_{-}=1|Z_{1},D_{1})$, $\widehat{P}(S^{I}=0|Z_{1},D_{1})$) is a series logistic regression estimator of $P(S^{I}_{+}=1|Z_{1},D_{1})$ (resp. $P(S^{I}_{-}=1|Z_{1},D_{1})$, $P(S^{I}=0|Z_{1},D_{1})$) defined analogously to the series logistic regression estimators in Section 4.3.
For any variable $X$, let
$\delta_{X}$ $=E\left[\text{sgn}(\Delta Z)\left(\Delta X-E(\Delta X|Z_{1},D_{1},S^{I}=0)\right)\right]$
$\psi_{X}$ $=\frac{1}{E(\left|\Delta Z\right|)}\left\{\left(S^{I}_{+}-S^{I}_{-}-E(S^{I}_{+}-S^{I}_{-}|Z_{1},D_{1})\frac{(1-S^{I})}{E(1-S^{I}|Z_{1},D_{1})}\right)\right.$
×$\qquad\times(\Delta X-E(\Delta X|Z_{1},D_{1},S^{I}=0))-\delta_{X}\left|\Delta Z\right|\bigg{\}}.$
Then, let
$\psi_{IV}$ $=\frac{\psi_{Y}-\delta_{IV}\psi_{D}}{\delta_{D}}.$
Under technical conditions similar to those in Assumptions 6 and 8, one can show that
$\sqrt{n}\left(\widehat{\delta}^{r}_{IV}-\delta_{IV}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{IV})),$
$\sqrt{n}\left(\widehat{\delta}^{ps}_{IV}-\delta_{IV}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{IV})).$
6 Extensions
In this section, we return to the case where the treatment, rather than an instrument, satisfies a parallel-trends condition. Combining the extensions below with the IV case is possible.
6.1 More than two time periods
In this section, we assume the representative unit is observed at $T>2$ time periods. Let $(D_{1},...,D_{T})$ denote the unit’s treatments and $\mathcal{D}_{t}=\text{Supp}(D_{t})$ for all $t\in\{1,...,T\}$. For any $t\in\{1,...,T\}$, and for any $d\in\mathcal{D}_{t}$ let $Y_{t}(d)$ denote the unit’s potential outcome at period $t$ with treatment $d$. Finally, let $Y_{t}$ denote their observed outcome at $t$. For any $t\in\{2,...,T\}$, let $S_{t}=1\{D_{t}\neq D_{t-1}\}$ be an indicator equal to 1 if the unit’s treatment switches from period $t-1$ to $t$. Let also $S_{+,t}=1\{D_{t}>D_{t-1}\}$ and $S_{-,t}=1\{D_{t}<D_{t-1}\}$. We assume that the assumptions made in the paper, rather than just holding for $t=1$ and $t=2$, actually hold for all pairs of consecutive time periods $(t-1,t)$. For instance, we replace Assumption 1 by:
Assumption 14
(Parallel trends) For all $t\geq 2$, for all $d\in\mathcal{D}_{t-1}$, $E(\Delta Y_{t}(d)|D_{t-1}=d,D_{t})=E(\Delta Y_{t}(d)|D_{t-1}=d)$.
Assumption 14 requires that $E(\Delta Y_{t}(d)|D_{t-1}=d,D_{t}=d^{\prime})$ be constant across $d^{\prime}$: groups of units with the same period-$t-1$ treatment but different period-$t$ treatments all have the same expected outcome evolution in the counterfactual where their period-$t-1$ treatment would not have changed. Importantly, note that because Assumption 14 is conditional on $D_{t-1}$, it cannot be “chained” across pairs of time periods: it requires parallel trends over pairs of consecutive time periods, not over the entire duration of the panel. To preserve space, we do not restate our other assumptions with more than two periods.
Let
$\delta_{1,t}$ $=E\left(\frac{Y_{t}(D_{t})-Y_{t}(D_{t-1})}{D_{t}-D_{t-1}}\middle|S_{t}=1\right),$
$\delta_{2,t}$ $=\frac{E\left(\text{sgn}(D_{t}-D_{t-1})(Y_{t}(D_{t})-Y_{t}(D_{t-1}))\right)}{E(|D_{t}-D_{t-1}|)}.$
Let
$\delta^{T\geq 3}_{1}$ $=\sum_{t=2}^{T}\frac{P(S_{t}=1)}{\sum_{k=2}^{T}P(S_{k}=1)}\delta_{1,t},$
$\delta^{T\geq 3}_{2}$ $=\sum_{t=2}^{T}\frac{E(|\Delta D_{t}|)}{\sum_{k=2}^{T}E(|\Delta D_{k}|)}\delta_{2,t}$
be generalizations of the AS and WAS effects to applications with more than two periods. Note that in line with the spirit of the two effects, we propose different weights to aggregate the AS and WAS across time periods. For the AS, the weights are just proportional to the proportion of switchers between $t-1$ and $t$. For the WAS, the weights are proportional to the average absolute value of the treatment switch from $t-1$ to $t$.
Theorem 7
If Assumption 14 and generalizations of Assumptions 2-5 to more than two periods hold,
$\delta^{T\geq 3}_{1}=$ $\sum_{t=2}^{T}\frac{P(S_{t}=1)}{\sum_{k=2}^{T}P(S_{k}=1)}E\left(\frac{\Delta Y_{t}-E(\Delta Y_{t}|D_{t-1},S_{t}=0)}{\Delta D_{t}}\middle|S_{t}=1\right).$
Theorem 8
If Assumption 14 and generalizations of Assumptions 2 and 7 to more than two periods hold,
$\delta^{T\geq 3}_{2}=$ $\sum_{t=2}^{T}\frac{E(|\Delta D_{t}|)}{\sum_{k=2}^{T}E(|\Delta D_{k}|)}\frac{E\left(\text{sgn}(\Delta D_{t})\left(\Delta Y_{t}-E(\Delta Y_{t}|D_{t-1},S_{t}=0)\right)\right)}{E(|\Delta D_{t}|)}$
$=$ $\sum_{t=2}^{T}\frac{E(|\Delta D_{t}|)}{\sum_{k=2}^{T}E(|\Delta D_{k}|)}\frac{E\left[\text{sgn}(\Delta D_{t})\Delta Y_{t}\right]-E\left[\Delta Y_{t}\frac{P(S_{+,t}=1|D_{t-1})-P(S_{-,t}=1|D_{t-1})}{P(S_{t}=0|D_{t-1})}P(S_{t}=0)\middle|S_{t}=0\right]}{E(|\Delta D_{t}|)}.$
Theorems 7 and 8 are straightforward generalizations of Theorems 1 and 4 to settings with more than two time periods.
Let
$\psi_{1,t}=$ $\frac{1}{E(S_{t})}\left\{\left(\frac{S_{t}}{\Delta D_{t}}-E\left(\frac{S_{t}}{\Delta D_{t}}\middle|D_{t-1}\right)\frac{(1-S_{t})}{E[1-S_{t}|D_{t-1}]}\right)[\Delta Y_{t}-E(\Delta Y_{t}|D_{t-1},S_{t}=0)]-\delta_{1,t}S_{t}\right\},$
$\psi_{2,t}=$ $\frac{1}{E(|\Delta D_{t}|)}\left\{\left(S_{+,t}-S_{-,t}-E(S_{+,t}-S_{-,t}|D_{t-1})\frac{(1-S_{t})}{E(1-S_{t}|D_{t-1})}\right)(\Delta Y_{t}-E(\Delta Y_{t}|D_{t-1},S_{t}=0))-\delta_{2,t}|\Delta D_{t}|\right\}.$
After some algebra, one can show that the influence function of the AS estimator with several periods is
$\psi^{T\geq 3}_{1}$ $:=\frac{\sum_{t=2}^{T}(P(S_{t}=1)\psi_{1,t}+(\delta_{1,t}-\delta^{T\geq 3}_{1})(S_{t}-P(S_{t}=1)))}{\sum_{t=2}^{T}P(S_{t}=1)},$ (18)
while the influence function of the WAS estimators with several periods is
$\psi^{T\geq 3}_{2}$ $:=\frac{\sum_{t=2}^{T}E(\left|\Delta D_{t}\right|)\psi_{2,t}+(\delta_{2,t}-\delta^{T\geq 3}_{2})(\left|\Delta D_{t}\right|-E(\left|\Delta D_{t}\right|))}{\sum_{t=2}^{T}E(|\Delta D_{t}|)}.$ (19)
Importantly, those influence functions allow the unit’s treatments and outcomes to be arbitrarily serially correlated.
6.2 Placebo tests
With several time periods, one can test the following condition, which is closely related to Assumption 14:
Assumption 15
(Testable parallel trends) For all $t\geq 3,t\leq T$, for all $d\in\mathcal{D}_{t-1}$, $E(\Delta Y_{t-1}(d)|D_{t-2}=D_{t-1}=d,D_{t})=E(\Delta Y_{t-1}(d)|D_{t-2}=D_{t-1}=d)$.
To test that condition, one can compute a placebo version of the estimators described in the previous subsection, replacing $\Delta Y_{t}$ by $\Delta Y_{t-1}$, and restricting the sample, for each pair of consecutive time periods $(t-1,t)$, to units whose treatment did not change between $t-2$ and $t-1$. Thus, the placebo compares the average $\Delta Y_{t-1}$ of the $t-1$-to-$t$ switchers and stayers, restricting attention to $t-2$-to-$t-1$ stayers. If one finds that from $t-2$-to-$t-1$, $t-1$-to-$t$ switchers and stayers are on parallel trends, this lends credibility to Assumption 14.
Assumption 14 can only be placebo tested among $t-2$-to-$t-1$ stayers. Then, as a robustness check one may restrict the estimation of $\delta_{1}$ and $\delta_{2}$ to $t-2$-to-$t-1$ stayers, to ensure that effects are only estimated in a subsample for which the identifying assumption can be placebo tested. The resulting estimator relies on the following identifying assumption:
$\forall t\geq 3,t\leq T,d\in\mathcal{D}_{t-1}:\leavevmode\nobreak\ E(\Delta Y_{t}(d)|D_{t-2}=D_{t-1}=d,D_{t})=E(\Delta Y_{t}(d)|D_{t-2}=D_{t-1}=d),$
the exact analogue of Assumption 15 but one period ahead.
6.3 Estimators robust to dynamic effects up to a pre-specified treatment lag.
Importantly, the robustness check in the previous section also yields an estimator robust to dynamic effects up to one treatment lag. If units’ current and first treatment lag affect their current outcome, our $t-1$-to-$t$ estimators in the subsample of $t-2$-to-$t-1$ stayers are unbiased for effects of the current treatment on the outcome under the following assumption:
$\forall t\geq 3,t\leq T,d\in\mathcal{D}_{t-1}:\leavevmode\nobreak\ E(Y_{t}(d,d)-Y_{t-1}(d,d)|D_{t-2}=D_{t-1}=d,D_{t})=E(Y_{t}(d,d)-Y_{t-1}(d,d)|D_{t-2}=D_{t-1}=d).$
Similarly, if one wants to allow for effects of the first and second treatment lags on the outcome, one just needs to restrict the estimation sample to $t-3$-to-$t-1$ stayers. However, the more robustness to dynamic effects one would like to have, the smaller the estimation sample becomes.
7 Application
Data and research questions.
We use the yearly 1966-to-2008 panel dataset of Li et al. (2014), covering 48 US states (Alaska and Hawaii are excluded). In view of the long duration of this panel, it is important to keep in mind that our estimators only assume parallel trends across pairs of consecutive years, not over the panel’s entire duration. For each state×$\times$×year cell $(i,t)$, the data contains $Z_{i,t}$, the total (state plus federal) gasoline tax in cents per gallon, $D_{i,t}$, the log tax-inclusive price of gasoline, and $Y_{i,t}$, the log gasoline consumption per adult. Our goal is to estimate the effect of gasoline taxes on gasoline consumption and prices, and to estimate the price-elasticity of gasoline consumption, using taxes as an instrument. Instead, Li et al. (2014) jointly estimate the effect of gasoline taxes and tax-exclusive prices on consumption, using a TWFE regression with two treatments. Between each pair of consecutive periods, the tax-exclusive price changes in all states, so this treatment does not have stayers and its effect cannot be estimated using the estimators proposed in this paper. Thus, our estimates cannot be compared to those of Li et al. (2014).
Switching cells, and how they compare to the entire sample.
Let $\mathcal{S}$ be the set of switching $(i,t)$ cells such that $Z_{i,t}\neq Z_{i,t-1}$ but $Z_{i^{\prime},t}=Z_{i^{\prime},t-1}$ for some $i^{\prime}$. The second condition drops from the estimation seven pairs of consecutive time periods between which the federal gasoline tax changed, thus implying that all states experienced a change of their tax. $\mathcal{S}$ includes 384 cells, so effects of taxes on gasoline prices and consumptions can be estimated for 19% of the 2,016 state×$\times$×year cells for which $Z_{i,t}-Z_{i,t-1}$ can be computed. Table 1 below compares some observable characteristics of switchers and stayers. Switchers seem slightly over-represented in the later years of the panel: $t$ is on average 2.5 years larger for switchers than for stayers, and the difference is significant. On the other hand, switchers are not more populated than stayers, and their gasoline consumption and gasoline price in 1966 are not significantly different from that of stayers. Thus, there is no strong indication that the cells in $\mathcal{S}$ are a very selected subgroup.
Table 1: Comparing switchers and stayers
Dependent Variables: $t$ Adult Population log(quantity)1966 log(price)1966
Constant 1,986.7 3,691,608.0 -0.5161 3.471
(0.2739) (577,164.0) (0.0210) (0.0054)
$\mathbf{1}\{Z_{i,t}\neq Z_{i,t-1}\}$ 2.481 39,588.0 -0.0099 0.0014
(0.7519) (320,342.1) (0.0096) (0.0029)
N 2,016 2,016 2,016 2,016
Notes: The table show the results of regressions of some dependent variables on a constant and an indicator for switching cells. The standard errors shown in parentheses are clustered at the state level.
Distribution of taxes.
As an example, the top panel of Figure 1 below shows the distribution of $Z_{g,1987}$ for 1987-to-1988 stayers, while the bottom panel shows the distribution for 1987-to-1988 switchers. The figure shows that there are many values of $Z_{g,1987}$ such that only one or two states have that value, so $Z_{g,1987}$ is close to being continuously distributed. Moreover, all switchers $g$ are such that
$\min_{g^{\prime}:Z_{g^{\prime},1988}=Z_{g^{\prime},1987}}Z_{g^{\prime},1987}\leq Z_{g,1987}\leq\max_{g^{\prime}:Z_{g^{\prime},1988}=Z_{g^{\prime},1988}}Z_{g^{\prime},1987}.$
Thus, Assumption 4 seems to hold for this pair of years. $(1987,1988)$ is not atypical. While $Z_{i,t}$ varies less across states in the first years of the panel, there are many other years where $Z_{i,t}$ is close to being continuously distributed. Similarly, almost 95% of cells in $\mathcal{S}$ are such that $\min_{g^{\prime}:Z_{i^{\prime},t}=Z_{i^{\prime},t-1}}Z_{i^{\prime},t-1}\leq Z_{i,t-1}\leq\max_{g^{\prime}:Z_{i^{\prime},t}=Z_{i^{\prime},t-1}}Z_{i^{\prime},t-1}$. Dropping the few cells that do not satisfy this condition barely changes the results presented below.
Figure 1: Gasoline tax in 1987 among 1987-to-1988 switchers and stayers
Distribution of tax changes.
Figure 2 below shows the distribution of $Z_{i,t}-Z_{i,t-1}$ for the 384 cells in $\mathcal{S}$. The majority experience an increase in their taxes, but 38 cells experience a decrease. The average value of $|Z_{i,t}-Z_{i,t-1}|$ is equal to 1.61 cents, while prior to the tax change, switchers’ average gasoline price is equal to 112 cents: our estimators leverage small changes in taxes relative to gasoline prices. Finally, $\min_{(i,t)\in\mathcal{S}}|Z_{i,t}-Z_{i,t-1}|=0.05$ : some switchers experience a very small change in their taxes.
Figure 2: Distribution of tax changes between consecutive periods
Reduced-form and first-stage AS and WAS estimates.
Table 2 below shows the AS and doubly-robust WAS estimates of the reduced-form (Panel A) and first-stage (Panel B) effects of taxes on quantities and prices. We follow results from Section 5, and estimate the reduced-form and the first-stage controlling for lagged prices $D_{t-1}$, to ensure that the resulting IV estimator is robust to heterogeneous effects over time. Reduced-form and first-stage estimators where $D_{t-1}$ is not controlled for are not very different, but controlling for $D_{t-1}$ reduces the standard error of the first-stage estimator. In Column (1), the estimators are computed using a polynomial of order 1 in $(Z_{t-1},D_{t-1})$ to estimate $E(\Delta Y_{t}|Z_{t-1},D_{t-1},S_{t}=0)$, $E(\Delta D_{t}|Z_{t-1},D_{t-1},S_{t}=0)$, and the propensity scores $P(S_{+,t}=1|Z_{t-1},D_{t-1})$, $P(S_{-,t}=1|Z_{t-1},D_{t-1})$, and $P(S_{t}=0|Z_{t-1},D_{t-1})$. In Column (2), a polynomial of order 2 is used in those estimations. 10-folds cross-validation selects a polynomial of order two for $E(\Delta D_{t}|Z_{t-1},D_{t-1},S_{t}=0)$, and a polynomial of order one for all the other conditional expectations. Thus, polynomials of order 1 and 2 are in line with those selected by cross validation. Standard errors clustered at the state level, computed following (18) and (19), are shown below the estimates, between parentheses. All estimations use 1632 (48$\times$ 35) first-difference observations: 7 periods have to be excluded as they do not have stayers. Finally, the last line of each panel shows the p-value of a test that the AS and WAS effects are equal. In Panel A Column (1), the AS estimate indicates that increasing gasoline tax by 1 cent decreases quantities consumed by 0.55 percent on average for the switchers. That effect is significant at the 5% level, but it becomes smaller and insignificant when one uses a quadratic model to estimate $E(\Delta Y_{t}|Z_{t-1},D_{t-1},S_{t}=0)$. The WAS estimates are slightly lower than, but close to, the AS estimates, and they are significant irrespective of the polynomial order used in the estimation. As predicted by Proposition 1, the standard errors of the WAS estimators are almost 3 times smaller than that of the AS estimators. Equality tests that the AS and WAS effects are equal are not rejected. In Panel B, the AS estimates of the first-stage effect are insignificant. Importantly, this implies that an IV-AS estimator of the price elasticity of gasoline consumption cannot be used: this estimator does not have a significant first stage. The WAS estimates are significant, and they indicate that if gasoline tax increases by 1 cent on average, prices increase by around 0.5 percent on average for the switchers. Again, the differences between the AS and WAS effects of taxes on prices are insignificant.
Table 2: Effects of gasoline tax on quantities consumed and prices
(1) Linear model (2) Quadratic model
Panel A: Reduced-form effect of taxes on quantities consumed
AS -0.0055 -0.0034
(0.0027) (0.0032)
WAS -0.0038 -0.0034
(0.0010) (0.0011)
Observations 1,632 1,632
P-value 0.4482 0.9974
Panel B: First-stage effect of taxes on prices
AS 0.0042 0.0047
(0.0024) (0.0025)
WAS 0.0056 0.0056
(0.0009) (0.0008)
Observations 1,632 1,632
P-value 0.4729 0.6798
Notes: All estimators in the table are computed using the data of Li et al. (2014). Panel A (resp. B) shows the AS and doubly-robust WAS estimates of the reduced-form (resp. first-stage) effect of taxes on quantities (resp. prices). All estimates control for the lag of prices. In Column (1), estimates are computed using a polynomial of order 1 in $(Z_{t-1},D_{t-1})$ to estimate $E(\Delta Y_{t}|Z_{t-1},D_{t-1},S_{t}=0)$ and the propensity scores $P(S_{+,t}=1|Z_{t-1},D_{t-1})$, $P(S_{-,t}=1|Z_{t-1},D_{t-1})$, and $P(S_{t}=0|Z_{t-1},D_{t-1})$. In Column (2), estimates are computed using a polynomial of order 2 in those estimations. Standard errors clustered at the state level, computed following (18) and (19) are shown below the estimates, between parentheses. All estimations use 1632 (48$\times$ 35) first-difference observations: 7 periods have to be excluded as they do not have stayers. Finally, the last line of each panel shows the p-value of a test that the AS and WAS effects are equal.
Placebo analysis.
Table 3 below shows placebo AS and doubly-robust WAS estimates of the reduced-form and first-stage effects. The placebo estimators are analogous to the actual estimators, but they replace $\Delta Y_{t}$ by $\Delta Y_{t-1}$, and they restrict the sample, for each pair of consecutive time periods $(t-1,t)$, to states whose taxes did not change between $t-2$ and $t-1$. The placebo WAS estimates are small and insignificant, both for quantities and prices. The placebo AS estimates are larger for quantities, but they are insignificant, and less precisely estimated. This placebo analysis shows that before switchers change their gasoline taxes, switchers’ and stayers’ consumption of gasoline and gasoline prices do not follow detectably different evolutions. As a robustness check, we reestimate the AS and WAS in the placebo subsample, to ensure that effects are estimated in a subsample for which the identifying assumption can be placebo tested, and also because in that subsample estimators remain valid if the first lag of taxes affect current gasoline prices and quantities. WAS reduced-form effects are very close to those in Table 2. WAS first-stage effects are 25 to 35% smaller, though they are still positive and highly significant.
Table 3: Placebo effects of gasoline tax on quantities consumed and prices
(1) Linear model (2) Quadratic model
Panel A: Reduced-form placebo effect of taxes on quantities consumed
AS 0.0039 0.0055
(0.0035) (0.0036)
WAS 0.0001 0.0012
(0.0017) (0.0017)
Observations 1,059 1,059
Panel B: First-stage placebo effect of taxes on prices
AS 0.0006 0.0009
(0.0056) (0.0053)
WAS 0.0014 0.0013
(0.0017) (0.0015)
Observations 1,059 1,059
Notes: The table shows the placebo AS and doubly-robust WAS estimates of the reduced-form and first-stage effects of taxes on quantities and prices. The estimators and their standard errors are computed as the actual estimators, replacing $\Delta Y_{t}$ by $\Delta Y_{t-1}$, and restricting the sample, for each pair of consecutive time periods $(t-1,t)$, to states whose taxes did not change between $t-2$ and $t-1$.
IV-WAS estimate of the price-elasticity of gasoline consumption.
Table 4 shows doubly-robust IV-WAS estimates of the price-elasticity of gasoline consumption. As the instrument’s first stage is not very strong and the sample effectively only has 48 observations, asymptotic approximations may not be reliable for inference. In line with that conjecture, we find that the bootstrap distributions of the three estimators in Table 4 are non-normal, with some outliers. Therefore, we use the percentile bootstrap for inference, clustering the bootstrap at the state level. Reassuringly, these confidence intervals have nominal coverage in simulations tailored to our application. [^7] Here is the DGP used in our simulations. We estimate TWFE regressions of $Y_{i,t}$ on state and year fixed effects and $Z_{i,t}$, and of $D_{i,t}$ on state and year fixed effects and $Z_{i,t}$. We let $\hat{\gamma}^{Y}_{i}+\hat{\lambda}^{Y}_{t}+\hat{\beta}^{Y}Z_{i,t}+\epsilon^{Y}_{i,t}$ and $\hat{\gamma}^{D}_{i}+\hat{\lambda}^{D}_{t}+\hat{\beta}^{D}Z_{i,t}+\epsilon^{D}_{i,t}$ denote the resulting regression decompositions. In each simulation, the simulated instrument is just the actual instrument, while the simulated outcomes and treatments are respectively equal to $Y^{s}_{i,t}=\hat{\gamma}^{Y}_{i}+\hat{\lambda}^{Y}_{t}+\hat{\beta}^{Y}Z_{i,t}+\epsilon^{Y,s}_{i,t},$ and $D^{s}_{i,t}=\hat{\gamma}^{D}_{i}+\hat{\lambda}^{D}_{t}+\hat{\beta}^{D}Z_{i,t}+\epsilon^{D,s}_{i,t},$ where the vector of simulated residuals $(\epsilon^{Y,s}_{g,1},...,\epsilon^{Y,s}_{g,T},\epsilon^{D,s}_{g,1},...,\epsilon^{D,s}_{g,T})$ is drawn at random and with replacement from the estimated vectors of residuals $((\epsilon^{Y}_{g^{\prime},1},...,\epsilon^{Y}_{g^{\prime},T},\epsilon^{D}_{g^{\prime},1},...,\epsilon^{D}_{g^{\prime},T}))_{g^{\prime}\in\{1,...,G\}}$. Thus, the first-stage and reduced-form effects, the correlation between the reduced-form and first-stage residuals, and the residuals’ serial correlation are the same as in the sample. The IV-WAS estimates are negative, significant, and larger than -1, though their confidence intervals contain -1. We compare those estimates to a 2SLS-TWFE estimator, computed via a 2SLS regression of $Y_{i,t}$ on $D_{i,t}$ and state and year fixed effects, using $Z_{i,t}$ as the instrument. The 2SLS-TWFE coefficient is equal to -1.0836 (bootstrap confidence interval=$[-2.1207,-0.4405]$), which is 60% larger in absolute value than the IV-WAS estimate in Column (1), and almost 80% larger than that in Column (2), though the 2SLS-TWFE coefficient does not significantly differ from the two IV-WAS estimates (P-value = 0.320 and 0.232 , respectively). Interestingly, the confidence interval of the 2SLS-TWFE coefficient is almost 80% wider than that of the IV-WAS coefficient in Column (1) and 27% wider than that in Column (2), thus showing that using a more robust estimator does not always come with a substantive precision cost.
Table 4: IV estimators of the price-elasticity of gasoline consumption
(1) Linear model (2) Quadratic model
IV-WAS -0.6773 -0.6130
[-1.2101,-0.2622] [-1.3183,-0.0004]
Observations 1,632 1,632
Notes: The table shows doubly-robust IV-WAS estimates of the price-elasticity of gasoline consumption, computed using the data of Li et al. (2014). Bootstrap confidence intervals are shown below the estimates. They are computed with 500 bootstrap replications, clustered at the state level.
8 Conclusion
We propose new difference-in-difference (DID) estimators for continuous treatments. We assume that between pairs of consecutive periods, the treatment of some units, the switchers, changes, while the treatment of other units, the stayers, does not change. We propose a parallel trends assumption on the outcome evolution of switchers and stayers with the same baseline treatment. Under that assumption, two target parameters can be estimated. Our first target is the average slope of switchers’ period-two potential outcome function, from their period-one to their period-two treatment, referred to as the AS. Our second target is a weighted average of switchers’ slopes, where switchers receive a weight proportional to the absolute value of their treatment change, referred to as the WAS. Economically, the AS and WAS serve different purposes, so neither parameter dominates the other. On the other hand, when it comes to estimation, the WAS unambiguously dominates the AS. First, it can be estimated at the parametric rate even if units can experience an arbitrarily small treatment change. Second, under some conditions, its asymptotic variance is strictly lower than that of the AS estimator. Third, unlike the AS, it is amenable to doubly-robust estimation. In our application, we use US-state-level panel data to estimate the effect of gasoline taxes on gasoline consumption. The standard error of the WAS is almost three times smaller than that of the AS, and the two estimates are close.
We also consider the instrumental-variable case, as there are instances where units experiencing/not experiencing a treatment change are unlikely to be on parallel trends, but one has at hand an instrument such that units experiencing/not experiencing an instrument change are more likely to be on parallel trends. Then, we propose widely applicable IV-DID estimators, that are robust to heterogeneous effects over time but impose some restrictions on effects’ heterogeneity across units.

### v5

@@
 Abstract
-We propose difference-in-differences estimators in designs where the treatment is continuously distributed at every period, as is often the case when one studies the effects of taxes, tariffs, or prices.
-We assume that between consecutive periods, the treatment of some units, the switchers, changes, while the treatment of other units remains constant.
-We show that under a placebo-testable parallel-trends assumption, averages of the slopes of switchers’ potential outcomes can be nonparametrically estimated.
-We generalize our estimators to the instrumental-variable case.
-We use our estimators to estimate the price-elasticity of gasoline consumption.
-Keywords: differences-in-differences, continuous treatment, two-way fixed effects regressions, heterogeneous treatment effects, panel data, policy evaluation, instrumental variable.
-JEL Codes: C21, C23
+We propose difference-in-differences (DID) estimators in designs where the treatment is continuously distributed in every period, as is often the case when one studies the effects of taxes, tariffs, or prices.
+We assume that between consecutive periods, the treatment of some units, the switchers, changes, while the treatment of other units, the stayers, remains constant.
+We show that under a parallel-trends assumption, weighted averages of the slopes of switchers’ potential outcomes are nonparametrically identified by difference-in-differences estimands comparing the outcome evolutions of switchers and stayers with the same baseline treatment.
+Controlling for the baseline treatment ensures that our estimands remain valid if the treatment’s effect changes over time.
+We highlight two possible ways of weighting switcher’s slopes, and discuss their respective advantages.
+For each weighted average of slopes, we propose a doubly-robust, nonparametric, $\sqrt{n}$-consistent, and asymptotically normal estimator.
+We generalize our results to the instrumental-variable case.
+Finally, we apply our method to estimate the price-elasticity of gasoline consumption.
+Keywords: differences-in-differences, continuous treatment, instrumental variable.
 1 Introduction
-A popular method to estimate the effect of a treatment on an outcome is to estimate a two-way fixed effects (TWFE) regression that controls for unit and time fixed effects:
+To estimate a treatment’s effect, researchers often run a two-way fixed effects (TWFE) regression:
 $Y_{i,t}=\alpha_{i}+\gamma_{t}+\beta_{TWFE}D_{i,t}+u_{i,t},$
-where $D_{i,t}$ is the treatment of unit $i$ at time $t$.
-de Chaisemartin and D’Haultfœuille (2023a) find that 26 of the 100 most cited papers published by the American Economic Review from 2015 to 2019 have estimated at least one TWFE regression.
-de Chaisemartin and D’Haultfœuille (2020), Goodman-Bacon (2021), and Borusyak et al.
-(2021) have shown that under a parallel trends assumption, TWFE regressions are not robust to heterogeneous effects: they may estimate a weighted sum of treatment effects across periods and units, with some negative weights.
-Owing to the negative weights, $\beta_{TWFE}$ could be, say, negative, even if the treatment effect is positive for every unit $\times$ period.
-Importantly, the result in de Chaisemartin and D’Haultfœuille (2020) applies to binary, discrete, and continuous treatments.
-Several alternative heterogeneity-robust difference-in-difference (DID) estimators have been proposed (see Table 2 of de Chaisemartin and D’Haultfœuille, 2023b).
-Some apply to binary and staggered treatments (see Sun and Abraham, 2021; Callaway and Sant’Anna, 2021; Borusyak et al., 2021).
-Some apply to designs where all units start with a treatment equal to 00, and then get treated with heterogeneous, potentially continuously distributed treatment intensities (see de Chaisemartin and D’Haultfœuille, 2023a; Callaway et al., 2021; de Chaisemartin and D’Haultfœuille, 2024).
-However, treatments continuously distributed at every period, including the first one in the data, are ubiquitous in applied work.
-For instance, taxes (see Li et al., 2014) or tariffs (see Fajgelbaum et al., 2020) are often continuously distributed at all periods.
-No difference-in-difference estimator robust to heterogeneous effects is available for such designs.
-Proposing such estimators is the purpose of this paper.
-We assume that we have a panel data set, whose units could be geographical locations such as counties.
-We start by considering the case where the panel has two time periods.
-From period one to two, the treatment of some units, hereafter referred to as the switchers, changes.
-On the other hand, the treatment of other units, hereafter referred to as the stayers, does not change.
-We consider two target parameters.
-The first one is the average slope of switchers’ period-two potential outcome function, from their period-one to their period-two treatment, hereafter referred to as the Average of Slopes (AS).
-Our second target is a weighted average of switchers’ slopes, where switchers receive a weight proportional to the absolute value of their treatment change, hereafter referred to as the Weighted Average of Slopes (WAS).
-We propose a novel parallel trends assumption on the outcome evolution of switchers and stayers with the same period-one treatment, in the counterfactual where switchers’ treatment would not have changed.
-Because it conditions on units’ period-one treatment, this parallel-trends assumption does not impose any restriction on effect heterogeneity.
-This parallel trends assumption is also placebo testable, by comparing switchers’ and stayers’ outcome evolutions before switchers’ treatment changes.
-We view the possibility of placebo-testing it as an important advantage of our assumption, as placebo tests are an essential step in establishing the credibility of an identifying assumption in observational studies (Imbens et al., 2001; Imbens and Xu, 2024).
-We show that under our placebo-testable parallel-trends assumption, the AS and the WAS are identified.
-This contrasts with other target parameters, like the dose-response function or the average marginal effect, which can only be identified under non-placebo-testable assumptions.
-Economically, the AS and the WAS can serve different purposes, so neither parameter dominates the other.
-Under shape restrictions on the potential outcome function, the AS can be used to infer the effect of other treatment changes than those that took place from period one to two.
-Instead, the WAS can be used to conduct a cost-benefit analysis of the treatment changes that effectively took place.
-On the other hand, when it comes to estimation, the WAS unambiguously dominates the AS.
-First, we show that it can be estimated at the standard parametric rate even if switchers can experience an arbitrarily small change of their treatment between consecutive periods.
-Second, we show that under some conditions, the asymptotic variance of the WAS estimator is strictly lower than that of the AS estimator.
-Third, unlike the AS, the WAS is amenable to doubly-robust estimation, which comes with a number of advantages.
-Then, we consider the instrumental-variable (IV) case.
-For instance, one may be interested in estimating the price-elasticity of a good’s consumption.
-If prices respond to demand shocks, the counterfactual consumption trends of units experiencing and not experiencing a price change may not be parallel.
-On the other hand, the counterfactual consumption trends of units experiencing and not experiencing a tax change may be parallel.
-Then, taxes may be used as an instrument for prices.
-In such cases, we show that the reduced-form WAS effect of the instrument on the outcome divided by the first-stage WAS effect is equal to a weighted average of switchers’ outcome-slope with respect to the treatment, where switchers with a larger first-stage effect receive more weight.
-Hereafter, we refer to this effect as the IV-WAS effect.
-The ratio of the reduced-form and first-stage AS effects is also equal to a weighted average of slopes, with arguably less natural weights, so in the IV case the WAS seems both economically and statistically preferable to the AS.
-Importantly, we show that the reduced-form parallel-trends assumption implicitly restricts treatment-effect heterogeneity.
-Such restrictions can be alleviated by controlling for groups baseline treatment in the IV specification, but the resulting estimator still restricts effects’ heterogeneity across units.
-We consider other extensions.
-First, we extend our results to applications with more than two time periods.
-Importantly, with several time periods our estimators rely on a parallel-trends assumption over consecutive periods, rather than over the entire duration of the panel.
-Second, we propose a placebo estimator of the parallel-trends assumption underlying our estimators.
-Finally, we use the yearly, 1966 to 2008 US state-level panel dataset of Li et al.
+where $D_{i,t}$ is the treatment of unit $i$ at time $t$, and $\alpha_{i}$ and $\gamma_{t}$ are unit and period fixed effects.
+de Chaisemartin and D’Haultfœuille (2024) find that 26 of the 100 most cited papers published by the American Economic Review from 2015 to 2019 estimate TWFE regressions.
+Chiu et al.
+(2024) find that 52 papers published by the top three political science journals from 2017 to 2022 estimate a TWFE regression.
+Researchers have long believed that those regressions only rely on a parallel-trends assumption, requiring that without treatment, all units would have experienced the same outcome evolutions.
+However, recent papers have shown that under that assumption alone, TWFE regressions may estimate a non-convex sum of treatment effects across units and periods (see, e.g., de Chaisemartin and D’Haultfœuille, 2020; Borusyak et al., 2024).
+Then, $\beta_{TWFE}$ could be, say, negative, even if the treatment effect is positive for every $(i,t)$.
+In reaction to this finding, several “heterogeneity-robust difference-in-difference” estimators, consistent for convex combinations of $(i,t)$-specific effects under a parallel-trends assumption, have been proposed.
+However, none can be applied to treatments continuously distributed at every period.
+Yet, such treatments are commonly found.
+In economics, taxes (see Li et al., 2014) or tariffs (see Fajgelbaum et al., 2020) are often continuously distributed at all periods.
+Similarly, Peterson (2021), Aklin and Kern (2021), and McNamara et al.
+(2022) are examples of papers in political sciences and epidemiology studying treatments continuous in all periods.
+With continuous treatments, due to the lack of an untreated group, TWFE regressions often estimate a highly non-convex combination of effects.
+For instance, with two time periods, exactly 50% of $(i,t)$-specific effects are weighted negatively (de Chaisemartin and Lei, 2021).
+Then, proposing robust estimators for such treatments is important.
+We assume that we have a two-periods panel data set.
+From period one to two, the treatment of some units, referred to as the switchers, changes, while the treatment of some other units, the stayers, does not change.
+We propose a novel parallel-trends assumption, which requires that switchers and stayers with the same period-one treatment would have experienced the same average outcome evolutions from period one to two, if switchers’ treatment had not changed.
+Our assumption has two desirable features.
+First, because it conditions on units’ period-one treatment, it does not restrict effects’ heterogeneity.
+Instead, as explained later, assuming parallel trends for switchers and stayers with different period-one treatments would essentially amount to assuming that the treatment effect is constant over time.
+Second, when another prior period of data, period zero, is available, the plausibility of our assumption can be assessed via a “pre-trends” or “placebo” test, comparing the period-zero-to-one outcome evolutions of period-one-to-two switchers and stayers.
+[^1] As explained later, the pre-trends estimator has to be computed in the subsample of units whose treatment does not change from period zero to one.
+Then, one may also restrict the main estimators to that subsample.
+Thus, one can assess if switchers and stayers are on parallel trends before switchers’ treatment changes.
+Then, we consider two target parameters.
+The first is the average slope of switchers’ period-two potential outcome function, from their period-one to their period-two treatment, referred to as the Average of Slopes (AS).
+The second is a weighted average of switchers’ slopes, where switchers receive a weight proportional to the absolute value of their treatment change, referred to as the Weighted Average of Slopes (WAS).
+The AS and WAS serve different purposes.
+Under shape restrictions on the potential outcome function, the AS can be used to identify or bound the effect of other treatment changes than those that took place.
+Instead, the WAS can be used to conduct a cost-benefit analysis of the treatment changes that took place.
+Under our parallel-trends assumption, whose plausibility can be assessed using a pre-trends test, the AS and the WAS are identified by DID estimands that compare the outcome evolution of switchers and stayers with the same period-one treatment.
+This is in contrast with other targets considered in the literature like the dose-response function or the average marginal effect, which can only be identified under assumptions whose plausibility cannot be assessed via a pre-trends test.
+Turning to estimation, the challenge with a continuous treatment is that the sample does not contain switchers and stayers with the same period-one treatment.
+In other words, the estimands identifying the AS and WAS depend on nonparametric regression functions.
+Therefore, we derive doubly-robust moment conditions identifying the AS and WAS, and we propose doubly-robust estimators.
+We show that the AS estimator converges at the parametric rate, provided switchers cannot experience arbitrarily small treatment changes.
+We then show that the WAS can always be estimated at the parametric rate.
+Under some conditions, the asymptotic variance of the WAS estimator is strictly lower than that of the AS estimator.
+We extend those results in several directions.
+First, we consider the instrumental-variable (IV) case, where one makes a parallel-trends assumption with respect to an instrument rather than the treatment.
+For instance, one may be interested in estimating the price-elasticity of a good.
+If prices respond to demand shocks, the counterfactual consumption trends of units experiencing and not experiencing a price change may not be the same, so a parallel-trends assumption with respect to prices may not hold.
+But instead, one can make a parallel-trends assumption with respect to taxes, used as an instrument for prices.
+We start by noting that parallel-trends with respect to an instrument restricts treatment-effect heterogeneity.
+Then, we show that under that assumption, the reduced-form WAS effect of the instrument on the outcome divided by the first-stage WAS effect is equal to a weighted average of switchers’ outcome-slope with respect to the treatment, where switchers with a larger first-stage effect receive more weight.
+We also consider other extensions.
+First, we consider applications with more than two time periods.
+Second, we propose a pre-trends estimator to assess the plausibility of our parallel-trends assumption.
+Third, with more than two time periods, our estimators rule out dynamic effects: they assume that units’ outcome at period $t$ only depends on their period-$t$ treatment, not on their prior treatments.
+To alleviate this concern, we propose a modified version of our estimators, robust to dynamic effects up to a pre-specified treatment lag.
+Finally, we propose estimators allowing for control variables.
+The did_multiplegt_stat Stata package, under development but already available from the SSC repository, computes some of our estimators.
+As an illustration, we use the yearly, 1966 to 2008 US state-level panel dataset of Li et al.
 (2014) to estimate the effect of gasoline taxes on gasoline consumption and prices.
 Using the WAS estimators, we find a significantly negative effect of taxes on gasoline consumption, and a significantly positive effect on prices.
-The AS estimators are close to, and not significantly different from, the WAS estimators, but they are also markedly less precise: their standard errors are almost three times larger than that of the WAS estimators.
-This shows that our theoretical result on the precision ranking of the WAS and AS estimators, which we derive under strong assumptions, still holds in a real-life application where those assumptions probably do not hold.
-The precision losses attached to using the AS have consequences.
-The AS estimator of the effect of taxes on prices is not statistically significant, so with that estimator one cannot use taxes as an instrument to estimate the price-elasticity of consumption, because the instrument does not have a first-stage.
-This contrasts with the WAS, whose first-stage t-stat is around 7.
-We compute an IV-WAS estimator of the price elasticity of gasoline consumption, and find a fairly small elasticity of -0.67, in line with previous literature (for instance, Hausman and Newey, 1995, find a long-run elasticity of -0.81).
-Our placebo estimators are small, insignificant, and fairly precisely estimated, thus suggesting that our parallel trends assumption is plausible.
-Our estimators are computed by the did_multiplegt_stat Stata package, available from the SSC repository.
-Our package allows estimators with control variables and weights, see the help file and the package’s companion paper for further details.
+The AS estimators are close to, and not significantly different from, the WAS estimators, but they are insignificant because they are markedly less precise: their standard errors are almost three times larger.
+We compute an IV estimator of the price elasticity of gasoline consumption, and find a significantly negative elasticity of -0.66, which is more negative than, but not significantly different from, that reported in recent literature (see, e.g., Coglianese et al., 2017).
+Our pre-trend estimators are small and insignificant, thus lending credibility to our parallel-trends assumption.
 Related Literature.
-Our paper builds upon several previous papers in the panel data literature.
+Our paper builds upon several papers in the panel data literature.
 Chamberlain (1982) seems to be the first paper to have proposed an estimator of the AS parameter.
 Under the assumption of no counterfactual time trend, the estimator therein is a before-after estimator.
-Then, our paper is closely related to the work of Graham and Powell (2012), who also propose DID estimators of the AS (see their Equation (21)) when the treatment is continuously distributed at every time period.
-Their estimators rely on a linear effect assumption and assume that units experience the same evolution of their treatment effect over time, a parallel-trends-on-treatment-effects assumption.
+Our paper is also closely related to the work of Graham and Powell (2012), who propose DID estimators of the AS (see their Equation (21)) when the treatment is continuously distributed at every time period.
+Their estimators rely on a linear effect assumption and assume that units experience the same evolution of their treatment effect over time.
 By contrast, our estimator of the AS does not place any restriction on treatment effects.
-But our main contribution to this literature is to introduce the WAS, and to contrast the pros and cons of the AS and WAS estimators.
-Our results are also related to Hoderlein and White (2012), who consider the average marginal effect of a continuous treatment with panel data.
-However, their target parameters and identifying assumptions are different.
-For instance, they rule out systematic changes of the outcome over time.
+But our main contribution to this literature is to introduce the WAS, and to contrast the different purposes served by the two parameters.
 With respect to the aforementioned heterogeneity-robust DID literature, we make two contributions.
-First, in the non-IV case we propose estimators that can be used even if units’ treatment varies at baseline.
-Thus we usefully complement previous literature, that has mostly focused on the case where all units have a baseline treatment equal to zero (see Sun and Abraham, 2021; Callaway and Sant’Anna, 2021; Borusyak et al., 2021; Callaway et al., 2021; de Chaisemartin and D’Haultfœuille, 2024).
-One exception predating this paper is de Chaisemartin and D’Haultfœuille (2020), who allow for a non-binary discrete treatment at baseline in their Web Appendix, and propose estimators comparing switchers and stayers with the same baseline treatment.
-However, that paper does not allow for continuously distributed treatments, and comparing switchers and stayers with the same baseline treatment is no longer feasible with a continuously distributed treatment.
-[^1] Another exception, posterior to this paper, is de Chaisemartin and D’Haultfœuille (2023a), who extend the estimators in this paper to models with dynamic effects in Section 1.10 of their Web Appendix.
-Second, in the IV case, previous IV-DID literature has only considered classical designs with two periods and binary instrument and treatment (de Chaisemartin, 2010; Hudson et al., 2017), as well as fuzzy DID designs, a special case of IV-DIDs (de Chaisemartin and D’Haultfœuille, 2018).
-Instead, this paper proposes broadly applicable IV-DID estimators, and also highlights that IV-DID estimators impose restrictions on treatment effect-heterogeneity, which can be mitigated by controlling for the baseline treatment.
-Importantly, our estimators require that there be some stayers, whose treatment does not change between consecutive periods.
-This assumption is unlikely to be met when the treatment is say, precipitations: for instance, US counties never experience the exact same precipitations over two consecutive years.
-In de Chaisemartin et al.
-(2023), we discuss the (non-trivial) extension of the results in this paper to applications without stayers.
+First, in the non-IV case we propose estimators that can be used even if units receive heterogeneous treatment doses at baseline.
+Thus we complement previous literature, that has mostly focused on the case where all units receive the same dose at baseline (see Sun and Abraham, 2021; Callaway and Sant’Anna, 2021; Liu et al., 2024; Borusyak et al., 2024; de Chaisemartin et al., 2024; Callaway et al., 2021).
+One exception predating this paper is de Chaisemartin and D’Haultfœuille (2020), who, in their web appendix, consider designs where units receive heterogeneous doses of a non-binary discrete treatment at baseline.
+With a discrete treatment, they can propose DID estimators comparing switchers and stayers with the exact same baseline treatment.
+This solution is no longer applicable with a continuously distributed treatment, hence the estimation challenge we tackle in this paper.
+Another exception, contemporaneous of this paper, is de Chaisemartin and D’Haultfœuille (2024).
+They propose estimators robust to dynamic effects up to any lag, focusing mostly on discrete treatments.
+In Section 1.10 of their web appendix, they build upon the current paper to sketch an extension of their estimators to continuous treatments.
+However, contrary to here, the estimators therein are parametric and not doubly-robust, and the estimators’ asymptotic distribution is not derived.
+Moreover, they show that with dynamic effects, under a parallel-trends assumption one can identify weighted averages of potential outcome’s slopes with respect to various treatment lags.
+Those effects are harder to interpret than the effects identified in this paper.
+In the IV case, we extend de Chaisemartin (2010) and Hudson et al.
+(2017), who consider designs with a binary instrument and treatment.
+Our paper is also the first to note that parallel-trends with respect to an instrument restricts treatment-effect heterogeneity.
 Organization of the paper.
-In Section 2, we present the set-up, introduce notation and discuss our main assumptions.
-In Section 3, we introduce the AS and discuss its identification and estimation.
-Section 4 then turns to the WAS.
-Section 5 extends our previous results to an instrumental variable set-up.
-We consider other extensions in Section 6.
-Finally, our application is developed in Section 7.
-The proofs are collected in the appendix.
+Section 2 presents the set-up, our main assumptions, and a building-block identification result.
+Section 3 considers identification and estimation of the AS.
+Section 4 turns to the WAS.
+Section 5 sketches some extensions, covered in our web appendix.
+Finally, Section 6 presents our application.
+The web appendix also collects all proofs.
 2 Set-up, assumptions, and building-block identification result
@@
 2), and let $\mathcal{D}_{1}$ (resp.
-$\mathcal{D}_{2}$) be the set of values $D_{1}$ (resp.
-$D_{2}$) can take, i.e.
-its support.
+$\mathcal{D}_{2}$) be its support.
 Let $S=1\{D_{2}\neq D_{1}\}$ be an indicator equal to 1 if the unit’s treatment changes from period one to two, i.e.
 if they are a switcher.
-For any $d\in\mathcal{D}_{1}\cup\mathcal{D}_{2}$, let $Y_{1}(d)$ and $Y_{2}(d)$ respectively denote the unit’s potential outcomes at periods 1 and 2 with treatment $d$, and let $Y_{1}$ and $Y_{2}$ denote their observed outcomes at periods 1 and 2.
+For any $d\in\mathcal{D}_{1}\cup\mathcal{D}_{2}$, let $Y_{1}(d)$ and $Y_{2}(d)$ respectively denote the unit’s potential outcomes at periods 1 and 2 with treatment $d$, and let $Y_{1}$ and $Y_{2}$ denote their observed outcomes.
 Our potential outcome notation assumes that $Y_{1}$ does not depend on units’ period-two treatment, thus ruling out anticipation effects, a commonly-made assumption in the DID literature.
 Our notation also assumes that $Y_{2}$ does not depend on units’ period-one treatment, thus ruling out dynamic effects.
-When the treatment is continuously distributed at period one, allowing for dynamic effects opens up the so-called initial-conditions problem.
-As units receive heterogeneous doses at period one, they may have experienced treatment changes before period one.
-With dynamic effects such changes may still affect their outcome over the study period, but they cannot be accounted for because they are not observed.
-Ruling out dynamic effects allows us to abstract from this thorny issue, but could yield misleading results if dynamic effects are present.
-To alleviate this concern, in Section 6.3 we propose a modified version of our estimators, robust to dynamic effects up to a pre-specified treatment lag.
+This is not of essence in the two-periods case we focus on for now.
+This is of essence when we extend our estimators to the case where $T>2$, as we will explain in Section 5.
+Therein, we propose a modified version of our estimators, which partly alleviates this concern by allowing for dynamic effects up to a pre-specified treatment lag.
 In what follows, all equalities and inequalities involving random variables are required to hold almost surely.
@@
 2.2 Assumptions
-We make the following assumptions. [^3] Throughout the paper, we implicitly assume that all potential outcomes have an expectation.
+We make the following assumptions.
 Assumption 1
-(Parallel trends) For all $d_{1}\in\mathcal{D}_{1}$, $E(\Delta Y(d_{1})|D_{1}=d_{1},D_{2})=E(\Delta Y(d_{1})|D_{1}=d_{1})$.
-Assumption 1 is a parallel trends assumption, requiring that $\Delta Y(d_{1})$ be mean independent of $D_{2}$, conditional on $D_{1}=d_{1}$.
+(Parallel trends) $\forall$ $d_{1}\in\mathcal{D}_{1}$, $E(\Delta Y(d_{1})|D_{1}=d_{1},D_{2})=E(\Delta Y(d_{1})|D_{1}=d_{1})$.
+Assumption 1 is a parallel-trends assumption, requiring that $\Delta Y(d_{1})$ be mean independent of $D_{2}$, conditional on $D_{1}=d_{1}$.
 Assumption 2
-(Bounded treatment, Lipschitz and bounded potential outcomes)
+(Bounded treatment and regularity of potential outcomes)
   1. 1.
@@
   2. 2.
-    For all $t\in\{1,2\}$ and for all $(d,d^{\prime})\in\mathcal{D}_{t}^{2}$, there is a random variable $\overline{Y}\geq 0$ such that $|Y_{t}(d)-Y_{t}(d^{\prime})|\leq\overline{Y}|d-d^{\prime}|$, with $\sup_{(d_{1},d_{2})\in\text{Supp}(D_{1},D_{2})}E[\overline{Y}|D_{1}=d_{1},D_{2}=d_{2}]<\infty$.
+    $\sup_{d_{1}\in\mathcal{D}_{1}}E[(\Delta Y(d_{1}))^{2}|D_{1}=d_{1}]<\infty$.
+  3. 3.
+    $\forall$ $t\in\{1,2\}$ and $\forall$ $(d,d^{\prime})\in\mathcal{D}_{t}^{2}$, there is a random variable $\overline{Y}\geq 0$ such that $|Y_{t}(d)-Y_{t}(d^{\prime})|\leq\overline{Y}|d-d^{\prime}|$, with $\sup_{(d_{1},d_{2})\in\text{Supp}(D_{1},D_{2})}E[\overline{Y}^{2}|D_{1}=d_{1},D_{2}=d_{2}]<\infty$.
 Assumption 2 ensures that all the expectations below are well defined.
@@
 It also requires that the potential outcome functions be Lipschitz (with a unit-specific Lipschitz constant).
-This will automatically hold if $d\mapsto Y_{2}(d)$ is differentiable with respect to $d$ and has a bounded derivative.
+This holds if $d\mapsto Y_{2}(d)$ is differentiable with respect to $d$ and has a bounded derivative.
 Finally, for estimation and inference we assume we observe an iid sample with the same distribution as $(Y_{1},Y_{2},D_{1},D_{2})$:
 Assumption 3
-(iid sample) We observe $(Y_{i,1},Y_{i,2},D_{i,1},D_{i,2})_{1\leq i\leq n}$, that are independent and identically distributed vectors with the same probability distribution as $(Y_{1},Y_{2},D_{1},D_{2})$.
+(iid sample) We observe $(Y_{i,1},Y_{i,2},D_{i,1},D_{i,2})_{1\leq i\leq n}$, which are independent and identically distributed vectors with the same probability distribution as $(Y_{1},Y_{2},D_{1},D_{2})$.
 Importantly, Assumption 3 allows for the possibility that $Y_{1}$ and $Y_{2}$ (resp. $D_{1}$ and $D_{2}$) are serially correlated, as is commonly assumed in DID studies (see Bertrand et al., 2004).
@@
 Lemma 1
-If Assumption 1 holds, then for all $(d_{1},d_{2})\in\mathcal{D}_{1}\times\mathcal{D}_{2}$ such that $d_{1}\neq d_{2}$ and $P(S|D_{1}=d_{1})<1$,
+If Assumption 1 holds, then $\forall$ $(d_{1},d_{2})\in\mathcal{D}_{1}\times\mathcal{D}_{2}$ such that $d_{1}\neq d_{2}$ and $P(S|D_{1}=d_{1})<1$,
 $\text{TE}(d_{1},d_{2}|d_{1},d_{2}):=$ $E\left(\frac{Y_{2}(d_{2})-Y_{2}(d_{1})}{d_{2}-d_{1}}\middle|D_{1}=d_{1},D_{2}=d_{2}\right)$
 $=$ $E\left(\frac{\Delta Y-E(\Delta Y|D_{1}=d_{1},S=0)}{d_{2}-d_{1}}\middle|D_{1}=d_{1},D_{2}=d_{2}\right).$
-Proof:
+Proof: we have
 $E\left(Y_{2}(d_{2})-Y_{2}(d_{1})\middle|D_{1}=d_{1},D_{2}=d_{2}\right)$
@@
 Then, a DID estimand comparing switchers’ and stayers’ outcome evolutions identifies $E\left(Y_{2}(d_{2})-Y_{2}(d_{1})\middle|D_{1}=d_{1},D_{2}=d_{2}\right)$, and we can scale that effect by $d_{2}-d_{1}$ to identify a slope rather than an unnormalized effect.
-Note that in a canonical DID design where $\mathcal{D}_{1}=0$ and $\mathcal{D}_{2}\in\{0,1\}$, Lemma 1 only applies to $(d_{1},d_{2})=(0,1)$, $\text{TE}(0,1|0,1)$ reduces to the ATT, and the estimand reduces to the canonical DID estimand comparing the outcome evolutions of treated and untreated units.
-Note also that in a design where all units are untreated at period one, $d_{1}=0$, and
+In a canonical DID design where $\mathcal{D}_{1}=\{0\}$ and $\mathcal{D}_{2}\in\{0,1\}$, Lemma 1 only applies to $(d_{1},d_{2})=(0,1)$, $\text{TE}(0,1|0,1)$ reduces to the average treatment on the treated (ATT), and the estimand reduces to the canonical DID estimand comparing the outcome evolutions of treated and untreated units.
+Also, if all units are untreated at period one ($\mathcal{D}_{1}=\{0\}$), we have
 $\text{TE}(0,d_{2}|0,d_{2})=E\left(\frac{Y_{2}(d_{2})-Y_{2}(0)}{d_{2}}\middle|D_{2}=d_{2}\right),$
-an effect closely related to the $ATT(d|d)$ effect in Callaway et al. (2021). Thus, the effects we consider are extensions of those effects to applications with a treatment continuous at all periods.
-Lemma 1 shows that under Assumption 1,
-$(d_{1},d_{2})\mapsto\text{TE}(d_{1},d_{2}|d_{1},d_{2})$
-is identified. Of course, one may be interested in other parameters, like
-$(d,d^{\prime})\mapsto\text{TE}(d,d^{\prime}):=E\left(\frac{Y_{2}(d)-Y_{2}(d^{\prime})}{d-d^{\prime}}\right),$
-a function which, unlike $(d_{1},d_{2})\mapsto\text{TE}(d_{1},d_{2}|d_{1},d_{2}),$ applies to the entire population rather than to specific subpopulation that depend on $(d_{1},d_{2})$.
-Alternatively, one could also be interested in the average marginal effect
-$E\left(Y^{\prime}_{2}(D_{2})\right).$
-What is the appeal of $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ with respect to those other parameters?
-Conditional on $D_{1}=d_{1},D_{2}=d_{2}$, $Y_{2}(d_{2})$ is observed, so estimating $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ only requires estimating $Y_{2}(d_{1})$, switchers’ counterfactual outcomes if their treatment had not changed.
-By definition, $Y_{1}(d_{1})$ is observed.
-If the data contains a third period 00 and the treatment of some units does not change from period 00 to $1$, then $Y_{0}(d_{1})$ is also observed for some switchers and stayers.
-Then, as explained in further details in Section 6.2, one can placebo-test Assumption 1, by comparing the outcome evolutions of switchers and stayers from period 00 to $1$.
-This shows that $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ is identified under a placebo-testable parallel-trends assumption.
-On the other hand, estimating $\text{TE}(d,d^{\prime})$ requires estimating, for most units, two unobserved counterfactual outcomes.
-This cannot be achieved under a placebo-testable assumption as we only observe one potential outcome at each date.
-When $D_{1}=0$, Callaway et al.
-(2021) propose a “strong parallel-trends” assumption under which the dose-response function, a parameter closely related to $\text{TE}(0,d^{\prime})$ are identified, but their “strong parallel-trends” assumption is not placebo testable.
-Similarly, estimating $E\left(\frac{Y_{2}(D_{2})-Y_{2}(d^{\prime})}{D_{2}-d^{\prime}}\right)$ requires estimating $Y_{2}(d^{\prime})$, which cannot be achieved under a placebo-testable assumption because $Y_{1}(d^{\prime})$ is not observed for all units.
-As $Y^{\prime}_{2}(D_{2})=\lim_{d^{\prime}\rightarrow D_{2}}\frac{Y_{2}(D_{2})-Y_{2}(d^{\prime})}{D_{2}-d^{\prime}}$, the same issue applies to $E\left(Y^{\prime}_{2}(D_{2})\right)$.
-Variability in $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ across values of $(d_{1},d_{2})$ conflates a dose-response relationship that may be of economic interest, and a selection bias due to the fact that units with different period one and two treatments may have heterogeneous treatment effects (Callaway et al., 2021).
-Moreover, Lemma 1 shows that estimating $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ requires estimating the values of two conditional expectations with respect to continuous variables, at points $D_{1}=d_{1},D_{2}=d_{2}$ and $D_{1}=d_{1}$.
-Unless one is willing to make parametric functional-form assumptions, the resulting estimator will converge at a slower rate than the standard $\sqrt{n}-$ parametric rate.
-For these two reasons, in this paper we focus on averages of the slopes $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$, that can be estimated non-parametrically at the standard $\sqrt{n}-$ parametric rate, and we do not focus on the function $(d_{1},d_{2})\mapsto\text{TE}(d_{1},d_{2}|d_{1},d_{2}).$
-Finally, our DID estimands compare switchers and stayers with the same period-one treatment.
+an effect closely related to the $ATT(d|d)$ effect in Callaway et al. (2021). Thus, the effects we consider are extensions of those effects to applications with a treatment continuous in all periods.
+We do not propose estimators of $(d_{1},d_{2})\mapsto\text{TE}(d_{1},d_{2}|d_{1},d_{2})$, for two reasons.
+First, variability in $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ across values of $(d_{1},d_{2})$ conflates a dose-response relationship that may be of interest, and a selection bias, typically not of interest, due to the fact that units with different period one and two treatments may have different treatment effects (Callaway et al., 2021).
+Moreover, Lemma 1 shows that estimating $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ requires estimating two bivariate nonparametric regressions.
+Unless one is willing to make parametric functional-form assumptions, the resulting estimator converges at most at the $n^{1/3}$ rate (Stone, 1982), and may therefore be very imprecise, given that DID applications often have a relatively small number of units (e.g.
+the 50 US states in our empirical application).
+For these two reasons, in the next sections our target parameters will be two averages of the slopes $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$, which can be estimated nonparametrically at the $\sqrt{n}-$ parametric rate.
+In a recent paper, posterior to ours, Haddad et al.
+(2024) consider a parallel-trends assumption similar to Assumption 1, and propose an estimator of a functional parameter similar to $(d_{1},d_{2})\mapsto\text{TE}(d_{1},d_{2}|d_{1},d_{2})$.
+Instead of averages of the slopes $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$, one may be interested in other parameters, like the average dose-slope function $(d,d^{\prime})\mapsto\text{TE}(d,d^{\prime}):=E\left((Y_{2}(d)-Y_{2}(d^{\prime}))/(d-d^{\prime})\right),$ or the average marginal effect $E\left(Y^{\prime}_{2}(D_{2})\right)$.
+However, as discussed in Section LABEL:appendixsec:alternatives of the web appendix, while averages of $\text{TE}(d_{1},d_{2}|d_{1},d_{2})$ can be identified under an assumption whose plausibility can be assessed via a pre-trends test, $\text{TE}(d,d^{\prime})$ and $E\left(Y^{\prime}_{2}(D_{2})\right)$ cannot be identified under assumptions whose plausibility can be assessed via a pre-trends test.
+2.4 Alternative parallel-trends assumption
+The DID estimands in Lemma 1 compare switchers and stayers with the same period-one treatment.
 Instead, one could propose estimands comparing switchers and stayers, without conditioning on their period-one treatment.
@@
   2. (ii)
-    For all $(d,d^{\prime})\in\mathcal{D}_{1}^{2}$, $E(\Delta Y(d))=E(\Delta Y(d^{\prime}))$.
+    $\forall$ $(d,d^{\prime})\in\mathcal{D}_{1}^{2}$, $E(\Delta Y(d))=E(\Delta Y(d^{\prime}))$.
 (i) requires that all units experience the same evolution of their potential outcome with treatment $d$, while Assumption 1 only imposes that requirement for units with the same baseline treatment.
@@
 (ii) requires that the trend affecting all potential outcomes be the same: to rationalize a DID estimand comparing a switcher going from two to five units of treatment to a stayer with treatment equal to three, $E(\Delta Y(2))$ and $E(\Delta Y(3))$ should be equal.
-Rearranging, (ii) is equivalent to
-$E(Y_{2}(d)-Y_{2}(d^{\prime}))=E(Y_{1}(d)-Y_{1}(d^{\prime})):$ (1)
-the treatment effect should be constant over time, a strong restriction on treatment effect heterogeneity.
-Assumption 1, on the other hand, does not impose any restriction on treatment effect heterogeneity, as it only restricts one potential outcome per unit.
+Rearranging, (ii) is equivalent to $E(Y_{2}(d)-Y_{2}(d^{\prime}))=E(Y_{1}(d)-Y_{1}(d^{\prime}))$, so the treatment effect should be constant over time, a strong restriction on treatment effect heterogeneity.
+In contrast, our Assumption 1 does not restrict treatment effect heterogeneity.
 3 Estimating the average of switchers’ slopes
@@
 For instance, in Fajgelbaum et al.
-(2020), one can use this strategy to derive a lower bound of the effect of increasing tariffs’ to even higher levels than those decided by the Trump administration.
-Under convexity, one can also use the AS to derive an upper bound of the effect of changing the treatment from $D_{1}$ to a lower value than $D_{1}$.
+(2020), one can use this strategy to derive a lower bound of the effect of increasing tariffs’ to even higher levels than those imposed in 2018 by the Trump administration.
+Under convexity, one can also use the AS to derive an upper bound of the effect of changing the treatment from $D_{1}$ to a lower value.
 And under concavity, one can derive an upper (resp.
@@
 (2023) for bounds of the same kind obtained under concavity or convexity.
-Importantly, the AS is identified even if those linearity or convexity/concavity conditions fail.
-But those non-placebo testable conditions are necessary to use the AS to identify or bound the effects of alternative policies.
+The AS is identified even if those linearity or convexity/concavity conditions fail, but those conditions are necessary to use the AS to identify or bound the effects of alternative policies.
 3.2 Identification
-To identify the AS, we use a DID estimand comparing switchers and stayers with the same period-one treatment.
-This requires that there be no value of the period-one treatment $D_{1}$ such that only switchers have that value, as stated formally below.
+The AS is identified under the two following assumptions on the design.
 Assumption 4
-(Support condition for AS identification) $P(S=1)>0$, $P(S=1|D_{1})<1$.
-Assumption 4 implies that $P(S=0)>0$, meaning that there are stayers whose treatment does not change.
-While we assume that $D_{1}$ and $D_{2}$ are continuous, we also assume that the treatment is persistent, and thus $\Delta D$ has a mixed distribution with a mass point at zero.
-To identify the AS, we also start by assuming that there are no quasi-stayers: the treatment of all switchers changes by at last $c$ from period one to two, for some strictly positive $c$.
+(Overlap condition for AS identification) $P(S=1)>0$ and $P(S=0|D_{1})>\zeta$ for some $\zeta>0$.
 Assumption 5
-(No quasi-stayers) $\exists c>0$: $P(|\Delta D|>c|S=1)=1$.
-We relax Assumption 5 just below.
+(No quasi-stayers) $P(|\Delta D|>\kappa|S=1)=1$ for some $\kappa>0$.
+The estimand identifying the AS compares the outcome evolutions of switchers and stayers with the same period-one treatment.
+Then, this estimand requires that there be no value of $D_{1}$ such that only switchers have that value, as assumed in Assumption 4.
+Assumption 5 requires that there are no quasi-stayers: the treatment of all switchers changes by at least $\kappa>0$.
+Define the functions:
+$\mu_{0}(d):=E[\Delta Y|S=0,D_{1}=d]\quad\text{ and }\alpha_{0}(s,d):=E\left(\left.\frac{S}{\Delta D}\right|D_{1}=d\right)\frac{1-s}{E(1-S|D_{1}=d)}.$
 Theorem 1
-If Assumptions 1-5 hold,
-$\delta_{1}$ $=E\left(\frac{\Delta Y-E(\Delta Y|D_{1},S=0)}{\Delta D}\middle|S=1\right).$
-If there are quasi-stayers, the AS is still identified.
+If Assumptions 1, 2, 4, and 5 hold, then for any functions $(s,d)\mapsto\alpha(s,d)$ and $d\mapsto\mu(d)$, respectively defined on $\{0,1\}\times\mathcal{D}_{1}$ and $\mathcal{D}_{1}$, such that $\alpha$ is bounded and $E[|\mu(D_{1})|]<\infty$:
+$\delta_{1}=$ $E\left(\frac{1}{E[S]}\left(\frac{S}{\Delta D}-\alpha(S,D_{1})\right)\left[\Delta Y-\mu_{0}(D_{1})\right]\right)$
+$=$ $E\left(\frac{1}{E[S]}\left(\frac{S}{\Delta D}-\alpha_{0}(S,D_{1})\right)\left[\Delta Y-\mu(D_{1})\right]\right).$
+It readily follows from Lemma 1 and the law of iterated expectations that $\delta_{1}$ is identified by
+$E\left[\left(\frac{\Delta Y-E[\Delta Y|S=0,D_{1}]}{\Delta D}\right)\middle|S=1\right]=E\left[\left(\frac{\Delta Y-\mu_{0}(D_{1})}{\Delta D}\right)\middle|S=1\right],$
+where the numerator of the ratio inside the expectation is a DID comparing the outcome evolutions of switchers and stayers with the same $D_{1}$. Theorem 1 implies that a related estimand,
+$E\left(\frac{1}{E[S]}\left(\frac{S}{\Delta D}-\alpha_{0}(S,D_{1})\right)\left[\Delta Y-\mu_{0}(D_{1})\right]\right),$
+also identifies $\delta_{1}$, and is doubly-robust in the functions $\mu_{0}(D_{1})$ and $\alpha_{0}(S,D_{1})$.
+[^4] We are grateful to Andres Santos for pointing out to us that the AS is identified by a doubly-robust estimand.
+If there are quasi-stayers, the AS is still identified, but estimation may be challenging.
 For any $\eta>0$, let $S_{\eta}=1\{|\Delta D|>\eta\}$ be an indicator for switchers whose treatment changes by at least $\eta$ from period one to two.
 Theorem 2
-If Assumptions 1-4 hold,
+If Assumptions 1, 2, and 4 hold,
 $\delta_{1}$ $=\lim_{\eta\downarrow 0}E\left(\frac{\Delta Y-E(\Delta Y|D_{1},S=0)}{\Delta D}\middle|S_{\eta}=1\right).$
@@
 On the other hand,
-$\Delta Y-E(\Delta Y|D_{1},S=0)$
-$=$ $Y_{2}(D_{2})-Y_{2}(D_{1})+\Delta Y(D_{1})-E(\Delta Y(D_{1})|D_{1},S=0)$
+$\Delta Y-E(\Delta Y|D_{1},S=0)$ $=Y_{2}(D_{2})-Y_{2}(D_{1})+\Delta Y(D_{1})-E(\Delta Y(D_{1})|D_{1},S=0)$
 $\approx\Delta Y(D_{1})-E(\Delta Y(D_{1})|D_{1},S=0),$
@@
 Accordingly, with quasi-stayers the AS is irregularly identified by a limiting estimand.
+Then, we conjecture that the AS cannot be estimated at the $\sqrt{n}-$rate, as in Graham and Powell (2012), and as is often the case with target parameters identified by limiting estimands.
+Therefore, we do not consider estimation of the AS with quasi stayers.
 3.3 Estimation and inference
-With no quasi-stayers, $E\left((\Delta Y-E(\Delta Y|D_{1},S=0))/\Delta D\middle|S=1\right)$ can be estimated in three steps.
-First, one estimates $E(\Delta Y|D_{1},S=0)$ using a non-parametric regression of $\Delta Y_{i}$ on $D_{i,1}$ among stayers.
-Second, for each switcher, one computes $\widehat{E}(\Delta Y|D_{1}=D_{i,1},S=0)$, its predicted outcome evolution given its baseline treatment, according to the non-parametric regression estimated among stayers.
-Third, one lets
-$\widehat{\delta}_{1}:=$ $\frac{1}{n_{s}}\sum_{i:S_{i}=1}\frac{\Delta Y_{i}-\widehat{E}(\Delta Y|D_{1}=D_{i,1},S=0)}{\Delta D_{i}},$
-where $n_{s}=\#\{i:S_{i}=1\}$.
-To estimate $E(\Delta Y|D_{1},S=0)$, we consider a series estimator based on polynomials in $D_{1}$, $(p_{k,K_{n}}(D_{1}))_{1\leq k\leq K_{n}}$. We make the following technical assumption.
+Let
+$g_{0}(d):=E[S/\Delta D|D_{1}=d]\quad\text{and }p_{0}(d):=E[1-S|D_{1}=d].$
+It follows from Theorem 1 that estimating $\delta_{1}$ requires estimating the nuisance functions $g_{0}(d)$, $p_{0}(d)$, and $\mu_{0}(d)$ in a first step, before averaging the variable
+$\left(\frac{S_{i}}{\Delta D_{i}}-\frac{\hat{g}(D_{i,1})}{\hat{p}(D_{i,1})}(1-S_{i})\right)\left(\Delta Y_{i}-\hat{\mu}(D_{i,1})\right).$
+We propose to use cross-fitting to perform this two-step estimation.
+For any set $A$, let #$\#A$ denote its cardinality.
+Let the sample be split randomly into two sub-samples, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$, and let $I_{k}=\#\mathcal{I}_{k}$ for $k=1,2$.
+We let $\hat{g}^{(2)}$, $\hat{p}^{(2)}$ and $\hat{\mu}^{(2)}$ be nonparametric estimators of $g_{0}$, $p_{0}$ and $\mu_{0}$ based on $\mathcal{I}_{2}$.
+Then the estimator $\hat{\delta}_{1,\mathsf{DR}}^{(1)}$ is computed in $\mathcal{I}_{1}$ as follows:
+$\hat{\delta}_{1,\mathsf{DR}}^{(1)}=\frac{1}{n_{S}^{(1)}}\sum_{i\in\mathcal{I}_{1}}\left(\frac{S_{i}}{\Delta D_{i}}-\frac{\hat{g}^{(2)}(D_{i,1})}{\hat{p}^{(2)}(D_{i,1})}(1-S_{i})\right)\left(\Delta Y_{i}-\hat{\mu}^{(2)}(D_{i,1})\right),$
+where $n_{S}^{(1)}:=\sum_{i\in\mathcal{I}_{1}}S_{i}$ and we use the convention that $\hat{\delta}_{1,\mathsf{DR}}^{(1)}=0$ if $n_{S}^{(1)}=0$.
+The estimator $\hat{\delta}_{1,\mathsf{DR}}^{(2)}$ is computed similarly, permuting the roles of $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$.
+Finally the cross-fitting doubly-robust estimator is computed as:
+$\hat{\delta}_{1,\mathsf{DR}}=\frac{n_{S}^{(1)}}{n_{S}^{(1)}+n_{S}^{(2)}}\hat{\delta}_{1,\mathsf{DR}}^{(1)}+\frac{n_{S}^{(2)}}{n_{S}^{(1)}+n_{S}^{(2)}}\hat{\delta}_{1,\mathsf{DR}}^{(2)}.$
+We impose the following condition:
 Assumption 6
-(Conditions for asymptotic normality of AS estimator)
-  1. 1.
-    $D_{1}$ is continuously distributed on a compact interval $I$, with $\inf_{d\in I}f_{D_{1}}(d)>0$.
-  2. 2.
-    $E[\Delta Y^{2}]<\infty$ and $d\mapsto E[\Delta Y^{2}|D_{1}=d]$ is bounded on $I$.
-  3. 3.
-    $P(S=1)>0$ and $\sup_{d\in I}P(S=1|D_{1}=d)<1$.
-  4. 4.
-    The functions $d\mapsto E[(1-S)\Delta Y|D_{1}=d]$, $d\mapsto E[S|D_{1}=d]$ and $d\mapsto E[S/\Delta D|D_{1}=d]$ are four times continuously differentiable.
-  5. 5.
-    The polynomials $d\mapsto p_{k,K_{n}}(d)$, $1\leq k\leq K_{n}$, are orthonormal on $I$ and $K_{n}^{12}/n\to+\infty$, $K_{n}^{7}/n\to 0$.
-Point 3 is a slight reinforcement of Assumption 4.
-In Point 5, $K_{n}^{12}/n\to\infty$ requires that $K_{n}$, the order of the polynomial in $D_{1}$ we use to approximate $E(\Delta Y|D_{1},S=0)$, goes to +$\infty$+ ∞ when the sample size grows, thus ensuring that the bias of our series estimator of $E(\Delta Y|D_{1},S=0)$ tends to zero.
-$K_{n}^{7}/n\to 0$ ensures that $K_{n}$ does not go to infinity too fast, thus preventing overfitting.
+$\mathcal{I}_{1}\perp\!\!\!\perp(D_{i,2},Y_{i,1},Y_{i,2})_{i=1,...,n}|(D_{i,1})_{i=1,...,n}$ and $I_{1}/n\to\pi\in(0,1)$.
+We also require that the estimators of the nonparametric nuisance functions satisfy some convergence-rate requirements.
+For any function $h(\cdot)$, $\hat{h}^{(1)}(.),\hat{h}^{(2)}(.)$ estimators of $h(.)$ based on subsample $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ respectively and $k\in\{1,2\}$, define:
+$\left\|\hat{h}-h\right\|_{2,\mathcal{I}_{k}}:=\left(\frac{1}{I_{k}}\sum_{i\in\mathcal{I}_{k}}(\hat{h}^{(3-k)}(D_{1,i})-h(D_{1,i}))^{2}\right)^{1/2},\quad\left\|\hat{h}-h\right\|_{k,\infty}:=\sup_{d\in\mathcal{D}_{1}}\left|\hat{h}^{(3-k)}(d)-h(d)\right|.$
+Assumption 7
+For $k=1,2$,
+  1. (a)
+    $\left\|\hat{g}-g_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$, $\left\|\hat{\mu}-\mu_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$, $\left\|\hat{p}-p_{0}\right\|_{k,\infty}=o_{P}(1)$.
+  2. (b)
+    $\sqrt{n}\left\|\hat{g}-g_{0}\right\|_{2,\mathcal{I}_{k}}\times\left\|\hat{\mu}-\mu_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$, $\sqrt{n}\left\|\hat{p}-p_{0}\right\|_{2,\mathcal{I}_{k}}\times\left\|\hat{\mu}-\mu_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$.
+As all nuisance functions are conditional expectations with respect to a single variable, these convergence-rate requirements are satisfied by classical nonparametric estimators.
+For instance, one can use series estimators with a polynomial order chosen by cross validation within subsample $\mathcal{I}_{k}$ to estimate $\hat{\mu}^{(k)}(d)$, $\hat{g}^{(k)}(d)$, and $\hat{p}^{(k)}(d)$.
+Results in Section 4 of Andrews (1991) imply that those estimators satisfy the rate conditions in Assumption 7, as long as the population functions are twice differentiable.
+Note also that the mean-square convergence requirements can be established under regularity conditions and appropriate choices of tuning parameters for machine learning estimators such as random forests (Biau, 2012; Wager and Athey, 2018; Syrgkanis and Zampetakis, 2020) and neural networks (Chen and White, 1999; Farrell et al., 2021).
+[^5] We rely on uniform convergence of the propensity score to ensure that the inverse propensity score is consistent.
+While uniform convergence may be too strong for machine learning estimators, this condition may be relaxed using automatic debiased machine learning, as proposed in Chernozhukov et al.
+(2024).
 Theorem 3
-If Assumptions 1-3 and 5-6 hold,
-$\sqrt{n}\left(\widehat{\delta}_{1}-\delta_{1}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{1})),$
-where
-$\psi_{1}$ $:=\frac{1}{E(S)}\left\{\left(\frac{S}{\Delta D}-E\left(\left.\frac{S}{\Delta D}\right|D_{1}\right)\frac{(1-S)}{E[1-S|D_{1}]}\right)[\Delta Y-E(\Delta Y|D_{1},S=0)]-\delta_{1}S\right\}.$
-Theorem 3 shows that without quasi-stayers, the AS can be estimated at the $\sqrt{n}-$rate, and gives an expression of its estimator’s asymptotic variance.
-With quasi-stayers, we conjecture that the AS cannot be estimated at the $\sqrt{n}-$rate.
-This conjecture is based on a result from Graham and Powell (2012).
-Though their result applies to a broader class of estimands, it implies in particular that with quasi-stayers,
-$\lim_{\eta\downarrow 0}E\left(\frac{\Delta Y-E(\Delta Y|S=0)}{\Delta D}\middle|S_{\eta}=1\right)$
-cannot be estimated at a faster rate than $n^{1/3}$. The estimand in the previous display is closely related to our estimand
-$\lim_{\eta\downarrow 0}E\left(\frac{\Delta Y-E(\Delta Y|D_{1},S=0)}{\Delta D}\middle|S_{\eta}=1\right)$
-in Theorem 2, and is equal to it if $E(\Delta Y|D_{1},S=0)=E(\Delta Y|S=0)$.
-Then, even though the assumptions in Graham and Powell (2012) differ from ours, it seems reasonable to assume that their general conclusion still applies to our set-up: here as well, owing to $\delta_{1}$’s irregular identification, this parameter can probably not be estimated at the parametric $\sqrt{n}-$rate with quasi-stayers.
-This is one of the reasons that lead us to consider, in the next section, another target parameter that can be estimated at the parametric $\sqrt{n}-$rate with quasi-stayers.
+If Assumptions 1-7 hold, $\sqrt{n}(\hat{\delta}_{1,\mathsf{DR}}-\delta_{1})\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{1}))$ where
+$\psi_{1}:=\frac{1}{E[S]}\left\{\left(\frac{S}{\Delta D}-E\left(\left.\frac{S}{\Delta D}\right|D_{1}\right)\frac{1-S}{E(1-S|D_{1})}\right)\left[\Delta Y-E(\Delta Y|S=0,D_{1})\right]-\delta_{1}S\right\}.$
 4 Estimating a weighted average of switchers’ slopes
@@
 In this section, our target parameter is
-$\delta_{2}:=$ $E\left(\frac{|D_{2}-D_{1}|}{E(|D_{2}-D_{1}||S=1)}\times\frac{Y_{2}(D_{2})-Y_{2}(D_{1})}{D_{2}-D_{1}}\middle|S=1\right)$
-$=$ $\frac{E\left(\text{sgn}(D_{2}-D_{1})(Y_{2}(D_{2})-Y_{2}(D_{1}))\middle|S=1\right)}{E(|D_{2}-D_{1}||S=1)}$
-$=$ $\frac{E\left(\text{sgn}(D_{2}-D_{1})(Y_{2}(D_{2})-Y_{2}(D_{1}))\right)}{E(|D_{2}-D_{1}|)}.$
+$\delta_{2}:=$ $E\left(\frac{|D_{2}-D_{1}|}{E(|D_{2}-D_{1}||S=1)}\times\frac{Y_{2}(D_{2})-Y_{2}(D_{1})}{D_{2}-D_{1}}\middle|S=1\right).$
 $\delta_{2}$ is a weighted average of the slopes of switchers’ potential outcome functions from their period-one to their period-two treatments, where slopes receive a weight proportional to switchers’ absolute treatment change from period one to two.
 Accordingly, we refer to $\delta_{2}$ as the Weighted Average of Slopes (WAS).
-$\delta_{2}=\delta_{1}$ if and only if
-$\text{cov}\left(\frac{Y_{2}(D_{2})-Y_{2}(D_{1})}{D_{2}-D_{1}},|D_{2}-D_{1}|\middle|S=1\right)=0:$ (3)
 The WAS and AS are equal if and only if switchers’ slopes are uncorrelated with $|D_{2}-D_{1}|$.
-Economically, the AS and WAS serve different purposes.
-As discussed above, under shape restrictions on the potential outcome function, the AS can be used to identify or bound the effect of other treatment changes than the actual change switchers experienced from period one to two.
-The WAS cannot serve that purpose, but under some assumptions, it may be used to conduct a cost-benefit analysis of the treatment changes that took place from period one to two.
+The AS and WAS serve different purposes.
+Unlike the AS, the WAS may not be used to identify or bound the effect of other treatment changes than the actual changes that took place, even under shape restrictions on the potential outcome function.
+Instead, the WAS may be used to conduct a cost-benefit analysis of the treatment changes that took place from period one to two.
 To simplify the discussion, let us assume in the remainder of this paragraph that $D_{2}\geq D_{1}$.
@@
 Finally, assume that the treatment is costly, with a cost linear in dose, uniform across units, and known to the analyst: the cost of giving $d$ units of treatment to a unit at period $t$ is $c_{t}\times d$ for some known $(c_{t})_{t\in\{1,2\}}$.
-Then, $D_{2}$ is beneficial relative to $D_{1}$ if and only if $E(Y_{2}(D_{2})-c_{2}D_{2})>E(Y_{2}(D_{1})-c_{2}D_{1})$ or, equivalently,
-$\delta_{2}>c_{2}:$
-comparing $\delta_{2}$ to $c_{2}$ is sufficient to evaluate if changing the treatment from $D_{1}$ to $D_{2}$ was beneficial.
+Then, $D_{2}$ is beneficial relative to $D_{1}$ if and only if $E(Y_{2}(D_{2})-c_{2}D_{2})>E(Y_{2}(D_{1})-c_{2}D_{1})$, namely if and only if $\delta_{2}>c_{2}:$ comparing $\delta_{2}$ to $c_{2}$ is sufficient to evaluate if changing the treatment from $D_{1}$ to $D_{2}$ was beneficial.
 4.2 Identification
-Let $S_{+}=1\{D_{2}-D_{1}>0\}$, $S_{-}=1\{D_{2}-D_{1}<0\}$ and
-$\delta_{2+}$ $:=\frac{E\left(Y_{2}(D_{2})-Y_{2}(D_{1})|S_{+}=1\right)}{E(D_{2}-D_{1}|S_{+}=1)},$
-$\delta_{2-}$ $:=\frac{E\left(Y_{2}(D_{1})-Y_{2}(D_{2})|S_{-}=1\right)}{E(D_{1}-D_{2}|S_{-}=1)}.$
-Hereafter, units with $S_{+}=1$ are referred to as “switchers up”, while units with $S_{-}=1$ are referred to as “switchers down”.
-Thus, $\delta_{2+}$ is the WAS of switchers up, and $\delta_{2-}$ is the WAS of switchers down.
-One has
-$\delta_{2}=$ $\frac{P(S_{+}=1|S=1)E(D_{2}-D_{1}|S_{+}=1)}{E(|D_{2}-D_{1}||S=1)}\delta_{2+}$
-$+$ $\frac{P(S_{-}=1|S=1)E(D_{1}-D_{2}|S_{-}=1)}{E(|D_{2}-D_{1}||S=1)}\delta_{2-}.$ (4)
-To identify $\delta_{2+}$ (resp.
-$\delta_{2-}$) we use DID estimands comparing switchers up (resp.
-switchers down) to stayers with the same period-one treatment.
-This requires that there be no value of $D_{1}$ such that some switchers up (resp.
-switchers down) have that baseline treatment while there is no stayer with the same baseline treatment, as stated in Point 1 (resp.
-2) of Assumption 7 below.
+Let $S_{+}=1\{D_{2}-D_{1}>0\}$ and $S_{-}=1\{D_{2}-D_{1}<0\}$ respectively be indicators for “switchers up” and “switchers down” and define the function:
+$\gamma_{0}(s,d):=\frac{E(S_{+}|D_{1}=d)-E(S_{-}|D_{1}=d)}{E(1-S|D_{1}=d)}(1-s).$
+Theorem 4
+If Assumptions 1, 2, and 4 hold, then for any functions $(s,d)\mapsto\gamma(s,d)$ and $d\mapsto\mu(d)$, respectively defined on $\{0,1\}\times\mathcal{D}_{1}$ and $\mathcal{D}_{1}$ and such that $\gamma$ is bounded and $E[|\mu(D_{1})|]<\infty$:
+$\delta_{2}=$ $\frac{E\left[\left(S_{+}-S_{-}-\gamma(S,D_{1})\right)(\Delta Y-\mu_{0}(D_{1}))\right]}{E[\left|\Delta D\right|]}$
+$=$ $\frac{E\left[\left(S_{+}-S_{-}-\gamma_{0}(S,D_{1})\right)(\Delta Y-\mu(D_{1}))\right]}{E[\left|\Delta D\right|]}.$
+Here are the steps leading to Theorem 4. One has
+$\delta_{2}=$ $E\left(\frac{|D_{2}-D_{1}|}{E(|D_{2}-D_{1}||S=1)}\times\frac{Y_{2}(D_{2})-Y_{2}(D_{1})}{D_{2}-D_{1}}\middle|S=1\right)$
+$=$ $\frac{E\left(\text{sgn}(D_{2}-D_{1})(Y_{2}(D_{2})-Y_{2}(D_{1}))\middle|S=1\right)}{E(|D_{2}-D_{1}||S=1)}$
+$\qquad$== $\frac{E\left((S_{+}-S_{-})(Y_{2}(D_{2})-Y_{2}(D_{1}))\right)}{E(|\Delta D|)}$
+$=$ $\frac{E\left((S_{+}-S_{-})(\Delta Y-\Delta Y(D_{1}))\right)}{E(|\Delta D|)}$
+$=$ $\frac{E\left((S_{+}-S_{-})(E(\Delta Y|D_{1},D_{2})-E(\Delta Y(D_{1})|D_{1},D_{2}))\right)}{E(|\Delta D|)}$
+$=$ $\frac{E\left((S_{+}-S_{-})(E(\Delta Y|D_{1},D_{2})-E(\Delta Y|D_{1},S=0))\right)}{E(|\Delta D|)}$
+$=$ $\frac{E\left[\left(S_{+}-S_{-}\right)(\Delta Y-E(\Delta Y|D_{1},S=0))\right]}{E[\left|\Delta D\right|]}$
+$=$ $\frac{E\left[\left(S_{+}-S_{-}\right)(\Delta Y-\mu_{0}(D_{1}))\right]}{E[\left|\Delta D\right|]}.$
+The fourth equality follows from adding and subtracting $Y_{1}(D_{1})$.
+The fifth equality follows from the law of iterated expectations.
+The sixth follows from Assumption 1 and the fact that $\Delta Y=\Delta Y(D_{1})$ for stayers.
+The seventh follows from the law of iterated expectations, and the last follows from the definition of $\mu_{0}$.
+The numerator of the estimand identifying $\delta_{2}$ in the previous display is a conditional DID comparing the outcome evolutions of switchers and stayers with the same $D_{1}$.
+Following similar steps, one can show that $\delta_{2}$ is also identified by
+$\frac{E\left[\left(S_{+}-S_{-}\right)\Delta Y\right]-E\left[\frac{E(S_{+}|D_{1}=d)-E(S_{-}|D_{1}=d)}{E(1-S|D_{1}=d)}(1-S)\Delta Y\right]}{E[\left|\Delta D\right|]}=\frac{E\left[\left(S_{+}-S_{-}-\gamma_{0}(S,D_{1})\right)\Delta Y\right]}{E[\left|\Delta D\right|]},$
+an estimand whose numerator is an unconditional DID with propensity score reweighting. Theorem 4 implies that an estimand combining those two estimands,
+$\frac{E\left[\left(S_{+}-S_{-}-\gamma_{0}(S,D_{1})\right)(\Delta Y-\mu_{0}(D_{1}))\right]}{E[\left|\Delta D\right|]},$
+also identifies $\delta_{2}$, and is doubly-robust in the functions $\mu_{0}(D_{1})$ and $\gamma_{0}(S,D_{1})$.
+4.3 Estimation and inference
+Let $h_{0}(d):=E[S_{+}-S_{-}|D_{1}=d]$. As before, split the sample into two subsamples $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$, and define:
+$\hat{\delta}_{2,\mathsf{DR}}^{(1)}:=\frac{1}{\Delta_{S}^{(1)}}\sum_{i\in\mathcal{I}_{1}}\left(S_{i+}-S_{i-}-\frac{\hat{h}^{(2)}(D_{i,1})}{\hat{p}^{(2)}(D_{i,1})}(1-S_{i})\right)(\Delta Y_{i}-\hat{\mu}^{(2)}(D_{i,1})).$
+where $\Delta_{S}^{(1)}:=\sum_{i\in\mathcal{I}_{1}}\left|\Delta D_{i}\right|$ and we use the convention that $\hat{\delta}_{2,\mathsf{DR}}^{(1)}=0$ if $\Delta_{S}^{(1)}=0$.
+The estimator $\hat{\delta}_{2,\mathsf{DR}}^{(2)}$ is constructed analogously, permuting the roles of $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$.
+Finally, let
+$\hat{\delta}_{2,\mathsf{DR}}:=\frac{\Delta_{S}^{(1)}}{\Delta_{S}^{(1)}+\Delta_{S}^{(2)}}\hat{\delta}_{2,\mathsf{DR}}^{(1)}+\frac{\Delta_{S}^{(2)}}{\Delta_{S}^{(1)}+\Delta_{S}^{(2)}}\hat{\delta}_{2,\mathsf{DR}}^{(2)}.$
+We impose the following rate conditions.
 Assumption 8
-(Support conditions for WAS identification)
-  1. 1.
-    $0<P(S_{+}=1)$, and $0<P(S_{+}=1|D_{1})$ implies that $0<P(S=0|D_{1})$.
-  2. 2.
-    $0<P(S_{-}=1)$, and $0<P(S_{-}=1|D_{1})$ implies that $0<P(S=0|D_{1})$.
+For $k=1,2$,
+  1. (a)
+    $\left\|\hat{h}-h_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$, $\left\|\hat{\mu}-\mu_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$, $\left\|\hat{p}-p_{0}\right\|_{k,\infty}=o_{P}(1)$.
+  2. (b)
+    $\sqrt{n}\left\|\hat{h}-h_{0}\right\|_{2,\mathcal{I}_{k}}\times\left\|\hat{\mu}-\mu_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$, $\sqrt{n}\left\|\hat{p}-p_{0}\right\|_{2,\mathcal{I}_{k}}\times\left\|\hat{\mu}-\mu_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$.
 Theorem 5
-  1. 1.
-    If Assumptions 1-2 and Point 1 of Assumption 7 hold,
-    $\delta_{2+}$ $=\frac{E\left(\Delta Y-E(\Delta Y|D_{1},S=0)\middle|S_{+}=1\right)}{E(\Delta D|S_{+}=1)}$ (5)
-    $=\frac{E\left(\Delta Y|S_{+}=1\right)-E\left(\Delta Y\frac{P(S_{+}=1|D_{1})}{P(S=0|D_{1})}\frac{P(S=0)}{P(S_{+}=1)}\middle|S=0\right)}{E(\Delta D|S_{+}=1)}.$ (6)
-  2. 2.
-    If Assumptions 1-2 and Point 2 of Assumption 7 hold,
-    $\delta_{2-}$ $=\frac{E\left(\Delta Y-E(\Delta Y|D_{1},S=0)\middle|S_{-}=1\right)}{E(\Delta D|S_{-}=1)}$ (7)
-    $=\frac{E\left(\Delta Y|S_{-}=1\right)-E\left(\Delta Y\frac{P(S_{-}=1|D_{1})}{P(S=0|D_{1})}\frac{P(S=0)}{P(S_{-}=1)}\middle|S=0\right)}{E(\Delta D|S_{-}=1)}.$ (8)
-  3. 3.
-    If Assumptions 1-2 and Assumption 7 hold,
-    $\delta_{2}=$ $\frac{E\left[\text{sgn}(\Delta D)\left(\Delta Y-E(\Delta Y|D_{1},S=0)\right)\right]}{E[|\Delta D|]}$ (9)
-    $=$ $\frac{E\left[\text{sgn}(\Delta D)\Delta Y\right]-E\left[\Delta Y\frac{P(S_{+}=1|D_{1})-P(S_{-}=1|D_{1})}{P(S=0|D_{1})}P(S=0)\middle|S=0\right]}{E[|\Delta D|]}.$ (10)
-Point 1 of Theorem 4 shows that $\delta_{2+}$, the WAS of switchers-up, is identified by two estimands, a regression-based and a propensity-score-based estimand.
-Point 2 of Theorem 4 shows that $\delta_{2-}$, the WAS of switchers down, is identified by two estimands similar to those identifying $\delta_{2+}$, replacing $S_{+}$ by $S_{-}$.
-Finally, if the conditions in Point 1 and 2 of Theorem 4 jointly hold, it directly follows from (4.2) that $\delta_{2}$, the WAS of all switchers, is identified by a weighted average of the estimands in Equations (5) and (7), and by a weighted average of the estimands in Equations (6) and (8).
-Those weighted averages simplify into the expressions given in Point 3 of Theorem 4.
-Point 3 of Theorem 4 also implies that $\delta_{2}$ is identified by the following doubly-robust estimand:
-$\frac{E\left[\left(S_{+}-S_{-}-\frac{P(S_{+}=1|D_{1})-P(S_{-}=1|D_{1})}{P(S=0|D_{1})}(1-S)\right)(\Delta Y-E(\Delta Y|D_{1},S=0))\right]}{E[\left|\Delta D\right|]}.$ (11)
-4.3 Estimation and inference
-The regression-based estimands identifying $\delta_{2+}$ and $\delta_{2-}$ can be estimated following almost the same steps as in Section 3.3. Specifically, let
-$\widehat{\delta}^{r}_{2+}:=$ $\frac{\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\left(\Delta Y_{i}-\widehat{E}(\Delta Y|D_{1}=D_{i,1},S=0)\right)}{\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta D_{i}}$
-$\widehat{\delta}^{r}_{2-}:=$ $\frac{\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\left(\Delta Y_{i}-\widehat{E}(\Delta Y|D_{1}=D_{i,1},S=0)\right)}{\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\Delta D_{i}},$
-where $n_{+}=\#\{i:S_{i+}=1\}$ and $n_{-}=\#\{i:S_{i-}=1\}$, and where $\widehat{E}(\Delta Y|D_{1},S=0)$ is the series estimator of $E(\Delta Y|D_{1},S=0)$ defined in Section 3.3 of the paper.
-Then, let
-$\widehat{w}_{+}=\frac{\frac{n_{+}}{n}\times\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta D_{i}}{\frac{n_{+}}{n}\times\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta D_{i}-\frac{n_{-}}{n}\times\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\Delta D_{i}},$
-and let
-$\widehat{\delta}^{r}_{2}=$ $\widehat{w}_{+}\widehat{\delta}^{r}_{2+}+(1-\widehat{w}_{+})\widehat{\delta}^{r}_{2-}$
-be the corresponding estimator of $\delta_{2}$.
-We now propose estimators of the propensity-score-based estimands identifying $\delta_{2+}$ and $\delta_{2-}$ in Equations (6) and (8).
-Let $\widehat{P}(S_{+}=1)=n_{+}/n$ (resp.
-$\widehat{P}(S_{-}=1)=n_{-}/n$, $\widehat{P}(S=0)=(n-n_{s})/n$) be an estimator of $P(S_{+}=1)$ (resp.
-$P(S_{-}=1)$, $P(S=0)$).
-Let $\widehat{P}(S_{+}=1|D_{1})$ (resp.
-$\widehat{P}(S_{-}=1|D_{1})$, $\widehat{P}(S=0|D_{1})$) be a non-parametric estimator of $P(S_{+}=1|D_{1})$ (resp.
-$P(S_{-}=1|D_{1})$, $P(S=0|D_{1})$) using a series logistic regression of $S_{i+}$ (resp.
-$S_{i-}$, $1-S_{i}$) on polynomials in $D_{1}$ $(p_{k,K_{n}}(D_{1}))_{1\leq k\leq K_{n}}$.
-We make the following technical assumption.
-Assumption 8
-(Conditions for asymptotic normality of propensity-score WAS estimator)
-  1. 1.
-    $D_{1}$ is continuously distributed on a compact interval $I$, with $\inf_{d\in I}f_{D_{1}}(d)>0$.
-  2. 2.
-    $E[\Delta Y^{2}]<\infty$ and $d\mapsto E[\Delta Y^{2}|D_{1}=d]$ is bounded on $I$
-  3. 3.
-    $10<E[S_{+}]<$ 1, $10<E[S_{-}]<$ 1, $E[S]>0$ and $\sup_{d\in I}E[S|D_{1}=d]<1$.
-  4. 4.
-    The functions $d\mapsto E[\Delta Y(1-S)|D_{1}=d]$, $d\mapsto E[S|D_{1}=d]$, $d\mapsto E[S_{+}|D_{1}=d]$ and $d\mapsto E[S_{-}|D_{1}=d]$ are four times continuously differentiable.
-  5. 5.
-    The polynomials $d\mapsto p_{k,K_{n}}(d)$, $k\leq 1\leq K_{n}$ are orthonormal on $I$ and $K_{n}=Cn^{\nu}$ where $61/10<\nu<1/$ 6.
-Let
-$\widehat{\delta}^{ps}_{2+}:=$ $\frac{\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta Y_{i}-\frac{1}{n-n_{s}}\sum_{i:S_{i}=0}\Delta Y_{i}\frac{\widehat{P}(S_{+}=1|D_{1}=D_{i1})}{\widehat{P}(S=0|D_{1}=D_{i1})}\frac{\widehat{P}(S=0)}{\widehat{P}(S_{+}=1)}}{\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta D_{i}}$
-$\widehat{\delta}^{ps}_{2-}:=$ $\frac{\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\Delta Y_{i}-\frac{1}{n-n_{s}}\sum_{i:S_{i}=0}\Delta Y_{i}\frac{\widehat{P}(S_{-}=1|D_{1}=D_{i1})}{\widehat{P}(S=0|D_{1}=D_{i1})}\frac{\widehat{P}(S=0)}{\widehat{P}(S_{-}=1)}}{\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\Delta D_{i}},$
-and let
-$\widehat{\delta}^{ps}_{2}=$ $\widehat{w}_{+}\widehat{\delta}^{ps}_{2+}+(1-\widehat{w}_{+})\widehat{\delta}^{ps}_{2-}$
-be the corresponding estimator of $\delta_{2}$. Let
-$\psi_{2+}$ $:=\frac{1}{E(\Delta DS_{+})}\left\{\left(S_{+}-E(S_{+}|D_{1})\frac{(1-S)}{E(1-S|D_{1})}\right)(\Delta Y-E(\Delta Y|D_{1},S=0))-\delta_{2+}\Delta DS_{+}\right\}$
-$\psi_{2-}$ $:=\frac{1}{E(\Delta DS_{-})}\left\{\left(S_{-}-E(S_{-}|D_{1})\frac{(1-S)}{E(1-S|D_{1})}\right)(\Delta Y-E(\Delta Y|D_{1},S=0))-\delta_{2-}\Delta DS_{-}\right\}$
-$\psi_{2}$ $:=\frac{1}{E(\left|\Delta D\right|)}\left\{\left(S_{+}-S_{-}-E(S_{+}-S_{-}|D_{1})\frac{(1-S)}{E(1-S|D_{1})}\right)\right.$
-×$\qquad\times(\Delta Y-E(\Delta Y|D_{1},S=0))-\delta_{2}\left|\Delta D\right|\bigg{\}}.$
-Theorem 5
-  1. 1.
-    If Assumptions 1-3 and 6 hold,
-    $\sqrt{n}\left((\widehat{\delta}^{r}_{2+},\widehat{\delta}^{r}_{2-})^{\prime}-(\delta_{2+},\delta_{2-})^{\prime}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V((\psi_{2+},\psi_{2-})^{\prime})).$
-    and
-    $\sqrt{n}\left(\widehat{\delta}^{r}_{2}-\delta_{2}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{2})).$
-  2. 2.
-    If Assumptions 1-3 and 8 hold,
-    $\sqrt{n}\left((\widehat{\delta}^{ps}_{2+},\widehat{\delta}^{ps}_{2-})^{\prime}-(\delta_{2+},\delta_{2-})^{\prime}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V((\psi_{2+},\psi_{2-})^{\prime})).$
-    and
-    $\sqrt{n}\left(\widehat{\delta}^{ps}_{2}-\delta_{2}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{2})).$
-Based on (11), we can also estimate $\delta_{2}$ using the following doubly-robust estimator:
-$\hat{\delta}^{dr}_{2}=\frac{\sum_{i}\left(S_{i+}-S_{i-}-\frac{\hat{P}(S_{+}=1|D_{1}=D_{1i})-P(S_{i-}=1|D_{1}=D_{1i})}{P(S_{i}=0|D_{1}=D_{1i})}(1-S_{i})\right)(\Delta Y_{i}-\hat{E}(\Delta Y_{i}|D_{1}=D_{1i},S_{i}=0))}{\sum_{i}\left|\Delta D_{i}\right|}.$
-This estimator has an important advantage.
-While our regression-based (resp.
-propensity-score-based) estimator is nominally non-parametric, in practice it requires choosing a polynomial order to estimate $E(\Delta Y|D_{1},S=0)$ (resp.
-$P(S_{+}=1|D_{1})$ and $P(S_{-}=1|D_{1})$), and the rate conditions on $K_{n}$ in Assumptions 6 (resp.
-8) do not give specific guidance on the choice of this tuning parameter.
-With the doubly-robust estimator above, one can choose this tuning parameter in a data-driven manner, using cross-validation (CV).
-Results in Section 4 of Andrews (1991) imply that a series estimator of a nonparametric regression model with a polynomial order chosen by CV is optimal: the ratio of its mean-squared error and that of an oracle estimator using the best polynomial order given the sample size converges to one.
-Then, as $D_{1}$ is a scalar variable, series estimators of $E(\Delta Y|D_{1},S=0)$, $P(S_{+}=1|D_{1})$, and $P(S_{-}=1|D_{1})$ with CV-chosen polynomial orders converge at a rate strictly faster than $n^{1/4}$, as long as one assumes that those nuisance functions are twice differentiable.
-Then, we conjecture that one can show, following arguments similar to those in Farrell (2015), that $\hat{\delta}^{dr}_{2}$ with CV-chosen polynomial orders in the estimation of the nuisance functions is $\sqrt{n}-$consistent, with asymptotic variance $V(\psi_{2})$.
+If Assumptions 1-4, 6, and 8 hold, $\sqrt{n}\left(\widehat{\delta}_{2,\mathsf{DR}}-\delta_{2}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{2}))$, where
+$\psi_{2}$ $:=\frac{1}{E(\left|\Delta D\right|)}\left\{\left(S_{+}-S_{-}-E(S_{+}-S_{-}|D_{1})\frac{(1-S)}{E(1-S|D_{1})}\right)\right.(\Delta Y-E(\Delta Y|D_{1},S=0))-\delta_{2}\left|\Delta D\right|\bigg{\}}.$
 Finally, we now show that under some assumptions, the asymptotic variance of the WAS estimator is lower than that of the AS estimator.
 Proposition 1
-If Assumption 1 holds, $(Y_{2}(D_{2})-Y_{2}(D_{1}))/(D_{2}-D_{1})=\delta$ for some real number $\delta$, $V(\Delta Y(D_{1})|D_{1},D_{2})=\sigma^{2}$ for some real number $\sigma^{2}>0$, $D_{2}\geq D_{1}$, and $\Delta D\perp\!\!\!\perp D_{1}$,
+If Assumption 1 holds, $(Y_{2}(D_{2})-Y_{2}(D_{1}))/(D_{2}-D_{1})=\delta$, $V(\Delta Y(D_{1})|D_{1},D_{2})=\sigma^{2}$ for some real number $\sigma^{2}>0$, $D_{2}\geq D_{1}$, and $\Delta D\perp\!\!\!\perp D_{1}$, then
 $V(\psi_{1})=$ $\sigma^{2}\left[\frac{E(1/(\Delta D)^{2}|S=1)}{P(S=1)}+\frac{\left(E(1/\Delta D|S=1)\right)^{2}}{P(S=0)}\right]$
@@
 with equality if and only if $V(\Delta D|S=1)=0$.
-Of course, the constant treatment effect and the homoscedasticity assumptions underlying Proposition 1 are strong, but one often has to make strong assumptions to be able to rank estimators’ variances.
-The question then is whether this ranking still holds in real-life applications, where those assumptions are unlikely to hold.
-Put differently, all models are wrong but some are useful, and the question is whether Proposition 1 is useful.
-In our empirical application, we find that the variance of $\hat{\delta}_{1}$ is indeed much larger than that of $\hat{\delta}^{dr}_{2}$, as predicted by Proposition 1.
-5 Instrumental-variable estimation
-There are instances where the parallel-trends condition in Assumption 1 is implausible, but one has at hand an instrument satisfying a similar parallel-trends condition.
-For instance, one may be interested in estimating the price-elasticity of a good’s consumption, but prices respond to supply and demand shocks, and therefore do not satisfy Assumption 1.
-On the other hand, taxes may not respond to supply and demand shocks and may satisfy a parallel-trends assumption.
-5.1 Notation and assumptions
-Let $(Z_{1},Z_{2})$ denote the instrument’s values at period one and two and $\mathcal{Z}_{t}$ be the support of $Z_{t}$.
-For any $z\in\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$, let $D_{1}(z)$ and $D_{2}(z)$ respectively denote the unit’s potential treatments at periods 1 and 2 with instrument $z$.
-Let $SC=1\{D_{2}(Z_{2})\neq D_{2}(Z_{1}),Z_{2}\neq Z_{1}\}$ be an indicator equal to 1 for switchers-compliers, namely units whose instrument changes from period one to two and whose treatment is affected by that change in the instrument.
-We replace Assumption 1 by the following assumption. [^5] Note that with our notation where potential outcomes do not depend on $z$, we also implicitly impose the usual exclusion restriction.
+The constant effect and homoscedasticity assumptions underlying Proposition 1 are admittedly strong, but ranking estimators’ variances typically requires such conditions.
+The question then is whether the ranking in Proposition 1 still holds in real-life applications, where those assumptions are likely to fail.
+In our empirical application, we find that the variance of $\widehat{\delta}_{1,\mathsf{DR}}$ is indeed much larger than that of $\widehat{\delta}_{2,\mathsf{DR}}$.
+5 Extensions
+We now briefly discuss several extensions, on which we provide more details in the web appendix.
+5.1 Instrumental-variable estimators
+There are instances where Assumption 1 is implausible, but one has an instrument that may satisfy a similar parallel-trends condition.
+This section sketches the extension of our estimators to the IV case.
+For further details, see Section LABEL:appendixsec:IV of the web appendix.
+Let $Z_{t}$ denote the instrumental variable at period $t\in\{1,2\}$ and $D_{t}(z)$ denote the unit’s potential treatment at $t$ with instrument value $z$.
+Let $SC=1\{D_{2}(Z_{2})\neq D_{2}(Z_{1})\}$ be an indicator equal to 1 for switchers-compliers, namely units whose instrument changes from period one to two and whose treatment is affected by that change in the instrument.
+We replace Assumption 1 by the following assumption. [^6] Our notation, where potential outcomes do not depend on $z$, implicitly imposes the usual exclusion restriction.
 Assumption 9
@@
 Point 1 of Assumption 9 requires that $Y_{2}(D_{2}(z))-Y_{1}(D_{1}(z))$, units’ outcome evolutions in the counterfactual where their instrument does not change from period one to two, be mean independent of $Z_{2}$, conditional on $Z_{1}$ and $D_{1}$.
-Unlike Assumption 1, this condition imposes some restrictions on treatment effect heterogeneity, and the goal of conditioning on $D_{1}$ is to minimize the stringency of those restrictions.
+This condition imposes some restrictions on treatment effect heterogeneity, and the goal of conditioning on $D_{1}$ is to alleviate those restrictions.
 To see this, note that the two following conditions are sufficient for Point 1 of Assumption 9 to hold:
@@
 (2) requires that $Y_{2}(D_{1}(z))-Y_{1}(D_{1}(z))$, units’ outcome evolutions in the counterfactual where their instrument and their treatment does not change from period one to two, be mean independent of $Z_{2}$, conditional on $Z_{1}$ and $D_{1}$.
-Thanks to the conditioning on $D_{1}$, (12) is a standard parallel trends assumption that does not impose any restriction on treatment effect heterogeneity, like Assumption 1.
-If $D_{1}$ was not conditioned upon, (12) would require parallel trends among units with different baseline treatments, which implicitly assumes homogeneous treatment effects over time, as discussed in Section 2.
+Thanks to the conditioning on $D_{1}$, (2) is a standard parallel-trends assumption that does not impose any restriction on treatment effect heterogeneity, like Assumption 1.
+If $D_{1}$ was not conditioned upon, (2) would require parallel trends among units with different baseline treatments, which implicitly assumes homogeneous treatment effects over time, as discussed in Section 2.4.
 (3), on the other hand, is a restriction on treatment effect heterogeneity across units.
-Essentially, it requires that switching the treatment from $D_{1}(Z_{1})$ to $D_{2}(Z_{1})$, the natural treatment change happening over time even without any change in the instrument, has an effect on the outcome that is mean independent of $Z_{2}$ conditional on $Z_{1}$ and $D_{1}$.
-Importantly, note that Point 1 of Assumption 9 is placebo testable, by comparing the outcome evolutions of instrument-switchers and instrument-stayers before instrument-switchers experience a change of their instrument.
-Finally, Point 2 of Assumption 9 requires that units’ treatment evolutions under $Z_{1}$ be mean independent of $Z_{2}$, conditional on $Z_{1}$ and $D_{1}$.
-Because $D_{1}$ is conditioned upon, this parallel trends condition is equivalent to a sequential exogeneity assumption (see Robins, 1986; Bojinov et al., 2021).
-Point 1 of Assumption 9 is related to identifying assumptions previously proposed in the literature.
+It requires that switching the treatment from $D_{1}(Z_{1})$ to $D_{2}(Z_{1})$, the natural treatment change happening over time even without any change in the instrument, has an effect on the outcome that is mean independent of $Z_{2}$ conditional on $Z_{1}$ and $D_{1}$.
 de Chaisemartin (2010) and Hudson et al.
 (2017) also consider IV-DID estimands, in classical designs with two periods and a binary instrument that turns on for some units at period two.
-Both papers introduce a “reduced-form” parallel trends assumption similar to Point 1 of Assumption 9, but without noting that it imposes restrictions on effects’ heterogeneity, even in the simple designs considered by those papers.
-We also make the following assumptions.
-Assumption 10
-(Monotonicity and strictly positive first-stage) i) For all $(z,z^{\prime})\in\mathcal{Z}_{2}^{2}$, $z\geq z^{\prime}\Rightarrow D_{2}(z)\geq D_{2}(z^{\prime})$, and ii) $E(|D_{2}(Z_{2})-D_{2}(Z_{1})|)>0$.
-i) is a monotonicity assumption similar to that in Imbens and Angrist (1994).
-It requires that increasing the period-two instrument weakly increases the period-two treatment.
-This condition is plausible when the instrument is taxes and the treatment is prices, as is the case in our application.
-ii) requires that the instrument has a strictly positive first stage.
-Assumption 11
-(Bounded instrument, Lipschitz and bounded reduced-form potential outcomes and potential treatments)
-  1. 1.
-    $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ are bounded subsets of $\mathbb{R}$.
-  2. 2.
-    For all $t\in\{1,2\}$ and for all $(z,z^{\prime})\in\mathcal{Z}_{t}^{2}$, there is a random variable $\overline{Y}\geq 0$ such that $|Y_{t}(D_{t}(z))-Y_{t}(D_{t}(z^{\prime}))|\leq\overline{Y}|z-z^{\prime}|$, with $\sup_{(z_{1},z_{2})\in\text{Supp}(Z_{1},Z_{2})}E[\overline{Y}|Z_{1}=z_{1},Z_{2}=z_{2}]<\infty$.
-  3. 3.
-    For all $t\in\{1,2\}$ and for all $(z,z^{\prime})\in\mathcal{Z}_{t}^{2}$, there is a random variable $\overline{D}\geq 0$ such that $|D_{t}(z)-D_{t}(z^{\prime})|\leq\overline{D}|z-z^{\prime}|$, with $\sup_{(z_{1},z_{2})\in\text{Supp}(Z_{1},Z_{2})}E[\overline{D}|Z_{1}=z_{1},Z_{2}=z_{2}]<\infty$.
-Assumption 11 is an adaptation of Assumption 2 to the IV setting we consider in this section.
-Assumption 12
-(iid sample) We observe $(Y_{i,1},Y_{i,2},D_{i,1},D_{i,2},Z_{i,1},Z_{i,2})_{1\leq i\leq n}$, that are independent and identically distributed with the same probability distribution as $(Y_{1},Y_{2},D_{1},D_{2},Z_{1},Z_{2})$.
-5.2 Target parameter
-In this section, our target parameter is
-$\delta_{IV}:=$ $E\left(\frac{|D_{2}(Z_{2})-D_{2}(Z_{1})|}{E(|D_{2}(Z_{2})-D_{2}(Z_{1})||SC=1)}\times\frac{Y_{2}(D_{2}(Z_{2}))-Y_{2}(D_{2}(Z_{1}))}{D_{2}(Z_{2})-D_{2}(Z_{1})}\middle|SC=1\right).$
-$\delta_{IV}$ is a weighted average of the slopes of compliers-switchers’ period-two potential outcome functions, from their period-two treatment under their period-one instrument, to their period-two treatment under their period-two instrument.
-Slopes receive a weight proportional to the absolute value of compliers-switchers’ treatment response to the instrument change.
-$\delta_{IV}$ is just equal to the reduced-form WAS effect of the instrument on the outcome, divided by the first-stage WAS effect of the instrument on the treatment.
-With a binary instrument, such that $Z_{1}=0$ and $Z_{2}\in\{0,1\}$, our IV-WAS effect coincides with that identified in Corollary 2 of Angrist et al.
-(2000), in a cross-sectional IV model.
-We could also consider a reduced-form AS divided by a first-stage AS.
-The resulting target is a weighted average of the slopes $\frac{Y_{2}(D_{2}(Z_{2}))-Y_{2}(D_{2}(Z_{1}))}{D_{2}(Z_{2})-D_{2}(Z_{1})}$, with weights proportional to $\frac{D_{2}(Z_{2})-D_{2}(Z_{1})}{Z_{2}-Z_{1}}$.
-It seems more natural to us to weight compliers-switchers’ slopes by the absolute value of their first-stage than by the slope of their first-stage.
-[^6] If the first-stage effect is homogenous and linear, the weights in the IV-AS effect reduce to one, and one recovers a standard AS effect.
-However, linearity and homogeneity of the first-stage effect are strong assumptions.
-5.3 Identification
-Let $S^{I}=1\{Z_{2}-Z_{1}\neq 0\}$, $S^{I}_{+}=1\{Z_{2}-Z_{1}>0\}$, and $S^{I}_{-}=1\{Z_{2}-Z_{1}<0\}$.
-Assumption 13
-(Support conditions for IV-WAS identification)
-  1. 1.
-    $0<P(S^{I}_{+}=1)$, and $0<P(S^{I}_{+}=1|Z_{1},D_{1})$ implies that $0<P(S^{I}=0|Z_{1},D_{1})$.
-  2. 2.
-    $0<P(S^{I}_{-}=1)$, and $0<P(S^{I}_{-}=1|Z_{1},D_{1})$ implies that $0<P(S^{I}=0|Z_{1},D_{1})$.
-Theorem 6
-If Assumptions 9-11 and 13 hold,
-$\delta_{IV}=$ $\frac{E\left[\text{sgn}(\Delta Z)\left(\Delta Y-E(\Delta Y|Z_{1},D_{1},S^{I}=0)\right)\right]}{E\left[\text{sgn}(\Delta Z)\left(\Delta D-E(\Delta D|Z_{1},D_{1},S^{I}=0)\right)\right]}$ (14)
-$=$ $\frac{E\left[\text{sgn}(\Delta Z)\Delta Y\right]-E\left[\Delta Y\frac{P(S^{I}_{+}=1|Z_{1},D_{1})-P(S^{I}_{-}=1|Z_{1},D_{1})}{P(S^{I}=0|Z_{1},D_{1})}P(S^{I}=0)\middle|S^{I}=0\right]}{E\left[\text{sgn}(\Delta Z)\Delta D\right]-E\left[\Delta D\frac{P(S^{I}_{+}=1|Z_{1},D_{1})-P(S^{I}_{-}=1|Z_{1},D_{1})}{P(S^{I}=0|Z_{1},D_{1})}P(S^{I}=0)\middle|S^{I}=0\right]}.$ (15)
-The regression-based (resp.
-propensity-score-based) estimand identifying $\delta_{IV}$ is just equal to the regression-based (resp.
-propensity-score-based) estimand identifying the reduced-form WAS effect of the instrument on the outcome controlling for $D_{1}$, divided by the regression-based (resp.
-propensity-score-based) estimand identifying the first-stage WAS effect controlling for $D_{1}$.
-5.4 Estimation and inference
-Let
-$\widehat{\delta}^{r}_{IV}=$ $\frac{\frac{1}{n}\sum_{i=1}^{n}\text{sgn}(\Delta Z_{i})\left(\Delta Y_{i}-\widehat{E}(\Delta Y|Z_{1}=Z_{i,1},D_{1}=D_{i,1},S^{I}=0)\right)}{\frac{1}{n}\sum_{i=1}^{n}\text{sgn}(\Delta Z_{i})\left(\Delta D_{i}-\widehat{E}(\Delta D|Z_{1}=Z_{i,1},D_{1}=D_{i,1},S^{I}=0)\right)},$ (16)
-where $\widehat{E}(\Delta Y|Z_{1},D_{1},S^{I}=0)$ and $\widehat{E}(\Delta D|Z_{1},D_{1},S^{I}=0)$ are series estimators of $E(\Delta Y|Z_{1},D_{1},S^{I}=0)$ and $E(\Delta D|Z_{1},D_{1},S^{I}=0)$ defined analogously to the series estimator in Section 3.3.
-Let us define
-$\widehat{\delta}^{ps}_{IV}=$ $\frac{\frac{1}{n}\sum_{i=1}^{n}\text{sgn}(\Delta Z_{i})\Delta Y_{i}-\frac{1}{n}\sum_{i:S^{I}_{i}=0}\Delta Y_{i}\frac{\widehat{P}(S^{I}_{+}=1|Z_{1}=Z_{i1},D_{1}=D_{i1})-\widehat{P}(S^{I}_{-}=1|Z_{1}=Z_{i1},D_{1}=D_{i1})}{\widehat{P}(S^{I}=0|Z_{1}=Z_{i1},D_{1}=D_{i1})}}{\frac{1}{n}\sum_{i=1}^{n}\text{sgn}(\Delta Z_{i})\Delta D_{i}-\frac{1}{n}\sum_{i:S^{I}_{i}=0}\Delta D_{i}\frac{\widehat{P}(S^{I}_{+}=1|Z_{1}=Z_{i1},D_{1}=D_{i1})-\widehat{P}(S^{I}_{-}=1|Z_{1}=Z_{i1},D_{1}=D_{i1})}{\widehat{P}(S^{I}=0|Z_{1}=Z_{i1},D_{1}=D_{i1})}},$ (17)
-where $\widehat{P}(S^{I}_{+}=1|Z_{1},D_{1})$ (resp.
-$\widehat{P}(S^{I}_{-}=1|Z_{1},D_{1})$, $\widehat{P}(S^{I}=0|Z_{1},D_{1})$) is a series logistic regression estimator of $P(S^{I}_{+}=1|Z_{1},D_{1})$ (resp.
-$P(S^{I}_{-}=1|Z_{1},D_{1})$, $P(S^{I}=0|Z_{1},D_{1})$) defined analogously to the series logistic regression estimators in Section 4.3.
-For any variable $X$, let
-$\delta_{X}$ $=E\left[\text{sgn}(\Delta Z)\left(\Delta X-E(\Delta X|Z_{1},D_{1},S^{I}=0)\right)\right]$
-$\psi_{X}$ $=\frac{1}{E(\left|\Delta Z\right|)}\left\{\left(S^{I}_{+}-S^{I}_{-}-E(S^{I}_{+}-S^{I}_{-}|Z_{1},D_{1})\frac{(1-S^{I})}{E(1-S^{I}|Z_{1},D_{1})}\right)\right.$
-×$\qquad\times(\Delta X-E(\Delta X|Z_{1},D_{1},S^{I}=0))-\delta_{X}\left|\Delta Z\right|\bigg{\}}.$
-Then, let
-$\psi_{IV}$ $=\frac{\psi_{Y}-\delta_{IV}\psi_{D}}{\delta_{D}}.$
-Under technical conditions similar to those in Assumptions 6 and 8, one can show that
-$\sqrt{n}\left(\widehat{\delta}^{r}_{IV}-\delta_{IV}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{IV})),$
-$\sqrt{n}\left(\widehat{\delta}^{ps}_{IV}-\delta_{IV}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{IV})).$
-6 Extensions
-In this section, we return to the case where the treatment, rather than an instrument, satisfies a parallel-trends condition. Combining the extensions below with the IV case is possible.
-6.1 More than two time periods
-In this section, we assume the representative unit is observed at $T>2$ time periods.
-Let $(D_{1},...,D_{T})$ denote the unit’s treatments and $\mathcal{D}_{t}=\text{Supp}(D_{t})$ for all $t\in\{1,...,T\}$.
-For any $t\in\{1,...,T\}$, and for any $d\in\mathcal{D}_{t}$ let $Y_{t}(d)$ denote the unit’s potential outcome at period $t$ with treatment $d$.
-Finally, let $Y_{t}$ denote their observed outcome at $t$.
-For any $t\in\{2,...,T\}$, let $S_{t}=1\{D_{t}\neq D_{t-1}\}$ be an indicator equal to 1 if the unit’s treatment switches from period $t-1$ to $t$.
-Let also $S_{+,t}=1\{D_{t}>D_{t-1}\}$ and $S_{-,t}=1\{D_{t}<D_{t-1}\}$.
-We assume that the assumptions made in the paper, rather than just holding for $t=1$ and $t=2$, actually hold for all pairs of consecutive time periods $(t-1,t)$.
-For instance, we replace Assumption 1 by:
-Assumption 14
-(Parallel trends) For all $t\geq 2$, for all $d\in\mathcal{D}_{t-1}$, $E(\Delta Y_{t}(d)|D_{t-1}=d,D_{t})=E(\Delta Y_{t}(d)|D_{t-1}=d)$.
-Assumption 14 requires that $E(\Delta Y_{t}(d)|D_{t-1}=d,D_{t}=d^{\prime})$ be constant across $d^{\prime}$: groups of units with the same period-$t-1$ treatment but different period-$t$ treatments all have the same expected outcome evolution in the counterfactual where their period-$t-1$ treatment would not have changed.
-Importantly, note that because Assumption 14 is conditional on $D_{t-1}$, it cannot be “chained” across pairs of time periods: it requires parallel trends over pairs of consecutive time periods, not over the entire duration of the panel.
-To preserve space, we do not restate our other assumptions with more than two periods.
-Let
-$\delta_{1,t}$ $=E\left(\frac{Y_{t}(D_{t})-Y_{t}(D_{t-1})}{D_{t}-D_{t-1}}\middle|S_{t}=1\right),$
-$\delta_{2,t}$ $=\frac{E\left(\text{sgn}(D_{t}-D_{t-1})(Y_{t}(D_{t})-Y_{t}(D_{t-1}))\right)}{E(|D_{t}-D_{t-1}|)}.$
-Let
-$\delta^{T\geq 3}_{1}$ $=\sum_{t=2}^{T}\frac{P(S_{t}=1)}{\sum_{k=2}^{T}P(S_{k}=1)}\delta_{1,t},$
-$\delta^{T\geq 3}_{2}$ $=\sum_{t=2}^{T}\frac{E(|\Delta D_{t}|)}{\sum_{k=2}^{T}E(|\Delta D_{k}|)}\delta_{2,t}$
-be generalizations of the AS and WAS effects to applications with more than two periods.
-Note that in line with the spirit of the two effects, we propose different weights to aggregate the AS and WAS across time periods.
-For the AS, the weights are just proportional to the proportion of switchers between $t-1$ and $t$.
-For the WAS, the weights are proportional to the average absolute value of the treatment switch from $t-1$ to $t$.
-Theorem 7
-If Assumption 14 and generalizations of Assumptions 2-5 to more than two periods hold,
-$\delta^{T\geq 3}_{1}=$ $\sum_{t=2}^{T}\frac{P(S_{t}=1)}{\sum_{k=2}^{T}P(S_{k}=1)}E\left(\frac{\Delta Y_{t}-E(\Delta Y_{t}|D_{t-1},S_{t}=0)}{\Delta D_{t}}\middle|S_{t}=1\right).$
-Theorem 8
-If Assumption 14 and generalizations of Assumptions 2 and 7 to more than two periods hold,
-$\delta^{T\geq 3}_{2}=$ $\sum_{t=2}^{T}\frac{E(|\Delta D_{t}|)}{\sum_{k=2}^{T}E(|\Delta D_{k}|)}\frac{E\left(\text{sgn}(\Delta D_{t})\left(\Delta Y_{t}-E(\Delta Y_{t}|D_{t-1},S_{t}=0)\right)\right)}{E(|\Delta D_{t}|)}$
-$=$ $\sum_{t=2}^{T}\frac{E(|\Delta D_{t}|)}{\sum_{k=2}^{T}E(|\Delta D_{k}|)}\frac{E\left[\text{sgn}(\Delta D_{t})\Delta Y_{t}\right]-E\left[\Delta Y_{t}\frac{P(S_{+,t}=1|D_{t-1})-P(S_{-,t}=1|D_{t-1})}{P(S_{t}=0|D_{t-1})}P(S_{t}=0)\middle|S_{t}=0\right]}{E(|\Delta D_{t}|)}.$
-Theorems 7 and 8 are straightforward generalizations of Theorems 1 and 4 to settings with more than two time periods.
-Let
-$\psi_{1,t}=$ $\frac{1}{E(S_{t})}\left\{\left(\frac{S_{t}}{\Delta D_{t}}-E\left(\frac{S_{t}}{\Delta D_{t}}\middle|D_{t-1}\right)\frac{(1-S_{t})}{E[1-S_{t}|D_{t-1}]}\right)[\Delta Y_{t}-E(\Delta Y_{t}|D_{t-1},S_{t}=0)]-\delta_{1,t}S_{t}\right\},$
-$\psi_{2,t}=$ $\frac{1}{E(|\Delta D_{t}|)}\left\{\left(S_{+,t}-S_{-,t}-E(S_{+,t}-S_{-,t}|D_{t-1})\frac{(1-S_{t})}{E(1-S_{t}|D_{t-1})}\right)(\Delta Y_{t}-E(\Delta Y_{t}|D_{t-1},S_{t}=0))-\delta_{2,t}|\Delta D_{t}|\right\}.$
-After some algebra, one can show that the influence function of the AS estimator with several periods is
-$\psi^{T\geq 3}_{1}$ $:=\frac{\sum_{t=2}^{T}(P(S_{t}=1)\psi_{1,t}+(\delta_{1,t}-\delta^{T\geq 3}_{1})(S_{t}-P(S_{t}=1)))}{\sum_{t=2}^{T}P(S_{t}=1)},$ (18)
-while the influence function of the WAS estimators with several periods is
-$\psi^{T\geq 3}_{2}$ $:=\frac{\sum_{t=2}^{T}E(\left|\Delta D_{t}\right|)\psi_{2,t}+(\delta_{2,t}-\delta^{T\geq 3}_{2})(\left|\Delta D_{t}\right|-E(\left|\Delta D_{t}\right|))}{\sum_{t=2}^{T}E(|\Delta D_{t}|)}.$ (19)
-Importantly, those influence functions allow the unit’s treatments and outcomes to be arbitrarily serially correlated.
-6.2 Placebo tests
-With several time periods, one can test the following condition, which is closely related to Assumption 14:
-Assumption 15
-(Testable parallel trends) For all $t\geq 3,t\leq T$, for all $d\in\mathcal{D}_{t-1}$, $E(\Delta Y_{t-1}(d)|D_{t-2}=D_{t-1}=d,D_{t})=E(\Delta Y_{t-1}(d)|D_{t-2}=D_{t-1}=d)$.
-To test that condition, one can compute a placebo version of the estimators described in the previous subsection, replacing $\Delta Y_{t}$ by $\Delta Y_{t-1}$, and restricting the sample, for each pair of consecutive time periods $(t-1,t)$, to units whose treatment did not change between $t-2$ and $t-1$.
-Thus, the placebo compares the average $\Delta Y_{t-1}$ of the $t-1$-to-$t$ switchers and stayers, restricting attention to $t-2$-to-$t-1$ stayers.
-If one finds that from $t-2$-to-$t-1$, $t-1$-to-$t$ switchers and stayers are on parallel trends, this lends credibility to Assumption 14.
-Assumption 14 can only be placebo tested among $t-2$-to-$t-1$ stayers.
-Then, as a robustness check one may restrict the estimation of $\delta_{1}$ and $\delta_{2}$ to $t-2$-to-$t-1$ stayers, to ensure that effects are only estimated in a subsample for which the identifying assumption can be placebo tested.
-The resulting estimator relies on the following identifying assumption:
-$\forall t\geq 3,t\leq T,d\in\mathcal{D}_{t-1}:\leavevmode\nobreak\ E(\Delta Y_{t}(d)|D_{t-2}=D_{t-1}=d,D_{t})=E(\Delta Y_{t}(d)|D_{t-2}=D_{t-1}=d),$
-the exact analogue of Assumption 15 but one period ahead.
-6.3 Estimators robust to dynamic effects up to a pre-specified treatment lag.
-Importantly, the robustness check in the previous section also yields an estimator robust to dynamic effects up to one treatment lag.
-If units’ current and first treatment lag affect their current outcome, our $t-1$-to-$t$ estimators in the subsample of $t-2$-to-$t-1$ stayers are unbiased for effects of the current treatment on the outcome under the following assumption:
-$\forall t\geq 3,t\leq T,d\in\mathcal{D}_{t-1}:\leavevmode\nobreak\ E(Y_{t}(d,d)-Y_{t-1}(d,d)|D_{t-2}=D_{t-1}=d,D_{t})=E(Y_{t}(d,d)-Y_{t-1}(d,d)|D_{t-2}=D_{t-1}=d).$
+Both papers introduce a “reduced-form” parallel-trends assumption similar to Point 1 of Assumption 9, but without noting that it imposes restrictions on effects’ heterogeneity, even in the simple designs considered by those papers.
+[^7] The realization that the “reduced-form” parallel-trends assumption in de Chaisemartin (2010) and Hudson et al.
+(2017) actually restricts effect heterogeneity explains the apparent discrepancy between those two papers and Blundell and Dias (2009) and de Chaisemartin and D’Haultfœuille (2018), two papers that show that IV-DIDs restrict treatment-effect heterogeneity under the standard parallel-trends assumption on the untreated outcome.
+Turning to Point 2 of Assumption 9, this condition requires that units’ treatment evolutions under $Z_{1}$ be mean independent of $Z_{2}$, conditional on $Z_{1}$ and $D_{1}$.
+Because $D_{1}$ is conditioned upon, this parallel trends condition is equivalent to a sequential exogeneity assumption (see Robins, 1986; Bojinov et al., 2021).
+Let us consider the parameter
+$\delta_{\mathsf{IV}}:=$ $E\left(\frac{|D_{2}(Z_{2})-D_{2}(Z_{1})|}{E(|D_{2}(Z_{2})-D_{2}(Z_{1})||SC=1)}\times\frac{Y_{2}(D_{2}(Z_{2}))-Y_{2}(D_{2}(Z_{1}))}{D_{2}(Z_{2})-D_{2}(Z_{1})}\middle|SC=1\right).$
+$\delta_{\mathsf{IV}}$ is a weighted average of the slopes of compliers-switchers’ period-two potential outcome functions, from their period-two treatment under their period-one instrument, to their period- two treatment under their period-two instrument.
+We show in Section LABEL:appendixsec:IV of the web appendix that under Assumption 9 and a monotonicity condition as in Imbens and Angrist (1994),
+$\delta_{\mathsf{IV}}=\frac{E\left[\left(S^{I}_{+}-S^{I}_{-}-\gamma^{I}_{0}(S^{I},Z_{1},D_{1})\right)(\Delta Y-\mu^{Y}_{0}(Z_{1},D_{1}))\right]}{E\left[\left(S^{I}_{+}-S^{I}_{-}-\gamma^{I}_{0}(S^{I},Z_{1},D_{1})\right)(\Delta D-\mu^{D}_{0}(Z_{1},D_{1}))\right]},$
+where $S^{I}=1\{Z_{2}-Z_{1}\neq 0\}$, $S^{I}_{+}=1\{Z_{2}-Z_{1}>0\}$, $S^{I}_{-}=1\{Z_{2}-Z_{1}<0\}$ and
+$\mu^{U}_{0}(z,d):=E[\Delta U|S^{I}=0,Z_{1}=z,D_{1}=d],\quad U\in\{D,Y\},$
+$\gamma^{I}_{0}(s,z,d):=\frac{E(S^{I}_{+}|Z_{1}=z,D_{1}=d)-E(S^{I}_{-}|Z_{1}=z,D_{1}=d)}{E(1-S^{I}|Z_{1}=z,D_{1}=d)}(1-s).$
+The estimand identifying $\delta_{\mathsf{IV}}$ is equal to the estimand identifying the reduced-form WAS effect of the instrument on the outcome controlling for $D_{1}$, divided by the estimand identifying the first-stage WAS effect of the instrument on the treatment controlling for $D_{1}$.
+5.2 Estimators with more than two time periods
+In this section and the next, we return to the case where the treatment, rather than an instrument, satisfies a parallel-trends condition.
+This section sketches the extension of our estimators to datasets with three periods or more.
+For further details, see Section LABEL:appendixsec:Tgeq3 of the web appendix.
+With $T>2$ time periods, we propose a simple extension of our estimators.
+Let $\widehat{\delta}_{1,t,\mathsf{DR}}$ and $\widehat{\delta}_{2,t,\mathsf{DR}}$ be defined exactly as $\hat{\delta}_{1,\mathsf{DR}}$ and $\hat{\delta}_{2,\mathsf{DR}}$ but on a pair of consecutive time periods $(t-1,t)$.
+Then, we consider a weighted average across $t$ of $\widehat{\delta}_{1,t,\mathsf{DR}}$, with weights proportional to the proportion of switchers from $t-1$ to $t$, thus yielding an AS estimator with more than two periods, denoted $\widehat{\delta}^{T\geq 3}_{1}$.
+We also consider a weighted average across $t$ of $\widehat{\delta}_{2,t,\mathsf{DR}}$, with weights proportional to the average absolute treatment change from $t-1$ to $t$, thus yielding a WAS estimator with more than two periods, denoted $\widehat{\delta}^{T\geq 3}_{2}$.
+Those estimators are consistent under the following parallel-trends assumption: $t-1$-to-$t$ switchers and stayers with the same $t-1$ treatment would have experienced the same average $t-1$-to-$t$ outcome evolutions if switchers had not switched.
+We propose “pre-trends” or “placebo” estimators, to assess the plausibility of that assumption.
+The placebo AS and WAS estimators are analogous to $\widehat{\delta}^{T\geq 3}_{1}$ and $\widehat{\delta}^{T\geq 3}_{2}$, but they compare the $t-2$-to-$t-1$ outcome evolutions of $t-1$-to-$t$ switchers and stayers, restricting the sample to $t-2$-to-$t-1$ stayers to avoid that those comparisons be contaminated by the treatment’s effect.
+If one finds that from $t-2$-to-$t-1$, $t-1$-to-$t$ switchers and stayers are on parallel trends, this lends credibility to the parallel-trends assumption underlying $\widehat{\delta}^{T\geq 3}_{1}$ and $\widehat{\delta}^{T\geq 3}_{2}$.
+With $T>2$ time periods, the assumption that units’ potential outcomes does not depend on their past treatments is of essence.
+If there are dynamic treatment effects, $\widehat{\delta}^{T\geq 3}_{1}$ and $\widehat{\delta}^{T\geq 3}_{2}$ may be biased.
+Those estimators compare the outcome evolutions of $t-1$-to-$t$ switchers and stayers.
+But there may be, say, $t-1$-to-$t$ stayers whose treatment changed from $t-2$ to $t-1$.
+If lagged treatments affect units’ current outcome, that change could still affect the $t-1$-to-$t$ outcome evolution of those stayers, thus leading them to violate the parallel-trends assumption underlying $\widehat{\delta}^{T\geq 3}_{1}$ and $\widehat{\delta}^{T\geq 3}_{2}$.
+To mitigate this concern, we propose the following robustness check.
+One can recompute $\widehat{\delta}^{T\geq 3}_{1}$ and $\widehat{\delta}^{T\geq 3}_{2}$, restricting, for each pair of consecutive periods $(t-1,t)$, the estimation sample to $t-2$-to-$t-1$ stayers, as in the placebo analysis.
+We show that the resulting estimator is robust to dynamic effects up to one treatment lag.
 Similarly, if one wants to allow for effects of the first and second treatment lags on the outcome, one just needs to restrict the estimation sample to $t-3$-to-$t-1$ stayers.
 However, the more robustness to dynamic effects one would like to have, the smaller the estimation sample becomes.
+5.3 Identification and estimation with covariates
+Our methods can be generalized to incorporate covariates.
+Consider for simplicity the case with two periods, $T=2$, where $X_{t}$ is a $c$-dimensional vector of covariates.
+We show in Section LABEL:appendixsec:covariates of the web appendix that the AS and WAS can be identified and estimated under an alternative parallel trends assumption that holds conditional on the first-period covariates $X_{1}$ (Assumption LABEL:hyp:parallel_trendsX).
+Because our estimation and inference methods are fully nonparametric, they can be readily applied to the case with covariates after redefining the nuisance functions to depend on the covariates, for instance, $\mu_{0}(x,d)=E(\Delta Y|X_{1}=x,D_{1}=d,S=0)$ and so on.
+The main difference is that with covariates, the nuisance functions are nonparametric objects of dimension $c+1$, so classical nonparametric estimators, such as series estimators, may not satisfy the convergence rate requirements for asymptotic normality of the AS and WAS estimators.
+Then, it may be preferable to use machine-learning estimators to estimate those nuisance functions.
 6 Application
-Data and research questions.
+6.1 Data and design
 We use the yearly 1966-to-2008 panel dataset of Li et al.
 (2014), covering 48 US states (Alaska and Hawaii are excluded).
-In view of the long duration of this panel, it is important to keep in mind that our estimators only assume parallel trends across pairs of consecutive years, not over the panel’s entire duration.
+This is a long panel, but recall that our estimators assume parallel trends only across pairs of consecutive years.
 For each state×$\times$×year cell $(i,t)$, the data contains $Z_{i,t}$, the total (state plus federal) gasoline tax in cents per gallon, $D_{i,t}$, the log tax-inclusive price of gasoline, and $Y_{i,t}$, the log gasoline consumption per adult.
-Our goal is to estimate the effect of gasoline taxes on gasoline consumption and prices, and to estimate the price-elasticity of gasoline consumption, using taxes as an instrument.
-Instead, Li et al.
+We seek to estimate the effect of taxes on gasoline consumption and prices, and the price-elasticity of gasoline consumption using taxes as an instrument.
+[^8] Instead, Li et al.
 (2014) jointly estimate the effect of gasoline taxes and tax-exclusive prices on consumption, using a TWFE regression with two treatments.
@@
 (2014).
-Switching cells, and how they compare to the entire sample.
 Let $\mathcal{S}$ be the set of switching $(i,t)$ cells such that $Z_{i,t}\neq Z_{i,t-1}$ but $Z_{i^{\prime},t}=Z_{i^{\prime},t-1}$ for some $i^{\prime}$.
 The second condition drops from the estimation seven pairs of consecutive time periods between which the federal gasoline tax changed, thus implying that all states experienced a change of their tax.
-$\mathcal{S}$ includes 384 cells, so effects of taxes on gasoline prices and consumptions can be estimated for 19% of the 2,016 state×$\times$×year cells for which $Z_{i,t}-Z_{i,t-1}$ can be computed.
-Table 1 below compares some observable characteristics of switchers and stayers.
-Switchers seem slightly over-represented in the later years of the panel: $t$ is on average 2.5 years larger for switchers than for stayers, and the difference is significant.
-On the other hand, switchers are not more populated than stayers, and their gasoline consumption and gasoline price in 1966 are not significantly different from that of stayers.
-Thus, there is no strong indication that the cells in $\mathcal{S}$ are a very selected subgroup.
-Table 1: Comparing switchers and stayers
-Dependent Variables: $t$ Adult Population log(quantity)1966 log(price)1966
-Constant 1,986.7 3,691,608.0 -0.5161 3.471
-(0.2739) (577,164.0) (0.0210) (0.0054)
-$\mathbf{1}\{Z_{i,t}\neq Z_{i,t-1}\}$ 2.481 39,588.0 -0.0099 0.0014
-(0.7519) (320,342.1) (0.0096) (0.0029)
-N 2,016 2,016 2,016 2,016
-Notes: The table show the results of regressions of some dependent variables on a constant and an indicator for switching cells.
-The standard errors shown in parentheses are clustered at the state level.
-Distribution of taxes.
-As an example, the top panel of Figure 1 below shows the distribution of $Z_{g,1987}$ for 1987-to-1988 stayers, while the bottom panel shows the distribution for 1987-to-1988 switchers.
-The figure shows that there are many values of $Z_{g,1987}$ such that only one or two states have that value, so $Z_{g,1987}$ is close to being continuously distributed.
-Moreover, all switchers $g$ are such that
-$\min_{g^{\prime}:Z_{g^{\prime},1988}=Z_{g^{\prime},1987}}Z_{g^{\prime},1987}\leq Z_{g,1987}\leq\max_{g^{\prime}:Z_{g^{\prime},1988}=Z_{g^{\prime},1988}}Z_{g^{\prime},1987}.$
+$\mathcal{S}$ includes 384 cells, namely 19% of the 2,016 state×$\times$×year cells for which $Z_{i,t}-Z_{i,t-1}$ can be computed.
+Table LABEL:table:balancing in the web appendix shows that cells in $\mathcal{S}$ are fairly similar to other cells on a number of observable characteristics.
+Thus, there is no strong indication that the effects we estimate below apply to a very selected subgroup.
+As an example, the top panel of Figure 1 below shows the distribution of $Z_{i,1987}$ for 1987-to-1988 stayers, while the bottom panel shows the distribution for 1987-to-1988 switchers.
+The figure shows that there are many values of $Z_{i,1987}$ such that only one or two states have that value, so $Z_{i,1987}$ is close to being continuously distributed.
+Moreover, all switchers $i$ are such that
+$\min_{i^{\prime}:Z_{i^{\prime},1988}=Z_{i^{\prime},1987}}Z_{i^{\prime},1987}\leq Z_{i,1987}\leq\max_{i^{\prime}:Z_{i^{\prime},1988}=Z_{i^{\prime},1988}}Z_{i^{\prime},1987}.$
 Thus, Assumption 4 seems to hold for this pair of years.
@@
 While $Z_{i,t}$ varies less across states in the first years of the panel, there are many other years where $Z_{i,t}$ is close to being continuously distributed.
-Similarly, almost 95% of cells in $\mathcal{S}$ are such that $\min_{g^{\prime}:Z_{i^{\prime},t}=Z_{i^{\prime},t-1}}Z_{i^{\prime},t-1}\leq Z_{i,t-1}\leq\max_{g^{\prime}:Z_{i^{\prime},t}=Z_{i^{\prime},t-1}}Z_{i^{\prime},t-1}$.
+Similarly, almost 95% of cells in $\mathcal{S}$ are such that $\min_{i^{\prime}:Z_{i^{\prime},t}=Z_{i^{\prime},t-1}}Z_{i^{\prime},t-1}\leq Z_{i,t-1}\leq\max_{i^{\prime}:Z_{i^{\prime},t}=Z_{i^{\prime},t-1}}Z_{i^{\prime},t-1}$.
 Dropping the few cells that do not satisfy this condition barely changes the results presented below.
 Figure 1: Gasoline tax in 1987 among 1987-to-1988 switchers and stayers
-Distribution of tax changes.
-Figure 2 below shows the distribution of $Z_{i,t}-Z_{i,t-1}$ for the 384 cells in $\mathcal{S}$.
-The majority experience an increase in their taxes, but 38 cells experience a decrease.
+Turning to the distribution of the tax changes $Z_{i,t}-Z_{i,t-1}$, 90% of the cells in $\mathcal{S}$ experience an increase in their taxes, and 10% experience a decrease.
 The average value of $|Z_{i,t}-Z_{i,t-1}|$ is equal to 1.61 cents, while prior to the tax change, switchers’ average gasoline price is equal to 112 cents: our estimators leverage small changes in taxes relative to gasoline prices.
 Finally, $\min_{(i,t)\in\mathcal{S}}|Z_{i,t}-Z_{i,t-1}|=0.05$ : some switchers experience a very small change in their taxes.
-Figure 2: Distribution of tax changes between consecutive periods
-Reduced-form and first-stage AS and WAS estimates.
-Table 2 below shows the AS and doubly-robust WAS estimates of the reduced-form (Panel A) and first-stage (Panel B) effects of taxes on quantities and prices.
-We follow results from Section 5, and estimate the reduced-form and the first-stage controlling for lagged prices $D_{t-1}$, to ensure that the resulting IV estimator is robust to heterogeneous effects over time.
-Reduced-form and first-stage estimators where $D_{t-1}$ is not controlled for are not very different, but controlling for $D_{t-1}$ reduces the standard error of the first-stage estimator.
-In Column (1), the estimators are computed using a polynomial of order 1 in $(Z_{t-1},D_{t-1})$ to estimate $E(\Delta Y_{t}|Z_{t-1},D_{t-1},S_{t}=0)$, $E(\Delta D_{t}|Z_{t-1},D_{t-1},S_{t}=0)$, and the propensity scores $P(S_{+,t}=1|Z_{t-1},D_{t-1})$, $P(S_{-,t}=1|Z_{t-1},D_{t-1})$, and $P(S_{t}=0|Z_{t-1},D_{t-1})$.
-In Column (2), a polynomial of order 2 is used in those estimations.
-10-folds cross-validation selects a polynomial of order two for $E(\Delta D_{t}|Z_{t-1},D_{t-1},S_{t}=0)$, and a polynomial of order one for all the other conditional expectations.
-Thus, polynomials of order 1 and 2 are in line with those selected by cross validation.
-Standard errors clustered at the state level, computed following (18) and (19), are shown below the estimates, between parentheses.
-All estimations use 1632 (48$\times$ 35) first-difference observations: 7 periods have to be excluded as they do not have stayers.
-Finally, the last line of each panel shows the p-value of a test that the AS and WAS effects are equal.
-In Panel A Column (1), the AS estimate indicates that increasing gasoline tax by 1 cent decreases quantities consumed by 0.55 percent on average for the switchers.
-That effect is significant at the 5% level, but it becomes smaller and insignificant when one uses a quadratic model to estimate $E(\Delta Y_{t}|Z_{t-1},D_{t-1},S_{t}=0)$.
-The WAS estimates are slightly lower than, but close to, the AS estimates, and they are significant irrespective of the polynomial order used in the estimation.
-As predicted by Proposition 1, the standard errors of the WAS estimators are almost 3 times smaller than that of the AS estimators.
-Equality tests that the AS and WAS effects are equal are not rejected.
-In Panel B, the AS estimates of the first-stage effect are insignificant.
-Importantly, this implies that an IV-AS estimator of the price elasticity of gasoline consumption cannot be used: this estimator does not have a significant first stage.
-The WAS estimates are significant, and they indicate that if gasoline tax increases by 1 cent on average, prices increase by around 0.5 percent on average for the switchers.
-Again, the differences between the AS and WAS effects of taxes on prices are insignificant.
-Table 2: Effects of gasoline tax on quantities consumed and prices
-(1) Linear model (2) Quadratic model
-Panel A: Reduced-form effect of taxes on quantities consumed
-AS -0.0055 -0.0034
-(0.0027) (0.0032)
-WAS -0.0038 -0.0034
-(0.0010) (0.0011)
-Observations 1,632 1,632
-P-value 0.4482 0.9974
-Panel B: First-stage effect of taxes on prices
-AS 0.0042 0.0047
-(0.0024) (0.0025)
-WAS 0.0056 0.0056
-(0.0009) (0.0008)
-Observations 1,632 1,632
-P-value 0.4729 0.6798
-Notes: All estimators in the table are computed using the data of Li et al.
-(2014).
-Panel A (resp.
-B) shows the AS and doubly-robust WAS estimates of the reduced-form (resp.
-first-stage) effect of taxes on quantities (resp.
+6.2 Reduced-form and first-stage estimates
+Table 6.2 below shows the AS and WAS estimates of the reduced-form (Panel A) and first-stage (Panel B) effects of taxes on gasoline quantities and prices.
+We control for lagged prices $D_{t-1}$, to ensure that the resulting IV estimate is robust to heterogeneous effects over time, as explained in Section 5.1.
+Estimates are computed with cross-fitting, using 10 splits.
+We use a polynomial of order 1 in $(Z_{t-1},D_{t-1})$ to estimate $E(\Delta Y_{t}|Z_{t-1},D_{t-1},S_{t}=0)$, $E(\Delta D_{t}|Z_{t-1},D_{t-1},S_{t}=0)$, $E[S/\Delta D|Z_{t-1},D_{t-1}]$, $P(S_{+,t}=1|Z_{t-1},D_{t-1})$, $P(S_{-,t}=1|Z_{t-1},D_{t-1})$, and $P(S_{t}=0|Z_{t-1},D_{t-1})$.
+10-folds cross-validation selects a polynomial of order one for all those conditional expectations, except for $E(\Delta D_{t}|Z_{t-1},D_{t-1},S_{t}=0)$ where it selects a polynomial of order two.
+Thus, polynomials of order one are in line with those selected by cross validation.
+[^9] In a previous version of this paper (de Chaisemartin et al., 2022), we had shown results very similar to those below with polynomials of order two and without cross fitting.
+With polynomials of order two and cross fitting, estimates become noisy: due to the small sample size, $\widehat{P}(S_{t}=0|Z_{t-1},D_{t-1})$ is close to zero for some $(i,t)$.
+Standard errors clustered at the state level, computed following (LABEL:eq:ASIFT>=3) and (LABEL:eq:WASIFT>=3) in the web appendix, are shown next to the estimates, between parentheses.
+Estimations use 1632 (48$\times$ 35) first-difference observations: 7 periods are excluded as they do not have stayers.
+The last line of each panel shows the p-value of a test that AS=WAS.
+In Panel A, the AS estimate indicates that increasing gasoline tax by 1 cent decreases quantities consumed by 0.43 percent on average for the switchers.
+That effect is insignificant.
+The WAS estimate indicates that increasing gasoline tax by 1 cent decreases quantities consumed by 0.36 percent, an effect that is significant.
+In Panel B, the AS estimate of the first-stage effect is positive and insignificant.
+The WAS estimate is significant, and indicates that if gasoline tax increases by 1 cent on average, prices increase by 0.58 percent.
+In both panels, the AS estimates are close to, and not significantly different from, the WAS estimates, but they are insignificant because they are markedly less precise: their standard errors are almost three times larger, as predicted by Proposition 1.
+Therefore, in what follows we focus on the WAS estimates.
+Table 1: Effects of gasoline tax on quantities and prices
+Notes: Panel A (resp.
+B) shows the AS and WAS estimates of the reduced-form (resp.
+first-stage) effect of taxes on gasoline quantities (resp.
 prices).
 All estimates control for the lag of prices.
-In Column (1), estimates are computed using a polynomial of order 1 in $(Z_{t-1},D_{t-1})$ to estimate $E(\Delta Y_{t}|Z_{t-1},D_{t-1},S_{t}=0)$ and the propensity scores $P(S_{+,t}=1|Z_{t-1},D_{t-1})$, $P(S_{-,t}=1|Z_{t-1},D_{t-1})$, and $P(S_{t}=0|Z_{t-1},D_{t-1})$.
-In Column (2), estimates are computed using a polynomial of order 2 in those estimations.
-Standard errors clustered at the state level, computed following (18) and (19) are shown below the estimates, between parentheses.
-All estimations use 1632 (48$\times$ 35) first-difference observations: 7 periods have to be excluded as they do not have stayers.
-Finally, the last line of each panel shows the p-value of a test that the AS and WAS effects are equal.
-Placebo analysis.
-Table 3 below shows placebo AS and doubly-robust WAS estimates of the reduced-form and first-stage effects.
-The placebo estimators are analogous to the actual estimators, but they replace $\Delta Y_{t}$ by $\Delta Y_{t-1}$, and they restrict the sample, for each pair of consecutive time periods $(t-1,t)$, to states whose taxes did not change between $t-2$ and $t-1$.
-The placebo WAS estimates are small and insignificant, both for quantities and prices.
-The placebo AS estimates are larger for quantities, but they are insignificant, and less precisely estimated.
-This placebo analysis shows that before switchers change their gasoline taxes, switchers’ and stayers’ consumption of gasoline and gasoline prices do not follow detectably different evolutions.
-As a robustness check, we reestimate the AS and WAS in the placebo subsample, to ensure that effects are estimated in a subsample for which the identifying assumption can be placebo tested, and also because in that subsample estimators remain valid if the first lag of taxes affect current gasoline prices and quantities.
-WAS reduced-form effects are very close to those in Table 2.
-WAS first-stage effects are 25 to 35% smaller, though they are still positive and highly significant.
-Table 3: Placebo effects of gasoline tax on quantities consumed and prices
-(1) Linear model (2) Quadratic model
-Panel A: Reduced-form placebo effect of taxes on quantities consumed
-AS 0.0039 0.0055
-(0.0035) (0.0036)
-WAS 0.0001 0.0012
-(0.0017) (0.0017)
-Observations 1,059 1,059
-Panel B: First-stage placebo effect of taxes on prices
-AS 0.0006 0.0009
-(0.0056) (0.0053)
-WAS 0.0014 0.0013
-(0.0017) (0.0015)
-Observations 1,059 1,059
-Notes: The table shows the placebo AS and doubly-robust WAS estimates of the reduced-form and first-stage effects of taxes on quantities and prices.
-The estimators and their standard errors are computed as the actual estimators, replacing $\Delta Y_{t}$ by $\Delta Y_{t-1}$, and restricting the sample, for each pair of consecutive time periods $(t-1,t)$, to states whose taxes did not change between $t-2$ and $t-1$.
-IV-WAS estimate of the price-elasticity of gasoline consumption.
-Table 4 shows doubly-robust IV-WAS estimates of the price-elasticity of gasoline consumption.
-As the instrument’s first stage is not very strong and the sample effectively only has 48 observations, asymptotic approximations may not be reliable for inference.
-In line with that conjecture, we find that the bootstrap distributions of the three estimators in Table 4 are non-normal, with some outliers.
-Therefore, we use the percentile bootstrap for inference, clustering the bootstrap at the state level.
-Reassuringly, these confidence intervals have nominal coverage in simulations tailored to our application.
+All conditional expectations are estimated using a polynomial of order 1 in $(Z_{t-1},D_{t-1})$, with cross-fitting, using ten splits.
+Standard errors clustered at the state level are shown next to the estimates, between parentheses.
+The last line of each panel shows the p-value of a test that the AS and WAS effects are equal.
+6.3 Placebo analysis
+Table LABEL:table:placeboRFandFS below shows “pre-trends” or “placebo” WAS estimates of the reduced-form and first-stage effects.
+The placebos are analogous to the actual estimates, but they replace $\Delta Y_{t}$ by $\Delta Y_{t-1}$, and they restrict the sample to states whose taxes did not change between $t-2$ and $t-1$.
+This reduces the sample size by around a third.
+The placebos are small and insignificant, both for quantities and prices.
+This shows that before switchers change their gasoline taxes, switchers’ and stayers’ consumption and prices do not follow detectably different evolutions.
+In Table LABEL:table:RFandFS_placebosample of the web appendix we reestimate the WAS in the placebo subsample.
+In that subsample, the WAS remains valid if the first lag of taxes affect current gasoline prices and quantities, as discussed in Section 5.2.
+The first-stage and reduced-form estimates therein are close to those in Table 6.2.
+Notes: The table shows placebo WAS estimates of the reduced-form and first-stage effects of taxes on quantities and prices.
+The estimates and their standard errors are computed as the actual estimates, replacing $\Delta Y_{t}$ by $\Delta Y_{t-1}$, and restricting the sample to states whose taxes did not change between $t-2$ and $t-1$.
+6.4 IV estimate
+Table LABEL:table:IV shows an IV-WAS estimate of the price-elasticity of gasoline consumption.
+As the instrument’s first stage is not very strong and the sample only has 48 observations, asymptotic approximations may not be reliable for inference.
+In line with that conjecture, we find that the bootstrap distribution of the estimator in Table LABEL:table:IV is non-normal, with some outliers.
+Therefore, for inference we use a percentile bootstrap confidence interval, clustered at the state level.
+Reassuringly, that confidence interval has close to nominal coverage in simulations tailored to our application.
 [^10] Here is the DGP used in our simulations.
@@
 We let $\hat{\gamma}^{Y}_{i}+\hat{\lambda}^{Y}_{t}+\hat{\beta}^{Y}Z_{i,t}+\epsilon^{Y}_{i,t}$ and $\hat{\gamma}^{D}_{i}+\hat{\lambda}^{D}_{t}+\hat{\beta}^{D}Z_{i,t}+\epsilon^{D}_{i,t}$ denote the resulting regression decompositions.
-In each simulation, the simulated instrument is just the actual instrument, while the simulated outcomes and treatments are respectively equal to $Y^{s}_{i,t}=\hat{\gamma}^{Y}_{i}+\hat{\lambda}^{Y}_{t}+\hat{\beta}^{Y}Z_{i,t}+\epsilon^{Y,s}_{i,t},$ and $D^{s}_{i,t}=\hat{\gamma}^{D}_{i}+\hat{\lambda}^{D}_{t}+\hat{\beta}^{D}Z_{i,t}+\epsilon^{D,s}_{i,t},$ where the vector of simulated residuals $(\epsilon^{Y,s}_{g,1},...,\epsilon^{Y,s}_{g,T},\epsilon^{D,s}_{g,1},...,\epsilon^{D,s}_{g,T})$ is drawn at random and with replacement from the estimated vectors of residuals $((\epsilon^{Y}_{g^{\prime},1},...,\epsilon^{Y}_{g^{\prime},T},\epsilon^{D}_{g^{\prime},1},...,\epsilon^{D}_{g^{\prime},T}))_{g^{\prime}\in\{1,...,G\}}$.
+In each simulation, the simulated instrument is just the actual instrument, while the simulated outcomes and treatments are respectively equal to $Y^{s}_{i,t}=\hat{\gamma}^{Y}_{i}+\hat{\lambda}^{Y}_{t}+\hat{\beta}^{Y}Z_{i,t}+\epsilon^{Y,s}_{i,t},$ and $D^{s}_{i,t}=\hat{\gamma}^{D}_{i}+\hat{\lambda}^{D}_{t}+\hat{\beta}^{D}Z_{i,t}+\epsilon^{D,s}_{i,t},$ where the vector of simulated residuals $(\epsilon^{Y,s}_{i,1},...,\epsilon^{Y,s}_{i,T},\epsilon^{D,s}_{i,1},...,\epsilon^{D,s}_{i,T})$ is drawn at random and with replacement from the estimated vectors of residuals $((\epsilon^{Y}_{i^{\prime},1},...,\epsilon^{Y}_{i^{\prime},T},\epsilon^{D}_{i^{\prime},1},...,\epsilon^{D}_{i^{\prime},T}))_{i^{\prime}\in\{1,...,n\}}$.
 Thus, the first-stage and reduced-form effects, the correlation between the reduced-form and first-stage residuals, and the residuals’ serial correlation are the same as in the sample.
-The IV-WAS estimates are negative, significant, and larger than -1, though their confidence intervals contain -1.
-We compare those estimates to a 2SLS-TWFE estimator, computed via a 2SLS regression of $Y_{i,t}$ on $D_{i,t}$ and state and year fixed effects, using $Z_{i,t}$ as the instrument.
-The 2SLS-TWFE coefficient is equal to -1.0836 (bootstrap confidence interval=$[-2.1207,-0.4405]$), which is 60% larger in absolute value than the IV-WAS estimate in Column (1), and almost 80% larger than that in Column (2), though the 2SLS-TWFE coefficient does not significantly differ from the two IV-WAS estimates (P-value = 0.320 and 0.232 , respectively).
-Interestingly, the confidence interval of the 2SLS-TWFE coefficient is almost 80% wider than that of the IV-WAS coefficient in Column (1) and 27% wider than that in Column (2), thus showing that using a more robust estimator does not always come with a substantive precision cost.
-Table 4: IV estimators of the price-elasticity of gasoline consumption
-(1) Linear model (2) Quadratic model
-IV-WAS -0.6773 -0.6130
-[-1.2101,-0.2622] [-1.3183,-0.0004]
-Observations 1,632 1,632
-Notes: The table shows doubly-robust IV-WAS estimates of the price-elasticity of gasoline consumption, computed using the data of Li et al.
-(2014).
-Bootstrap confidence intervals are shown below the estimates.
-They are computed with 500 bootstrap replications, clustered at the state level.
+The IV-WAS estimate is negative and significant at any conventional level.
+It is more negative than, but not significantly different from, estimates reported in recent literature (Coglianese et al., 2017).
+Notes: The table shows an IV-WAS estimate of the price-elasticity of gasoline consumption, controlling for the lag of prices.
+All conditional expectations are estimated using a polynomial of order 1 in $(Z_{t-1},D_{t-1})$, with cross-fitting, using ten splits.
+A percentile-bootstrap confidence interval, computed with 500 bootstrap replications clustered at the state level, is shown below the estimate.
 7 Conclusion
-We propose new difference-in-difference (DID) estimators for continuous treatments.
+We propose new, doubly-robust difference-in-difference (DID) estimators for treatments continuously distributed at every time period.
 We assume that between pairs of consecutive periods, the treatment of some units, the switchers, changes, while the treatment of other units, the stayers, does not change.
-We propose a parallel trends assumption on the outcome evolution of switchers and stayers with the same baseline treatment.
+We propose a parallel-trends assumption on the outcome evolution of switchers and stayers with the same baseline treatment.
 Under that assumption, two target parameters can be estimated.
@@
 Our second target is a weighted average of switchers’ slopes, where switchers receive a weight proportional to the absolute value of their treatment change, referred to as the WAS.
-Economically, the AS and WAS serve different purposes, so neither parameter dominates the other.
-On the other hand, when it comes to estimation, the WAS unambiguously dominates the AS.
+The AS and WAS serve different purposes.
+When it comes to estimation, the WAS has two advantages.
 First, it can be estimated at the parametric rate even if units can experience an arbitrarily small treatment change.
-Second, under some conditions, its asymptotic variance is strictly lower than that of the AS estimator.
-Third, unlike the AS, it is amenable to doubly-robust estimation.
+Second, under some conditions, the asymptotic variance of the WAS estimator is strictly lower than that of the AS.
 In our application, we use US-state-level panel data to estimate the effect of gasoline taxes on gasoline consumption.
@@
 Then, we propose widely applicable IV-DID estimators, that are robust to heterogeneous effects over time but impose some restrictions on effects’ heterogeneity across units.
+Throughout, we assume that there are some stayers, namely units whose treatment does not change between consecutive periods.
+de Chaisemartin et al.
+(2024) discuss the extension of the results in this paper to applications without stayers.